### v1

# Online Flow Time Minimization: Tight Bounds for Non-Preemptive Algorithms

Abstract
This paper studies the classical online scheduling problem of minimizing total flow time for $n$ jobs on $m$ identical machines. Prior work often cites the $\Omega(n)$ lower bound for non-preemptive algorithms to argue for the necessity of preemption or resource augmentation, which shows the trivial $O(n)$-competitive greedy algorithm is tight. However, this lower bound applies only to deterministic algorithms in the single-machine case, leaving several fundamental questions unanswered. Can randomness help in the non-preemptive setting, and what is the optimal online deterministic algorithm when $m\geq 2$? We resolve both questions. We present a polynomial-time randomized algorithm with competitive ratio $\Theta(\sqrt{n/m})$ and prove a matching randomized lower bound, settling the randomized non-preemptive setting for every $m$. This also improves the best-known offline approximation ratio from $O(\sqrt{n/m}\log(n/m))$ to $O(\sqrt{n/m})$. On the deterministic side, we present a non-preemptive algorithm with competitive ratio $O(n/m^{2}+\sqrt{n/m}\log m)$ and prove a nearly matching lower bound.
Our framework also extends to the kill-and-restart model, where we reveal a sharp transition of deterministic algorithms: we design an asymptotically optimal algorithm with the competitive ratio $O(\sqrt{n/m})$ for $m\geq 2$, yet establish a strong $\Omega(n/\log n)$ lower bound for $m=1$. Moreover, we show that randomization provides no further advantage, as the lower bound coincides with that of the non-preemptive setting.
While our main results assume prior knowledge of $n$, we also investigate the setting where $n$ is unknown. We show kill-and-restart is powerful enough to break the $O(n)$ barrier for $m\geq 2$ even without knowing $n$. Conversely, we prove randomization alone is insufficient, as no algorithm can achieve an $o(n)$ competitive ratio in this setting.
1 Introduction
The scheduling problem is a core combinatorial optimization problem across various disciplines in computer science. Its significance stems not only from its wide range of applications but also from its simple yet representative formulation. Numerous variants of scheduling problems exist, motivated by diverse real-world scenarios, such as different machine types (identical, related, or unrelated) and varying objectives (makespan, completion time, flow time). A prominent line of research in scheduling focuses on making irrevocable decisions for jobs that arrive online over time — a framework known as online scheduling, which was initiated in the 1960s by graham1966bounds. Consider an online service platform that must dispatch jobs submitted by multiple users as they arrive; this naturally leads to the objective of minimizing the total flow time, the sum over jobs of their completion time minus release time.
In this paper, we focus on the flow time minimization problem on $m$ identical machines. Most studies on the flow time objective allow additional capabilities for algorithms, such as preemption, resource augmentation, and rejection. Many papers (e.g., [DBLP:conf/focs/GargK07, DBLP:conf/esa/LucarelliTST16, DBLP:conf/stoc/ChadhaGKM09, DBLP:conf/soda/ChenIP25, DBLP:conf/icalp/GuptaKL18, DBLP:journals/dam/EpsteinS06]) have argued that they are necessary to surpass the trivial greedy algorithm [^1] The trivial greedy algorithm simply schedules any waiting job whenever any machine becomes available, achieving a competitive ratio of $\Theta(n/m)$. because of the following single-machine lower bound consisting of $n$ jobs: the adversary releases a unit job at time 0 and, upon scheduling, reveals n1n-1 $\varepsilon$-jobs, forcing any deterministic non-preemptive algorithm to incur $\Omega(n)$ flow time while the offline optimum remains $O(1)$ [DBLP:journals/siamcomp/KellererTW99, DBLP:conf/stoc/ChekuriKZ01]. However, this lower bound applies only to deterministic algorithms in the single-machine setting. The limitations of this lower bound motivate our exploration of the following open questions.
Utilizing Randomness.
Randomization is widely employed in the design of online algorithms. Against an oblivious adversary who does not adapt to the randomness used by the algorithm, randomization often breaks deterministic barriers, including the classic caching problem and matching problem [DBLP:journals/jal/FiatKLMSY91, DBLP:conf/stoc/KarpVV90].
In our problem, the previously known lower bound $\Omega(\sqrt{n})$ for oblivious adversaries by DBLP:journals/tcs/EpsteinS03 applies only to single-machine; no information-theoretic lower bound stronger than $\Omega(1)$ was previously known when $m\geq 2$. A gap remains between the current best-known $\Theta(n/m)$-competitive trivial greedy algorithm and the largely unanswered questions about lower bounds.
Question for Randomized Algorithms Can we use randomization to beat the trivial greedy algorithm? Is it possible to find an algorithm much better than $\Theta(\sqrt{n})$ (perhaps even a constant) when $m\geq 2$?
Deterministic Multi-Machine Case.
In fact, even without randomization, the problem remains unsolved when $m$ is large: the deterministic lower bound $\Omega(n/m^{2})$ established by DBLP:journals/dam/EpsteinS06 decreases much faster than the trivial upper bound $\Theta(n/m)$ achieved by the greedy algorithm, leaving a substantial gap when $m$ is comparable to $n$.
Question for Deterministic Algorithms Is greedy indeed the best deterministic algorithm for multi-machine case? What is the optimal competitive ratio when $m$ is large?
Kill-and-Restart Model.
Another direction that is closely related to the non-preemptive setting is the kill-and-restart capability, first introduced by DBLP:journals/siamcomp/ShmoysWW95, where the scheduler is allowed to kill a job and restart it from scratch, losing all processed work. Kill-and-restart lies between preemption and non-preemption. This model is particularly suitable in settings where partially processed work cannot be resumed due to system or application constraints (e.g., data corruption or non-checkpointable tasks). It also has a practical advantage over preemption, as it avoids the need to maintain and store potentially unbounded intermediate states of suspended jobs. This notion has been studied under objectives such as makespan and total completion time, where it has been shown to surpass previously known lower bounds [DBLP:conf/approx/0002KTWZ18, DBLP:journals/mp/JagerSWW25, DBLP:journals/jal/SteeP05]. Notably, DBLP:journals/mp/JagerSWW25 considered online non-clairvoyant scheduling with the objective of minimizing total completion time and demonstrated that kill-and-restart significantly improves the competitive ratio from $\Omega(n)$ to $O(1)$.
Unlike the completion time objective where kill-and-restart has proven powerful, its potential for flow time minimization has remained largely unexplored. No algorithms were previously known to utilize the kill-and-restart capability in the flow time minimization problem, and the only known lower bound $\Omega(\sqrt{n})$ by DBLP:journals/tcs/EpsteinS03 for the kill-and-restart model is also restricted to the single-machine case. Since the only known algorithm without preemption is the $O(n/m)$-competitive greedy algorithm, it remained unclear whether the kill-and-restart capability can improve competitiveness in flow-time minimization, and to what extent such improvement is possible.
Question for Kill-and-Restart Can we use kill-and-restart to beat the trivial greedy algorithm? Is it possible to find an algorithm much better than $\Theta(\sqrt{n})$ (perhaps even a constant) when $m\geq 2$?
1.1 Our Results
We give a complete set of answers to the above questions, summarized in Table 1 (the single-machine case) and Table 2 (the multi-machine case). A key technical feature of our lower bounds in the multi-machine case is their robustness: unlike bounds that only apply to specific choices of these parameters, they hold for every fixed pair of $(n,m)$.
The results presented in this section assume prior knowledge of $n$. This assumption is common when we study the non-preemptive version of the flow time minimization problem [DBLP:journals/dam/EpsteinS06, DBLP:journals/tcs/EpsteinS03], and we will further discuss this in Section 1.2 and present results on the algorithms without prior knowledge of $n$.
Randomized Algorithms: Beating the Deterministic Barrier
We first address whether randomization can overcome the $\Omega(n)$ deterministic barrier for non-preemptive algorithms. We answer this affirmatively by presenting an $O(\sqrt{n})$-competitive randomized algorithm for the single-machine case, matching the established lower bound by DBLP:journals/tcs/EpsteinS03. We then extend this to the multi-machine setting, achieving an $O(\sqrt{n/m})$ competitive ratio. This result is optimal, as we prove a matching $\Omega(\sqrt{n/m})$ lower bound that holds for every pair of $(n,m)$. This completely characterizes the power of randomization in the non-preemptive setting.
Deterministic Algorithms: Closing the Multi-Machine Gap
[…]
The Power of Kill-and-Restart
Perhaps our most surprising finding is non-uniform impact of the kill-and-restart capability on deterministic algorithms.
  • For the single-machine case, we show that kill-and-restart offers little advantage, proving a strong lower bound of $\Omega(n/\log n)$.
  • However, with just one additional machine, the problem’s competitive ratio collapses. We present a deterministic algorithm with kill-and-restart that achieves an $O(\sqrt{n/m})$ competitive ratio for all $m\geq 2$. This is a quadratic improvement over the single-machine case and remarkably matches the optimal bound achievable with randomization.
Conversely, we show that for randomized algorithms, kill-and-restart provides no additional power, as the $\Omega(\sqrt{n/m})$ lower bound holds even in this stronger model.
Improving the State-of-the-Art for Offline Scheduling
A direct consequence of our online results is a new state-of-the-art for the classic offline problem. We improve the long-standing best approximation ratio from $O(\sqrt{n/m}\log(n/m))$ by DBLP:journals/jcss/LeonardiR07 to $O(\sqrt{n/m})$. The improved bound can be realized by either derandomizing our randomized algorithm or applying our deterministic kill-and-restart algorithm in an offline context where restarts can be ignored at no cost.
Table 1: Summary of competitive ratios for the single-machine setting.
Single machine, $n$ jobs Non-Preemptive Kill-and-Restart
Deterministic Upper Bound $O(n)$ $O(n)$
(Trivial) (Trivial)
Lower Bound $\Omega(n)$ $\Omega(n/\log n)$
[DBLP:journals/siamcomp/KellererTW99, DBLP:conf/stoc/ChekuriKZ01] Theorem 7
Randomized Upper Bound $O(\sqrt{n})$ $O(\sqrt{n})$
Theorem 1 Theorem 1
Lower Bound $\Omega(\sqrt{n})$ $\Omega(\sqrt{n})$
[DBLP:journals/tcs/EpsteinS03] [DBLP:journals/tcs/EpsteinS03]
Table 2: Summary of competitive ratios for the multi-machine setting.
$m$ machines, $n$ jobs Non-Preemptive Kill-and-Restart
Deterministic Upper Bound $O(n/m^{2}+\sqrt{n/m}\log m)$ $O(\sqrt{n/m})$ for $m\geq 2$}
Theorem 2 Theorem 3
Lower Bound $\Omega(n/m^{2}+\sqrt{n/m})$ $\Omega(\sqrt{n/m})$
Theorem 8 Theorem 6
Randomized Upper Bound $O(\sqrt{n/m})$ $O(\sqrt{n/m})$
Theorem 10 Theorem 10
Lower Bound $\Omega(\sqrt{n/m})$ $\Omega(\sqrt{n/m})$
Theorem 5 Theorem 6
1.2 Further Discussion of the Prior Knowledge of $n$
In our main results, we follow the common convention in non-preemptive flow-time scheduling by assuming that $n$, the total number of jobs, is known in advance. This assumption is common to nearly all significant prior work in this area, including both upper and lower bounds [DBLP:journals/dam/EpsteinS06, DBLP:journals/tcs/EpsteinS03]. The necessity of this assumption is underscored by strong lower bounds showing that without it, no deterministic algorithm can beat the trivial $O(n)$ competitive ratio, even if the algorithm can use $\ell$ machines but the optimal solution can only use one [DBLP:journals/dam/EpsteinS06].
However, this barrier does not preclude the possibility that randomization or the kill-and-restart capability could yield improvements in the more challenging “unknown nn” setting. We investigate this question and show that these two tools have different capabilities:
  • Randomization alone is insufficient. We prove that for a single machine, no randomized algorithm can achieve a competitive ratio of $o(n)$ without knowing $n$. For the multi-machine setting, we further show that no randomized algorithm can achieve a competitive ratio of $o(n/m^{2}+\sqrt{n/m})$ without prior knowledge of $n$. This demonstrates that knowing $n$ is necessary for designing non-trivial randomized algorithms in the non-preemptive setting.
  • Kill-and-restart overcomes the barrier. In sharp contrast, we show that the kill-and-restart model is powerful enough to break the $O(n)$ barrier when $m\geq 2$. We extend our deterministic kill-and-restart algorithm to operate without knowledge of $n$ and prove that it achieves an $O(n^{\alpha}/\sqrt{m})$-competitive ratio, where $\alpha=(\sqrt{5}-1)/2\approx 0.618$ is the golden ratio. This demonstrates that the kill-and-restart capability provides a fundamental advantage over purely randomized approaches in this setting.
Finally, we note that our algorithms are robust to estimation. If only an approximate estimate $\hat{n}\in[n/c,cn]$ is known instead of the accurate value of $n$, as might be provided by heuristics or machine-learned advice, all algorithms’ guarantees hold within a constant factor of $c$, preserving their asymptotic performance.
1.3 Our Techniques
A central challenge in non-preemptive scheduling is to partition incoming jobs into two manageable groups: a small set of large jobs, and a set of small jobs whose processing times are provably limited. This distinction is key to circumventing the classic $\Omega(n)$ lower bound, where handling a single large job at the wrong time can delay a swarm of subsequent small jobs, leading to high flow time. Offline algorithms can address this by partitioning jobs based on global properties of the total job set (e.g., the offline approximation algorithms by DBLP:journals/siamcomp/KellererTW99 and DBLP:journals/jcss/LeonardiR07 relied on a pre-solved preemptive solution over all jobs, and use the flow time of jobs in the preemptive solution to partition jobs), but this is impossible to implement in an online context, since future jobs are unknown. Therefore, a fundamentally new approach is required for the online setting where decisions are irrevocable and based on incomplete information.
To overcome this barrier, we introduce a new online rank-based partitioning framework, which is the cornerstone of all our algorithms. Instead of using an unknown future metric (like the flow time in the preemptive solution over all jobs) to classify a job’s size, we use its processing time $p_{j}$, an immediately available property. Jobs are dynamically classified based on their rank: at any time, we designate the $k$ jobs with the largest processing times seen so far as large, with all others classified as small. This rank-based distinction allows the algorithm to adapt as new jobs arrive. However, it also introduces two significant technical challenges that we resolve:
  • Managing dynamic re-classification. A job initially classified as large may later become small as larger jobs arrive. If we start treating it as a small job from that point on, we must carefully bound its previous waiting time and analyze whether this transition delays future jobs by effectively leaving one more job behind OPT. To handle this transition gracefully, we introduce the concept of proxy jobs. Instead of re-label a large job when it should become small, we treat it as a newly arrived small job, called a proxy job. We then decompose its total flow time into two parts: the waiting time during its large phase, and the flow time of the proxy job. Note that a proxy job is not an actual job from the original instance, so we must further bound the additional impact introduced by creating such proxy jobs.
  • Scheduling small jobs without a global view. Prior offline approaches use a near-optimal preemptive schedule as a guide, and try to simulate the preemptive schedule of small jobs in a non-preemptive way. However, without the guide, we must handle the stream of small jobs using only local information. We demonstrate that a simple greedy strategy, Non-Preemptive Shortest Job First (NSJF), is remarkably effective. A key part of our technical contribution is a novel and robust analysis of NSJF. We analyze its performance under the influence of blocking periods, where machines may be temporarily occupied by large jobs. This analysis allows us to precisely bound the flow time of small jobs, even with interference.
With this online rank-based partitioning framework in place, subsequent algorithmic design can focus solely on coordinating a bounded number of large jobs with an efficiently scheduled set of small jobs. A key advantage of this framework is that it no longer requires a preemptive algorithm to exploit global properties of the entire job set. This distinction not only enables our algorithm to be implemented online, but also improves the offline approximation ratio by a logarithmic factor over the result of DBLP:journals/jcss/LeonardiR07, since their logarithmic loss arises from relying on a suboptimal preemptive algorithm in the multi-machine setting.
We believe that our online partitioning framework and the analysis of NSJF provide a general technical insight that can be applied to broader classes of non-preemptive scheduling problems, such as the flow time minimization problem in the related machine setting.
1.4 Paper Organization
The remainder of this paper is organized as follows. In Section 3, we introduce our core technical contribution: the online rank-based partitioning framework that serves as the foundation for all our algorithms. The analysis of the NSJF algorithm, used for handling small jobs within this framework, is detailed in Appendix A.
We then present our main algorithmic results. Section 4 develops our $O(\sqrt{n/m})$-competitive randomized non-preemptive algorithm. Section 5 presents our deterministic non-preemptive algorithm. Section 6 details our deterministic $O(\sqrt{n/m})$-competitive algorithm using kill-and-restart for $m\geq 2$. We then address the unknown-$n$ setting in Section 7 with an extension of this algorithm.
Finally, we establish the tightness of our results in Section 8 and Section 9, where we present a comprehensive set of matching lower bounds for both the known-$n$ and unknown-$n$ settings, respectively.
1.5 Related Works
The problem of minimizing total flow time has been a central and extensively studied topic in scheduling theory for decades, spanning a wide range of models and settings.
Non-Preemptive Scheduling.
In the classic non-preemptive offline setting, the problem is known to be NP-hard. A key result by DBLP:journals/jcss/LeonardiR07 provided an $O(\sqrt{n/m}\log(n/m))$-approximation algorithm, which stood as the best-known guarantee for many years. On the approximation hardness front, DBLP:journals/siamcomp/KellererTW99 established a polynomial-time lower bound of $\Omega(\sqrt{n})$ for the single-machine case, later complemented by DBLP:journals/jcss/LeonardiR07, presenting an $\Omega(n^{1/3-\varepsilon})$ lower bound for multiple machines when $n\geq m^{4/\varepsilon}$. In the online setting, the problem is notoriously difficult, with a simple greedy algorithm achieving a $\Theta(n/m)$ competitive ratio, which for a long time was the best-known result for deterministic algorithms.
Preemptive Scheduling.
To circumvent the hardness of non-preemptive scheduling, a significant body of work has focused on the preemptive model, where jobs can be suspended and resumed. In this setting, the folklore SRPT (Shortest Remaining Processing Time) policy is optimal on a single machine. For multiple machines, SRPT achieves a tight competitive ratio of $\Theta(\log(n/m))$ when job migration is allowed [DBLP:journals/jcss/LeonardiR07]. For the more restrictive model without migration, DBLP:journals/siamcomp/AwerbuchALR02 presented an algorithm with a competitive ratio of $O(\log(\min\{\log n,\log P\}))$.
Models with Algorithmic Relaxations.
Another prominent line of research has explored models that grant additional power to the online non-preemptive algorithm. One common approach is resource augmentation, where the algorithm is provided with faster or more machines than the adversary. DBLP:journals/algorithmica/PhillipsSTW02 showed that an $O(1)$-competitive ratio is achievable with $m\log P$ machines, where $P$ is the max-to-min job size ratio. Another popular relaxation is the rejection model, where the algorithm can choose to reject a small fraction of jobs. DBLP:journals/jcss/ChoudhuryDGK18 first introduced this notion for flow time, with subsequent works [DBLP:conf/icalp/GuptaKL18, DBLP:conf/esa/LucarelliMTST18] achieving an $O(1/\varepsilon^{3})$-competitive algorithm for rejecting an $\varepsilon$-fraction of jobs.
The Kill-and-Restart Model.
The kill-and-restart capability, introduced by DBLP:journals/siamcomp/ShmoysWW95 in the context of makespan minimization, represents a minimal form of preemption. This model has been successfully applied to other objectives, demonstrating its power to surpass non-preemptive lower bounds. For minimizing total completion time, DBLP:journals/jal/SteeP05 designed a $3/2$-competitive algorithm, breaking the non-preemptive barrier of $1.582$. Even more dramatically, for non-clairvoyant scheduling, DBLP:journals/mp/JagerSWW25 showed that kill-and-restart reduces the competitive ratio for total completion time from $\Omega(n)$ to $O(1)$. Despite these successes, the potential of kill-and-restart for flow time minimization has remained largely unexplored. The primary work in this specific area was by DBLP:journals/tcs/EpsteinS03, who established several lower bounds, including an $\Omega(\sqrt{n})$ bound for randomized algorithms on a single machine, but no upper bounds were known.
2 Preliminaries
In this section, we formally define the scheduling problem, the algorithmic models under consideration, and the metrics used for performance evaluation.
2.1 Problem Formulation
An instance of our scheduling problem is defined by a set of $m$ identical machines and a set of $n$ jobs, $J=\{1,2,\dots,n\}$. Each job $j\in J$ is characterized by [^3] In our algorithms, when comparing job release times or processing times, we break ties in an arbitrary but fixed manner. For convenience, one may assume that all values are distinct.:
  • A release time $r_{j}$, at which the job becomes available for processing.
  • A processing time $p_{j}$, representing the duration required to complete the job.
A job becomes known to the algorithm at its release time $r_{j}$ and must be assigned to a machine at or after $r_{j}$. The job must be processed on a single machine at a time, and must be processed for $p_{j}$ time in total to complete. One machine can serve at most one job at a time.
We introduce standard notation to describe the status of a job or a machine at a given time $t$. A machine is said to be busy at time $t$ if it is processing a job, and idle otherwise. Time $t$ is said to be busy if all machines are busy at time $t$. A job $j$ is said to be waiting at time $t$ if it has been released but not yet started, i.e., $r_{j}\leq t<s_{j}$. A job is said to be processing at time $t$ if it is being processed by a machine.
2.2 Scheduling Models
A schedule $\S$ assigns a start time $s_{j}\geq r_{j}$ to each job $j$ on one of the $m$ machines. This paper focuses on two primary models:
  • Non-Preemptive: Once a job begins execution on a machine, it must run to completion without interruption.
  • Kill-and-Restart: Algorithms are permitted to terminate (kill) a job that is currently running. The job then returns to the set of waiting jobs and can be scheduled again (restart) from the beginning at a later time, possibly on a different machine. All work done prior to the termination is lost.
For the purpose of analysis, we also refer to the following preemptive models.
  • Preemptive (without Migration): The algorithm is allowed to interrupt a job and later resume it on the same machine without losing any processed work.
  • Preemptive with Migration: The algorithm is allowed to interrupt a job and later resume it on any machine without losing any processed work.
2.3 Objective and Performance Analysis
Let $\S=\textup{{ALG}}(J)$ be the schedule produced by some algorithm ALG on job set $J$. We write $s_{j}(\S)$ for the start time of job $j$ in schedule $\S$, and define its completion time as $C_{j}(\S)=s_{j}(\S)+p_{j}$. The flow time of job $j$ is $F_{j}(\S)=C_{j}(\S)-r_{j}$, and the objective is to minimize the total flow time on the job set $J$:
$F(\S)=\sum_{j\in J}F_{j}(\S).$
When the schedule $\S$ is clear from context, we omit it from the notation and write $s_{j}$, $C_{j}$, and $F_{j}$ instead.
Competitive Analysis.
The performance of an online algorithm is measured by its competitive ratio. An algorithm ALG is $\Gamma$-competitive [^4] The competitive ratio, $\Gamma$, is often a function $\Gamma(n,m)$ of the number of jobs $n$ and machines $m$ in this paper. if for any job instance $J$:
$F(\textup{{ALG}}(J))\leq\Gamma\cdot F(\textup{{OPT}}(J))$
where $\textup{{OPT}}(J)$ is the total flow time of an optimal offline solution. For the purpose of analysis, we define OPT differently for our upper and lower bounds:
  • For our upper bounds, we compare our online algorithm against a more powerful adversary: OPT is the optimal offline preemptive schedule without migration.
  • For our lower bounds, we use the offline non-preemptive schedule as OPT.
For convenience, when the context is clear, we will use OPT and ALG to denote the costs $F(\textup{{OPT}}(J))$ and $F(\textup{{ALG}}(J))$, respectively.
3 Online Rank-Based Partitioning Framework
This section introduces our core technical contribution: a framework for partitioning jobs into large and small categories in an online manner. Unlike offline methods that rely on global knowledge, our approach uses a fixed rank-based threshold, $\ell$, whose value will be set by our algorithms based on $n$ and $m$.
In Section 3.1, we introduce this framework, centered around the primary challenge of dynamic job re-classification and our solution using proxy jobs. We formalize this logic in Algorithm 1 and prove its key properties. In Section 3.2, we show why the resulting set of small jobs is manageable and can be scheduled efficiently using a simple greedy algorithm.
3.1 The Challenge of Dynamic Re-Classification
Our framework is built on a dynamic partitioning strategy designed to maintain two key invariants at all times:
  • The number of active jobs classified as large is bounded by $\ell$.
  • The processing time of any job classified as small is bounded by $\textup{{OPT}}/\ell$.
To achieve this, we maintain a set of active large jobs, denoted AA, containing the $\ell$ jobs with the largest processing times seen so far. The logic of our framework defines a clear life cycle for every job based on its interaction with this set.
Arrival and Classification.
Upon arrival, a job is evaluated. If it is not large enough to enter the top-$\ell$ group, it is permanently classified as a small job. Otherwise, it is designated as an active large job and added to the set AA.
Retirement and the Proxy Mechanism.
The main complexity arises when a job $j\in A$ is forced out (or retired) at time $r_{i}$ due to the arrival of a new, larger job ii. The state of $j$ at this moment determines its new classification:
  • If job $j$ is still waiting, it becomes proxied. A new proxy job $j^{\prime}$ with the same processing time of $j$ is created immediately, which is then handled along with the small jobs. This is the primary mechanism our analysis relies on.
  • If job $j$ has already started processing by the online algorithm, it cannot be proxied, as the online algorithm must not revoke historical scheduling decisions. We call this a committed large job. Although no longer in the active set AA, it must still be tracked and scheduled as a large job. This is a subtle point in the analysis and can be safely ignored for a high-level understanding, but it is an important case required for the correctness of our online algorithm.
Data: Active large job set AA, and every job’s other current states
Function Partition(ii, $\ell$):
 if $|A|<\ell$ then
      Add $i$ to AA;
    return $i$ is large
 else if $p_{i}>\min_{j\in A}p_{j}$ then
      Let $j=\arg\min_{j\in A}p_{j}$;
      Remove $j$ from AA and add $i$ to AA;
    if job $j$ is still waiting then
         Create a new proxy job $j^{\prime}$ with $r_{j^{\prime}}=r_{i}$ and $p_{j^{\prime}}=p_{j}$;
         Mark $j$ as proxied by $j^{\prime}$;
       // The scheduler will actually process $j$ when it processes $j^{\prime}$.
    else
       // $j$ is a committed large job.
      end if
    return $i$ is large
 else
    return $i$ is small
   end if
end
Algorithm 1 Online Rank-Based Partitioning Step
We now present the pseudocode for this partitioning step in Algorithm 1. For simplicity, the function only explicitly manages the active large job set AA; the other sets (small, proxied, and committed jobs) are implicitly defined by the procedure. This partitioning mechanism guarantees our desired invariants along with a key monotonicity property: the set of all small jobs (including proxy jobs) forms an “easier” scheduling instance than the original job set. This is formalized in the following lemma.
Lemma 1.
Under the online rank-based partitioning method, the following properties hold:
  1. 1.
    The number of active large jobs (the set AA) is at most $\ell$ at any time.
  2. 2.
    The processing time of any job classified as small is upper bounded by $\textup{{OPT}}/\ell$.
  3. 3.
    Let S{S} be the set of all jobs ever classified as small, including all proxy jobs. Let $J$ be the original job set. The instance S{S} is no harder to schedule than $J$, meaning $F(\textup{{OPT}}({S}))\leq F(\textup{{OPT}}(J))$.
Proof.
  • (1)
    Holds by construction, as the algorithm explicitly maintains $|A|\leq\ell$.
  • (2)
    Consider any job $j$ classified as small. By definition of the algorithm, at the time of its arrival, there must have been at least $\ell$ other jobs released that were larger than $j$. The total processing time of these $\ell$ larger jobs alone is a lower bound on the optimal solution’s value, thus $p_{j}<\textup{{OPT}}/\ell$.
  • (3)
    This property holds due to the proxy mechanism. The set S{S} differs from $J$ by replacing some proxied large jobs with proxy jobs. Crucially, each proxy job $j^{\prime}$ is created at the exact moment $r_{j^{\prime}}$ that a new, larger job $i\in J$ ($p_{i}>p_{j^{\prime}}$) arrives and displaces it. We can therefore charge the existence of each proxy job in S{S} to a unique, strictly larger job in $J$ with the same release time. This implies that the instance S{S} is a “weaker” instance than $J$, and thus its optimal flow time can be no larger.
3.2 Handling Small Jobs with Non-Preemptive Shortest Job First
The partitioning framework effectively isolates a small, bounded number of active large jobs at any time, allowing our main algorithms to focus on managing them. This leaves the much larger set of small jobs, which must be handled efficiently. In this section, we show that a natural greedy algorithm, Non-Preemptive Shortest Job First (NSJF), is provably effective for this task.
The NSJF algorithm is simple: whenever a machine is idle and there are waiting jobs, it immediately starts processing the one with the smallest processing time. The following lemma guarantees the performance of NSJF on a set of jobs whose sizes are bounded. This result is central to our analysis, as our framework guarantees that all jobs in the set S{S} have bounded size.
Lemma 2.
Consider running NSJF on a set of jobs whose processing times are all bounded by $\tau$. Compared to an optimal offline preemptive schedule (even with migration), the total flow time of NSJF is bounded by: $F(\textup{{NSJF}})\leq\textup{{OPT}}+O(n\tau)$.
Lemma 2 demonstrates the power of this partition by bounding the flow time contribution from this large group of small jobs. For example, when we set the rank threshold $\ell=\sqrt{n}$, our framework ensures that all small jobs have a processing time $\tau\leq\textup{{OPT}}/\sqrt{n}$. For clarity, we will first prove the lemma as stated. In Appendix A, we prove a more general version of this result (Lemma 29) that is robust to interference from large jobs, where machines may be subject to initial “blocking periods.”, during which they are unavailable for scheduling small jobs. Specifically, given a vector of initial blocking times $\vec{b}$ that NSJF cannot use a machine $i$ in its blocking period $0,b_{i})$, we show that NSJF remains competitive with respect to the optimal flow time on an instance without blocking, up to an extra additive factor depending on the total blocking length. This general version is a fundamental component in the final analysis of our main algorithms.
Proof Strategy.
The proof of Lemma 2 follows a three-step structure. First, we show that the total volume of work started by NSJF quickly catches up to the volume of work processed by the optimal solution after a small time shift. Second, we lift this volume-based argument to show that the number of jobs completed by NSJF also catches up to OPT after a slightly larger time shift. Finally, we use this bound on the number of completed jobs to prove the lemma. Note that the general version of Lemma 29 follows the same strategy; however, the analysis must additionally account for how the available processing power varies over time in the presence of blocking periods.
Proof of Lemma 2.
To begin, we introduce notation to compare the progress of NSJF against OPT. For any processing time threshold $p$ and time $t$, let $J_{\leq p}(t)$ be the set of jobs with $p_{j}\leq p$ and release time $r_{j}\leq t$. We measure progress on this set differently for each algorithm:
  • For OPT, let $V_{\leq p}(t)$ denote the total volume of jobs in $J_{\leq p}(t)$ that has been processed by time $t$.
  • For NSJF, let $V^{\prime}_{\leq p}(t)$ denote the total volume of jobs in $J_{\leq p}(t)$ that have started by time $t$.
Our first step is to prove the following volume-matching lemma.
Lemma 3.
For every $t\geq 0$, and every upper bound $p$, we have: $V^{\prime}_{\leq p}(t+\tau)\geq V_{\leq p}(t).$
Proof.
Fix an arbitrary $p$, let $t_{0}$ be the last time some machine is idle by NSJF before $t+\tau$. Define $s_{j}$ as the latest start time, within the interval $(t_{0},\,t+\tau)$, at which NSJF schedules a job with $p_{j}>p$. If no such job exists, we set $s_{j}=t_{0}$ and conceptually assume that an infinitely large job starts at time $s_{j}$.
Since NSJF starts job $j$ with $p_{j}>p$ at time $s_{j}$, it must have already started all jobs with size at most $p$ released up to $s_{j}$ (due to its SJF policy). Thus, we have:
$V^{\prime}_{\leq p}(s_{j})=\sum_{h\in J_{\leq p}(s_{j})}p_{h}\geq V_{\leq p}(s_{j}).$
We now analyze two cases to conclude the lemma. If $s_{j}\geq t$. We straightforwardly have:
$V^{\prime}_{\leq p}(t+\tau)\geq V^{\prime}_{\leq p}(s_{j})\geq V_{\leq p}(s_{j})\geq V_{\leq p}(t).$
Otherwise, if $s_{j}<t$, the time interval $s_{j}+\tau,t+\tau)$ must be fully used by NSJF to process jobs with size at most $p$ (if any). We have:
$V^{\prime}_{\leq p}(t+\tau)\geq V_{\leq p}(s_{j})+m(t-s_{j})\geq V_{\leq p}(t).$
Denote $N_{\leq p}(t)$ (or $N^{\prime}_{\leq p}(t)$ below) as the number of jobs in $J_{\leq p}(t)$ completed by time $t$ by OPT (or NSJF, respectively). Note that without the subscription $\leq p$, we remove the condition on job size. We prove:
Lemma 4.
For every $t\geq 0$, and every upper bound $p$, we have $N^{\prime}_{\leq p}(t+2\tau)\geq N_{\leq p}(t)$.
Proof.
We discuss two timepoints $t+\tau$ and $t+2\tau$. By Lemma 3, we have the volume bound: $V^{\prime}_{\leq p}(t+\tau)\geq V_{\leq p}(t)$ for every pp.
We claim the following: the $i$-th smallest job (in terms of size) that has started (either in processing or completed) under NSJF by time $t+\tau$, denoted by size $p^{\prime}_{i}$, must not exceed the size of the $i$-th smallest completed job under OPT by time $t$, denoted $p_{i}$. That is, we aim to prove: $p^{\prime}_{i}\leq p_{i}$ for all rank ii.
To prove this, assume for contradiction that there exists the smallest rank $i$ such that $p^{\prime}_{i}>p_{i}$. Under this assumption, we have the following bound for OPT:
$V_{\leq p_{i}}(t)\geq\sum_{j=1}^{i}p_{j},$
since at least $i$ jobs of size at most $p_{i}$ are completed by time $t$ in OPT.
Now consider NSJF. Since the $i$-th smallest started job under NSJF has a size strictly greater than $p_{i}$, all started jobs of size $\leq p_{i}$ must be among the first i1i-1 jobs. Therefore, we have:
$V^{\prime}_{\leq p_{i}}(t+\tau)\leq\sum_{j=1}^{i-1}p^{\prime}_{j}\leq\sum_{j=1}^{i-1}p_{j},$
where the second inequality uses the assumption that $p^{\prime}_{j}\leq p_{j}$ for all $j<i$. Combining these two, we get: $V^{\prime}_{\leq p_{i}}(t+\tau)<V_{\leq p_{i}}(t),$ which contradicts the earlier volume bound from Lemma 3. Hence, our assumption must be false, and we conclude: $p^{\prime}_{i}\leq p_{i}$ for all ii.
Finally, this implies that the number of started jobs with $p_{j}\leq p$ by time $t+\tau$ under NSJF is at least the number of completed jobs with $p_{j}\leq p$ by time $t$ under OPT. Since all jobs started by time $t+\tau$ must be completed by time $t+2\tau$, it completes the proof of the lemma.
Proof of lemma 2.
Let $u$ denote the maximum completion time under NSJF. The flow times of NSJF and OPT can be expressed as:
$\textup{{OPT}}=\int_{0}^{\infty}\left(|J(t)|-N(t)\right)\,dt,\quad F(\textup{{NSJF}})=\int_{0}^{u}\left(|J(t)|-N^{\prime}(t)\right)\,dt,$
where $|J(t)|$ is the number of jobs released by time $t$.
$F(\textup{{NSJF}})$ $=\int_{0}^{u}\left(|J(t)|-N^{\prime}(t)\right)\,dt$
$\leq\int_{0}^{u-2\tau}\left(|J(t)|-N^{\prime}(t+2\tau)\right)\,dt+\int_{u-2\tau}^{u}|J(t)|\,dt$
$\leq\int_{0}^{u-2\tau}\left(|J(t)|-N(t)\right)\,dt+2n\tau$
$\leq\textup{{OPT}}+2n\tau.$
4 Randomized Non-Preemptive Algorithm
This section presents our $O(\sqrt{n/m})$-competitive randomized algorithm. We begin by detailing the core ideas in the single-machine setting before sketching the straightforward extension to multiple machines.
4.1 Single-Machine Case
Our algorithm is built on the online partitioning framework (Section 3) with the large-job threshold set to $\ell=\lfloor\sqrt{n}\rfloor$. This ensures that all small jobs have a processing time of at most $\textup{{OPT}}/\sqrt{n}$. The high-level strategy is to combine two principles:
  1. 1.
    Prioritize Small Jobs: A simple greedy strategy, NSJF, is used to schedule all small jobs, ensuring their contribution to the total flow time is bounded by $O(\sqrt{n})\cdot\textup{{OPT}}$ (Lemma 2).
  2. 2.
    Randomly Delay Large Jobs: Each large job is carefully inserted into the schedule created by the small jobs. To minimize the disruption, each large job $j$ is randomly assigned a “patience” level, delaying its start until a certain amount of machine idle time has passed.
Instead of presenting the online algorithm directly, we describe a dynamic algorithm that reconstructs a schedule upon each job’s arrival while keeping the decisions before unchanged. This dynamic view is a powerful tool for analysis, as it allows us to cleanly decompose the total flow time.
Include : Algorithm 1
Data: Every job’s current states as determined by Algorithm 1.
On job $i$ release:
   Classify job $i$ large or small based on Partition($i,\ell$) with $\ell=\lfloor\sqrt{n}\rfloor$ ;
 if $i$ is classified as large then
      Sample $w_{i}\sim\text{Unif}\{1,\dotsc,\lfloor\sqrt{n}\rfloor\}$
   end if
  Run NSJF on all small jobs to get schedule $\S_{1}$ ;
 // Including proxy jobs.
 $\S_{2}\leftarrow\S_{1}$ ;
 for each unproxied (active or committed) large job $j$ in order of release time $r_{j}$ do
    $t\leftarrow$ the first time in $\S_{2}$ where the cumulative idle time since $r_{j}$ is at least $w_{j}p_{j}$;
      Insert $j$ into $\S_{2}$ to start at $t$ ;
   end for
  Schedule jobs according to $\S_{2}$;
Algorithm 2 Dynamic (Online Stable) Non-Preemptive Randomized Algorithm for Single-Machine
The key requirement, online stability, is that the schedule reconstructed at time $t$ must remain consistent with the scheduling decisions made prior to $t$. The proof of this property is given below.
Lemma 5.
Algorithm 2 is online stable; that is, for every time $t$, its schedule before time $t$ depends solely on the jobs released before $t$.
Proof.
Consider the arrival of a new job $i$ at time $r_{i}$, which triggers a reconstruction of the schedule. Let $\S^{\prime}_{1}$ and $\S^{\prime}_{2}$ be the schedules produced in the previous round (before $i$’s arrival), and let $\S_{1}$ and $\S_{2}$ be the newly computed schedules. We must show that for any time $t<r_{i}$, the schedules are identical.
First, consider the base schedule $\S_{1}$. The set of small jobs is only augmented by jobs with release time $r_{i}$ (either $i$ itself, if it is small, or a new proxy job). Since the set of small and proxy jobs released before $r_{i}$ is unchanged, and NSJF is an online algorithm, $\S_{1}$ is identical to $\S^{\prime}_{1}$ on the interval $0,r_{i})$.
Next, consider the final schedule $\S_{2}$. The set of unproxied large jobs changes only if job $i$ is large and retires a waiting large job $j$. In this case, $j$ is part of the for-loop that constructs $\S^{\prime}_{2}$, but it is absent from the loop that constructs $\S_{2}$. However, because $j$ was retired while waiting at time $r_{i}$, its scheduled start time $s^{\prime}_{j}$ in $\S^{\prime}_{2}$ must be greater than or equal to $r_{i}$. The removal of $j$ from the insertion process for $\S_{2}$ can only affect the schedule at or after time $s^{\prime}_{j}$. Any jobs scheduled in $\S^{\prime}_{2}$ before time $s^{\prime}_{j}$ will be scheduled at the same time in $\S_{2}$. Since $s^{\prime}_{j}\geq r_{i}$, the schedule $\S_{2}$ is identical to $\S^{\prime}_{2}$ on the interval $0,r_{i})$.
In all other cases, the set of unproxied large jobs remains the same, so the construction is identical up to time $r_{i}$. Therefore, the schedule is online stable.
Theorem 1.
There exists an online polynomial time randomized non-preemptive algorithm for total flow time minimization on the single-machine case, e.g., Algorithm 2, is $O(\sqrt{n})$-competitive, against the preemptive offline solution.
Proof.
Recall that $F(\S)$ denotes the total flow time of a schedule $\S$. The final schedule produced by Algorithm 2 has total flow time $F(\S_{2})$. Our algorithm constructs this final schedule incrementally by transitioning through intermediate schedules $\S_{1}\rightarrow\S_{2}$. We analyze the additional flow time incurred during this transition. Let $J$ denote the real job set, which includes only actual jobs and excludes proxy jobs. Let $L$ and $S$ denote the sets of large jobs (including all active and retired large jobs) and small jobs (including proxy jobs), respectively. Note that a job belongs to $L$ or $S$ according to its classification, large or small, at its arrival time.
First, $\S_{1}$ is competitive with the optimal schedule for the job set $S$. By Lemma 1, all jobs in $S$ have size at most $\textup{{OPT}}/\lfloor\sqrt{n}\rfloor$. Therefore, by Lemma 2, we have:
$F(\S_{1})\leq O(\sqrt{n})\cdot F(\textup{{OPT}}(S)).$
Moreover, since $S$ is weaker than $J$ by Lemma 1, we further have:
$F(\S_{1})\leq O(\sqrt{n})\cdot F(\textup{{OPT}}(S))\leq O(\sqrt{n})\cdot\textup{{OPT}}.$ (1)
It remains to analyze the additional flow time incurred during the transition from $\S_{1}$ to $\S_{2}$ due to the insertion of large jobs.
Consider an unproxied large job $j$ inserted at time $t$. This insertion can delay previously scheduled jobs (large or small) by at most $p_{j}$ time units. Let $\delta_{j}$ denote the total increase in flow time due to such delays, which we refer to as the external delay of job $j$. If a large job is proxied, it incurs no external delay, and thus we define $\delta_{j}=0$. In Lemma 6, we prove that only $O(\sqrt{n})$ jobs are delayed in expectation by any unproxied large job, giving:
$\mathbb{E}[\delta_{j}]=O(\sqrt{n})\cdot p_{j}.$ (2)
We must also account for the self-delay of each large job. If $j$ is proxied by a virtual job $j^{\prime}$ released at $r_{j^{\prime}}$, then the actual flow time of $j$ is $C_{j^{\prime}}-r_{j}$. Since $C_{j^{\prime}}-r_{j^{\prime}}$ is already included in $F(\S_{1})$ via the flow time of $j^{\prime}$, the only additional term to consider is $r_{j^{\prime}}-r_{j}$.
On the other hand, if $j$ is not proxied and is directly inserted into the schedule, we only need to consider its current flow time upon insertion. Any subsequent increases to this flow time are charged to the external delays ($\delta_{j^{\prime}}$) caused by later jobs.
Let $\hat{F}_{j}$ denote the self-delay of $j$. We define:
[…]
where $C^{\prime}_{j}$ is the completion time of $j$ in $\S_{2}$ immediately after its insertion (in the last update round of our algorithm). In Lemma 7, we show:
$\sum_{j\in L}\hat{F}_{j}=O(\sqrt{n})\cdot\textup{{OPT}}.$ (3)
In conclusion, assuming the correctness of Lemma 6 (which proves (2)), Lemma 7 (which proves (3)), and using (1), we obtain:
$F(\S_{2})\leq F(\S_{1})+\sum_{j\in L}\delta_{j}+\sum_{j\in L}\hat{F}_{j}\leq O(\sqrt{n})\cdot\textup{{OPT}}+O(\sqrt{n})\cdot\sum_{j\in L}p_{j}.$
Since $\sum_{j\in L}p_{j}\leq\textup{{OPT}}$, it follows that:
$F(\S_{2})=O(\sqrt{n})\cdot\textup{{OPT}}.$
Thus, we conclude that our algorithm is $O(\sqrt{n})$-competitive.
Lemma 6 (External Delay).
For each large job $j$, the additional flow time arising from its delay of other previous jobs is bounded by $O(\sqrt{n})\cdot p_{j}$ in expectation, i.e., $\mathbb{E}(\delta_{j})\leq O(\sqrt{n})\cdot p_{j}$.
Proof.
Let $D_{k}$ be the set of jobs delayed by job $j$ when $w_{j}=k$. We aim to show that $\mathbb{E}[|D_{k}|]=O(\sqrt{n})$. The key claim is that for each job $i\neq j$, it appears in at most one $D_{k}$. This is because there is exactly $p_{j}$ idle time between any two adjacent insertion locations, since we collect $p_{j}$ units of idle time between them. This prevents a single job from being delayed by more than one possible insertion position of $j$.
Therefore, we have:
$\mathbb{E}[|D_{k}|]=\sum_{k=1}^{\sqrt{n}}|D_{k}|\cdot\Pr[w_{j}=k]\leq n\cdot\frac{1}{\lfloor\sqrt{n}\rfloor}\leq O(\sqrt{n}).$
Thus, the expected number of jobs delayed by $j$ is $O(\sqrt{n})$, and since each of them is delayed by at most $p_{j}$, this implies the lemma. ∎
Lemma 7 (Self-Delay).
For all large jobs, the total self-delay is bounded by $O(\sqrt{n})\cdot\textup{{OPT}}$, i.e., $\sum_{j\in L}\hat{F}_{j}=O(\sqrt{n})\cdot\textup{{OPT}}$.
Proof.
For a fixed large job $j$, the self-delay is either $r_{j^{\prime}}-r_{j}$, where $j^{\prime}$ is the proxied virtual job, or $C^{\prime}_{j}-r_{j}$ if $j$ is unproxied. Both types of self-delay arise from only two sources: collecting idle time or passing through busy time.
First, consider the idle time collected. Since $w_{j}\leq\sqrt{n}$, the total idle time accumulated for each large job $j$ is at most $\sqrt{n}\cdot p_{j}$. Summing over all large jobs, we have:
$\sum_{j\in L}\text{(idle contribution to self-delay)}\leq\sqrt{n}\cdot\sum_{j\in L}p_{j}\leq\sqrt{n}\cdot\textup{{OPT}}.$
Next, we analyze the busy time passed through by jobs. This may occur during the intervals $r_{j^{\prime}},r_{j})$ (for proxied jobs) or $r_{j},C^{\prime}_{j})$ (for unproxied jobs). In both cases, the large job remains unproxied throughout the interval. By Lemma 1, the number of unproxied large jobs at any time $t$ is at most $\sqrt{n}$. Since the total length of busy intervals is at most the total processing time of all jobs, we have:
$\sum_{j\in L}\text{(busy-time contribution to self-delay)}\leq\sqrt{n}\cdot\textup{{OPT}}.$
Combining both contributions, the total self-delay across all large jobs is bounded by
$\sum_{j\in L}\hat{F}_{j}=O(\sqrt{n})\cdot\textup{{OPT}}.$
4.2 Generalization to Multi-Machine Case
The analysis for the multi-machine setting directly parallels the single-machine case. The core proof structure — bounding the flow time of small jobs, external delay, and self-delay — remains identical. The generalization simply requires adjusting the key parameters to account for the presence of $m$ machines. Specifically, the threshold for large jobs becomes $\sqrt{nm}$, and we sample an additional random variable, $m_{j}$, for each large job $j$ to denote its assigned machine. These adjustments alter each component of the analysis as follows:
  • Small Job Flow Time: The large-job threshold $\ell$ is now set to $\sqrt{nm}$, which implies that the size of each small job is upper bounded by $\textup{{OPT}}/\sqrt{nm}$. Therefore, the total flow time incurred by small jobs is bounded by $O(\sqrt{n/m})\cdot\textup{{OPT}}$.
  • External Delay: The external delay for any large job $j$ is bounded by $O(\sqrt{n/m})\cdot p_{j}$. In the single-machine setting, each job has $\sqrt{n}$ insertion choices, resulting in an expected delay affecting $\sqrt{n}$ jobs. In the multi-machine setting, we allocate $\sqrt{n/m}$ choices to $w_{j}$ and $m$ choices to $m_{j}$, resulting in an expected delay affecting only $\sqrt{n/m}$ jobs.
  • Self-Delay (Idle Time): The self-delay incurred by collecting idle time is at most $\sqrt{n/m}\cdot p_{j}$, due to the smaller upper bound on $w_{j}$.
  • Self-Delay (Busy Time): The self-delay incurred by passing through busy periods is bounded by $O(\sqrt{n/m})\cdot\textup{{OPT}}$. This is because the number of waiting large jobs increases to $\sqrt{nm}$, while the total length of busy time across all machines decreases to $\textup{{OPT}}/m$.
For completeness, we present the full algorithm and its analysis in Appendix B.
5 Deterministic Non-Preemptive Algorithm
This section introduces a deterministic scheduling algorithm that operates without preemption or restart. In the single-machine setting, the trivial $O(n)$-competitive algorithm is known to be tight. However, when the number of machines $m$ is no longer constant, this $O(n)$ bound becomes loose. Our algorithm achieves a competitive ratio of $O(n/m^{2}+\sqrt{n/m}\log m)$.
The algorithm builds on the online rank-based partitioning framework introduced in Algorithm 1. Upon the arrival of a job $j$, it is classified as either large or small according to the rule specified in Algorithm 1 with $\ell=\lfloor\sqrt{nm}\rfloor$. Recall that this partitioning framework not only distinguishes between large and small jobs but may also release proxy jobs and mark certain jobs as proxied.
Machines are assigned specific roles as follows:
  • Machine 1: Reserved for large jobs.
  • Mixed machines (Machines 22 to $\lceil m/2\rceil+1$): Accept both small and large jobs.
  • Small-Only machines (Machines $\lceil m/2\rceil+2$ to $m$): Restricted to small jobs only.
Note that not all Mixed machines are available to large jobs at all times. Instead, we adopt a fine-grained policy to dynamically determine the number of active Mixed machines dedicated to large jobs. This is generated by a function $\gamma$, which depends on $k$, the number of active large jobs still waiting. Roughly speaking, at any given time, we activate $\gamma(k)$ Mixed machines to serve large jobs, where:
$\gamma(k)=\left\lfloor\frac{k}{\sqrt{n/m}}\right\rfloor.$
Finally, the scheduling policy follows a generalized version of NSJF, respecting the machine roles defined above:
  • If Machine 1 is idle, it schedules the smallest waiting large job, if available.
  • If a Mixed machine is idle, it first schedules the smallest waiting small job, if one is available. Otherwise, let $\theta$ be the number of Mixed machines currently processing large jobs. If $\theta<\gamma(k)$, where $k$ is the number of active large jobs still waiting, the algorithm schedules the smallest such job.
  • If a Small-Only machine is idle, it schedules the smallest waiting small job, if available.
We use $J$ to denote the real job set, which includes only real jobs and excludes proxy jobs. Let $L$ denote the set of large jobs, and let $S$ denote the set of small jobs, including proxy jobs. Recall that, for a proxied large job $j$ (proxied by $j^{\prime}$), its flow time consists of two components: $F_{j}=C_{j}-r_{j}=(C_{j^{\prime}}-r_{j^{\prime}})+(r_{j^{\prime}}-r_{j}).$ The first term $C_{j^{\prime}}-r_{j^{\prime}}$, the flow time of the proxy job $j^{\prime}$, is already included in $\sum_{j\in S}F_{j}$. Therefore, we define the following as a modified flow time:
[…]
We have: $\sum_{j\in L}\hat{F}_{j}+\sum_{j\in S}F_{j}=\sum_{j\in J}F_{j}.$ Let us begin by bounding the total flow time of jobs in $L$.
Lemma 8 (Large Jobs).
$\sum_{j\in L}\hat{F}_{j}\leq O(\sqrt{n/m})\cdot\textup{{OPT}}.$
Proof.
We first bound the waiting time component $r_{j},s_{j})$ (or $r_{j},r_{j^{\prime}})$ if proxied) of the flow time which does not include the $p_{j}$ part. At any time $t$ in this interval, $j$ is not scheduled because all active Mixed machines are busy. Let $\theta(t)$ be the number of Mixed machines currently processing jobs, and let $k(t)$ be the number of active large jobs waiting at time $t$. By the scheduling rule, we either have
$\gamma(k(t))\leq\theta(t)+1,\quad\text{or}\quad\theta(t)=\lceil m/2\rceil.$
In the second case, since $k(t)$ is at most $\lfloor\sqrt{nm}\rfloor$, and using the definition $\gamma(k)=\left\lfloor\frac{k}{\sqrt{n/m}}\right\rfloor$, we get:
$k(t)\leq O(\sqrt{n/m})\cdot(\theta(t)+1).$
Also, note that while a large job is waiting, the large-only machine (Machine 1) must be busy. Letting $\mathbb{I}_{k(t)\geq 1}(t)$ denote the indicator function for whether $k(t)\geq 1$, we have:
$\int_{0}^{\infty}(\theta(t)+1)\cdot\mathbb{I}_{k(t)\geq 1}(t)\,dt\leq\sum_{j\in J}p_{j}.$
Combining all these elements and reincorporating $p_{j}$, we have:
$\sum_{j\in L}\hat{F}_{j}$ $\leq\sum_{j\in L}p_{j}+\int_{0}^{\infty}k(t)\,dt$
$\leq\sum_{j\in L}p_{j}+O(\sqrt{n/m})\cdot\int_{0}^{\infty}(\theta(t)+1)\cdot\mathbb{I}_{k(t)\geq 0}(t)\,dt$
$\leq\sum_{j\in L}p_{j}+O(\sqrt{n/m})\cdot\sum_{j\in J}p_{j}$
$=O\left(\sqrt{\frac{n}{m}}\right)\cdot\textup{{OPT}}.$
Next, we analyze the contribution of small jobs. The first loss occurs because of the large-only machine. Ideally, even without any large job, our algorithm is a (m1)(m-1)-machine NSJF algorithm, which is competitive to the (m1)(m-1)-machine preemptive optimal solution. However, the (m1)(m-1)-machine preemptive optimal may already be larger than the $m$-machine optimal solution. To address it, we give the following lemma.
Lemma 9.
For every $0\leq k\leq m-1$, letting $\textup{{OPT}}_{m-k}$ be the optimal preemptive solution only using mkm-k machines, we have $\textup{{OPT}}_{m-k}=O\left(\frac{nk}{m(m-k)}\right)\cdot\textup{{OPT}}_{m}$.
Before proceeding, we recall that the preemptive optimal solutions discussed here, $\textup{{OPT}}_{m}$ and $\textup{{OPT}}_{m-k}$, operate under preemptive model without migration.
Proof.
Consider a preemptive optimal solution with $m$ machines. We will use it to construct a feasible solution using mkm-k machines and prove that its total flow time increases by at most $\frac{nk}{m(m-k)}\cdot\textup{{OPT}}_{m}$. This will complete the proof of the lemma.
First, we sort the machines in the optimal solution $\textup{{OPT}}_{m}$ by their total workloads. Let $J_{i}$ denote the set of jobs scheduled on machine $i$ in $\textup{{OPT}}_{m}$, and let $J_{[i]}=\bigcup_{h=1}^{i}J_{h}$ denote the union of jobs on the first $i$ machines. Machines are sorted so that
$\sum_{j\in J_{1}}p_{j}\leq\sum_{j\in J_{2}}p_{j}\leq\cdots\leq\sum_{j\in J_{m}}p_{j}.$
We consider the first $k$ machines and reassign all jobs in $J_{[k]}$ to the remaining mkm-k machines in arbitrary order. By construction, we have:
$\sum_{j\in J_{[k]}}p_{j}\leq\frac{k\sum_{j\in J}p_{j}}{m}.$
The reassignment proceeds as follows. For each job $j\in J_{[k]}$, we preserve its starting time $t$ from $\textup{{OPT}}_{m}$, and insert it at the same time $t$ on one of the machines among machines mk+1m-k+1 to $m$. As a result, the existing jobs on that machine are delayed by at most $p_{j}$.
To minimize the impact, we always place $j$ on the machine with the fewest currently assigned jobs. Since there are mkm-k machines and at most $n$ jobs, this ensures the number of jobs on that machine is at most n/(mk)n/(m-k). Therefore, the additional cost due to delaying existing jobs is at most $np_{j}/(m-k)$.
On the other hand, since we insert job $j$ at its original start time, its flow time is at most what it was in $\textup{{OPT}}_{m}$. Even though future delays may increase $j$’s flow time, these increases are charged to the delay costs of subsequent jobs. Thus, the increase in total flow time due to inserting a single job $j$ compared to $\textup{{OPT}}_{m}$ is at most
$\frac{np_{j}}{m-k}.$
Summing over all $j\in J_{[k]}$, the total increase in flow time is at most
$\frac{n\sum_{j\in J_{[k]}}p_{j}}{m-k}\leq\frac{nk\sum_{j\in J}p_{j}}{m(m-k)}\leq\frac{nk\cdot\textup{{OPT}}_{m}}{m(m-k)}.$
In conclusion, we obtain the desired bound:
$\textup{{OPT}}_{m-k}\leq\left(1+\frac{nk}{m(m-k)}\right)\cdot\textup{{OPT}}_{m}.$
As in Lemma 12, we illustrate the ideas again here for completeness. We partition the timeline into intervals based on when large jobs are started. Let $t_{h}$ denote the start time of the $h$-th large job, with $t_{0}=0$. Let $S_{h}$ be the set of small jobs released during the interval $t_{h-1},t_{h})$.
Let $\textup{{OPT}}_{m-1}(S_{h})$ denote the optimal total flow time for scheduling $S_{h}$ on m1m-1 machines. Since the intervals are disjoint, we have:
$\textup{{OPT}}_{m-1}\geq\sum_{h}F(\textup{{OPT}}_{m-1}(S_{h})).$
We compare our algorithm’s flow time on $S_{h}$ to $\textup{{OPT}}_{m-1}(S_{h})$. During $t_{h-1},t_{h})$, no new large jobs are started, and all jobs in $S_{h}$ arrive after $t_{h-1}$. The only difference from running $\textup{{NSJF}}(S_{h})$ on a fresh instance is that each machine may still be processing a job that started before $t_{h-1}$. We refer to these as initial blocking jobs, and define the corresponding initial blocking period on machine $i$ as $t_{h-1},t_{h-1}+b_{i})$, where $b_{i}$, the initial blocking time is the remaining processing time at $t_{h-1}$. Thus, our algorithm behaves like NSJF on m1m-1 machines with such initial blocking times periods.
This behavior is captured by Lemma 29. To apply it, we define a shifted instance $S^{\prime}_{h}$ by translating the time origin to $t_{h-1}$, i.e., replacing each release time $r_{j}$ with $r_{j}-t_{h-1}$. Then, the flow time of each job $j\in S_{h}$ under our algorithm matches its flow time in $\textup{{NSJF}}(S^{\prime}_{h},\vec{b})$, where each machine has initial blocking interval $0,b_{i})$.
Lemma 10.
For each small-job period $(t_{h-1},t_{h}]$, the total length of initial blocking time satisfies:
$\sum_{i=1}^{m-1}b_{i}\leq O\left(\sqrt{\frac{m}{n}}\log m\right)\cdot\textup{{OPT}}.$
Proof.
We split the analysis based on whether the blocking job is small or large.
We first consider the blocking caused by small jobs. By Lemma 1, each small job has size at most $O\left(\frac{\textup{{OPT}}}{\sqrt{nm}}\right)$. Since at most $m$ machines can be blocked, the total blocking time from small jobs is:
$O\left(\frac{\textup{{OPT}}}{\sqrt{nm}}\right)\cdot m=O\left(\sqrt{\frac{m}{n}}\cdot\textup{{OPT}}\right).$
For the blocking caused by large jobs, we proceed as follows. For large jobs causing blocking, let $p_{i}$ be the size of the $i$-th large job that causes blocking. When it starts, at least i1i-1 Mixed machines are already processing large jobs, so it must satisfy $\gamma(k)\geq i$ for the number of active large jobs still waiting $k$. Using $\gamma(k)=\lfloor k/(\sqrt{n/m})\rfloor$, we get $k\geq i\sqrt{n/m}$, and thus:
$p_{i}\leq\frac{\textup{{OPT}}}{k}\leq\frac{\textup{{OPT}}}{i\sqrt{n/m}}\leq\frac{1}{i}\cdot O\left(\sqrt{\frac{m}{n}}\cdot\textup{{OPT}}\right).$
Summing over $i=1$ to $m$, the total blocking from large jobs is:
$\sum_{i=1}^{m}p_{i}\leq O\left(\sqrt{\frac{m}{n}}\cdot\textup{{OPT}}\right)\cdot\sum_{i=1}^{m}\frac{1}{i}=O\left(\sqrt{\frac{m}{n}}\log m\cdot\textup{{OPT}}\right).$
Adding both parts completes the proof. ∎
With the bound on the initial blocking times $b_{i}$, we can apply the generalized NSJF analysis in Lemma 29. This lemma allows us to compare the performance of NSJF with initial blocking times against the optimal schedule without any blocking. We now prove the main result.
Theorem 2.
There exists an online polynomial time deterministic non-preemptive algorithm for online total flow time minimization with competitive ratio $O(n/m^{2}+\sqrt{n/m}\log m)$, against the preemptive offline optimal solution.
Proof.
By Lemma 8, the total modified flow time of large jobs is bounded by:
$\sum_{j\in L}\hat{F}_{j}=O\left(\sqrt{\frac{n}{m}}\right)\cdot\textup{{OPT}}.$
Let $\tau$ be the size upper bound for small jobs, i.e., $\tau=O(\textup{{OPT}}/\sqrt{nm})$. For each small-job segment $t_{h-1},t_{h})$, our algorithm performs the same as $\textup{{NSJF}}(S^{\prime}_{h})$ with initial blocking vector $\vec{b}$, denoted by $F_{j}(\textup{{NSJF}}(S^{\prime}_{h},\vec{b}))$. By Lemma 29, and letting $B=\sum_{i}b_{i}$, we have:
$\sum_{j\in S^{\prime}_{h}}F_{j}(\textup{{NSJF}}(S^{\prime}_{h},\vec{b}))\leq\textup{{OPT}}_{m-1}(S_{h})+2|S_{h}|\tau+\frac{|S_{h}|B}{m}.$
Summing over all disjoint small-job segments:
$\sum_{j\in S}F_{j}\leq F(\textup{{OPT}}_{m-1})+2n\tau+\frac{nB}{m}.$
Using the bounds $\tau=O(\textup{{OPT}}/\sqrt{nm})$ and $B=O(\sqrt{m/n}\log m\cdot\textup{{OPT}})$, we get:
$\sum_{j\in S}F_{j}\leq\textup{{OPT}}_{m-1}+O\left(\sqrt{\frac{n}{m}}\right)\cdot\textup{{OPT}}+O\left(\sqrt{\frac{n}{m}}\log m\right)\cdot\textup{{OPT}}.$
By Lemma 9, which bounds the loss in flow time from using m1m-1 instead of $m$ machines, we have:
$\textup{{OPT}}_{m-1}\leq\textup{{OPT}}+O(n/m^{2})\cdot\textup{{OPT}}.$
Combining all terms:
$\sum_{j\in S}F_{j}$ $\leq\textup{{OPT}}+O(n/m^{2})\cdot\textup{{OPT}}+O\left(\sqrt{\frac{n}{m}}\right)\cdot\textup{{OPT}}+O\left(\sqrt{\frac{n}{m}}\log m\right)\cdot\textup{{OPT}}$
$=O\left(\frac{n}{m^{2}}+\sqrt{\frac{n}{m}}\log m\right)\cdot\textup{{OPT}}.$
Finally, adding the contributions from large and small jobs:
$\sum_{j\in J}F_{j}=\sum_{j\in L}\hat{F}_{j}+\sum_{j\in S}F_{j}=O\left(\frac{n}{m^{2}}+\sqrt{\frac{n}{m}}\log m\right)\cdot\textup{{OPT}}.$
6 Deterministic Algorithm with Kill-and-Restart
To utilize the power of kill-and-restart without relying on randomization, we first introduce a fine-grained online partitioning framework with the following modification (see Algorithm 3 for details, and note that the threshold will be set to $\sqrt{nm}$):
  • Modified partition rule: Let $P$ be the total size of all previously released jobs (excluding proxy jobs), which serves as a lower bound on OPT. We classify a new job $j$ as a small job if $p_{j}\leq 4P/\ell$, rather than using the large job queue as in the original rule.
The modification is motivated by the need for a size-based lower bound in the restart setting, as opposed to a purely rank-based one. For instance, we aim to avoid killing jobs that are too small. These changes are mild and preserve all the original properties of the online partitioning framework.
Data: Active large job set AA, and every job’s other current states
Function Partition($i,\ell$):
 // New Rule: An absolute smallness rule to filter out tiny jobs.
   Let $P$ be the sum of processing times of all (original) jobs released so far;
 if $p_{i}\leq 4P/\ell$ then
    return $i$ is small
 else if $|A|<\ell$ then
      Add $i$ to AA;
    return $i$ is large
 else if $p_{i}>\min_{j\in A}p_{j}$ then
      Let $j=\arg\min_{j\in A}p_{j}$;
      Remove $j$ from AA and add $i$ to AA;
    if job $j$ is still waiting then
         Create a new proxy job $j^{\prime}$ with $r_{j^{\prime}}=r_{i}$ and $p_{j^{\prime}}=p_{j}$;
         Mark $j$ as proxied by $j^{\prime}$;
       // The scheduler will actually process $j$ when it processes $j^{\prime}$.
    else
       // $j$ is a committed large job.
      end if
    return $i$ is large
 else
    return $i$ is small
   end if
end
Algorithm 3 Refined Online Partitioning for Kill-and-Restart Models
Lemma 11.
Under the online rank-based partitioning framework (fine-grained for restart), the following properties hold:
  1. 1.
    The number of large jobs is at most $\ell$ at any time.
  2. 2.
    Let $S$ be the set of small jobs, including proxy jobs. Then $S$ is strictly weaker than the full job set $J$, meaning $F(\textup{{OPT}}(S))\leq F(\textup{{OPT}}(J))$.
  3. 3.
    The size of any small job, including any retired large job, is upper bounded by $O(\textup{{OPT}}/\ell)$.
  4. 4.
    New size-based lower bound: The size of any large job is at least $4P(r_{j})/\ell$, where $P(r_{j})$ denotes the total processing time of all jobs released up to $r_{j}$, excluding proxy jobs.
Proof.
The only property that requires verification under the modified framework is the size bound for small jobs. By the new partition rule, we classify a job $j$ as small if $p_{j}\leq 4P(r_{j})/\ell$. Since $P(r_{j})$ is a valid lower bound on OPT, it follows that $p_{j}\leq O\left(\textup{{OPT}}/\ell\right)$, as required. The remaining properties are preserved from the original online partitioning framework. ∎
After introducing the fine-grained partitioning framework, we now describe the basic design of our scheduling policy, which divides the $m$ machines into two roles:
  • Small-Only machines (Machines 11 to $\lfloor m/2\rfloor$): These machines accept only small jobs.
  • Mixed machines (Machines $\lfloor m/2\rfloor+1$ to $m$): These machines can process both small and large jobs.
Our scheduling policy is a generalized version of the NSJF algorithm, incorporating the specific machine assignment rule described above. In particular, the algorithm makes scheduling decisions based on the partitioning produced by Algorithm 1. Each job is classified as either large or small, and further as retired or active, and proxied or unproxied.
The scheduling rule proceeds as follows: the algorithm always selects the smallest unproxied waiting job to schedule. Additionally, Small-Only machines are restricted to processing only small jobs.
The kill-and-restart rule applies only to active large jobs. Intuitively, an active large job should be killed only if it is blocking too many small jobs. We say that an active large job blocks a small job if it is processing at the small job’s release time. To detect such situations, the algorithm monitors the number of blocked small jobs. Specifically, if at any point the number of blocked small jobs reaches the threshold $\sqrt{nm}$, we kill all currently processing active large jobs. At the moment of a kill event, as well as when a new large job starts processing, we reset the counter tracking the number of blocked small jobs.
With this, we complete the high-level description of our algorithmic design. The full pseudocode is presented in Algorithm 4.
Include : Algorithm 3
Data: Potential counter $\phi$ initiated as 0; every job’s current states as determined by Algorithm 3.
On job $i$ release:
[…]
 if Job $i$ is classified as small then
[…]
    $\phi\leftarrow\phi+1$ if $i$ is blocked ;
    if $\text{$\phi$}=\lfloor\sqrt{nm}\rfloor$ then
         Kill all processing active large jobs and $\phi\leftarrow 0$;
      end if
   end if
On any job completion or after any job release:
 for each idle Small-Only machine $i$ do
    if there exists a waiting small job then
       $j\leftarrow$ the smallest small job ;
         Schedule $j$ on $i$;
      end if
   end for
 for each idle Mixed machine $i$ do
    if there exists a waiting job that is not proxied then
       $j\leftarrow$ the smallest waiting unproxied job ;
         Schedule $j$ on $i$ ;
       $\phi\leftarrow 0$ if $j$ is large ;
      end if
   end for
Algorithm 4 Deterministic Algorithm with Kill-and-Restart on $m$ Machines
Moving into the analysis, we now distinguish between several types of job sets: The real job set $J$, which includes only actual jobs and excludes proxy jobs. The unblocked small job set $S$, and the blocked small job set $B$, both of which may include proxy jobs. The large job set $L$, which includes all large jobs, even if they become retired or are proxied at some point.
Note that we do not have $B\cup S\cup L\subseteq J$ due to the presence of proxy jobs. We analyze the total flow time contributed by $S$, $B$, and $L$ separately.
Lemma 12 (Unblocked Small Job).
Let $S$ be the set of unblocked small jobs (including proxy jobs). Then:
$\sum_{j\in S}F_{j}=O\left(\sqrt{\frac{n}{m}}\right)\cdot\textup{{OPT}}.$
Proof.
In the absence of blocked small jobs and retired large jobs, the flow time of all jobs in $S$ would be bounded by $F(\textup{{NSJF}}(S))$, which is at most $O\left(\sqrt{n/m}\right)\cdot F(\textup{{OPT}}(S))$ by Lemma 2. The reason is that the processing of active large jobs never affects the scheduling of unblocked small jobs. Whenever NSJF wants to schedule an unblocked small job $j$, the algorithm cannot have a large job smaller than $j$ outside of $S$. If such a large job is active, it must be larger than all small jobs by our online partition rule. If it is retired and still waiting, it is proxied and, therefore, never considered for scheduling.
However, the presence of blocked small jobs and retired large jobs may delay the scheduling of unblocked small jobs, because we mix them together. We define a set $U$, consisting of the retired large jobs, and recall the set $B$, consisting of the blocked small jobs. Correspondingly, we construct two modified instances $U^{\prime}$ and $B^{\prime}$ by updating the release times of all jobs $j\in U$ and $j\in B$ to $s_{j}$; note that our algorithm’s schedule on $S\cup U\cup B$ is identical to that of $\textup{{NSJF}}(S\cup U^{\prime}\cup B^{\prime})$. By Lemma 2, we know that
$F(\textup{{NSJF}}(S\cup U^{\prime}\cup B^{\prime}))\leq F(\textup{{OPT}}(S\cup U^{\prime}\cup B^{\prime}))+O(n\tau)\leq F(\textup{{OPT}}(S\cup U^{\prime}\cup B^{\prime}))+O(\sqrt{n/m})\cdot\textup{{OPT}}.$
This alone does not prove the lemma, because the jobs in $U^{\prime}$ and $B^{\prime}$ may have later release times than in the original instance, so $F(\textup{{OPT}}(S\cup U^{\prime}\cup B^{\prime}))$ is not necessarily upper bounded by OPT. However, for every retired job $j\in U^{\prime}$, by the same reason as Lemma 1, it is retired after $s_{j}$, and we can find a newly released job $j^{\prime}$, where $p_{j^{\prime}}>p_{j}$, and $r_{j^{\prime}}>s_{j}$, where $s_{j}$ is the release time of $j$ in $U^{\prime}$. So, at least $S\cup U^{\prime}$ is a weaker set of $J$, and $F(\textup{{OPT}}(S\cup U^{\prime}))\leq\textup{{OPT}}$.
We now aim to bound the difference between $F(\textup{{OPT}}(S\cup U^{\prime}\cup B^{\prime}))$ and $F(\textup{{OPT}}(S\cup U^{\prime}))$. To do so, we partition the timeline into intervals free of unproxied large jobs, denoted as $t_{h},t_{h+1})$ for $h=1,2,\dots$, where every $t_{h}$ can be viewed as the time we start a currently unproxied large job (it may be killed or retired later). In each interval $t_{h},t_{h+1})$, no unproxied large job is processed. Unblocked small jobs are only released in such intervals.
Let $S_{h}$ be the set of unblocked small jobs with $r_{j}\in[t_{h},t_{h+1})$ and retired large jobs who becomes small in $t_{h},t_{h+1})$. Clearly, $\bigcup_{h}S_{h}=S\cup U^{\prime}$ by definition. We similarly define $B^{\prime}_{h}\subseteq B^{\prime}$ to be the jobs in $B^{\prime}$ whose processing falls within $t_{h},t_{h+1})$. Observe that, because only $B^{\prime}_{h}$ can affect $S_{h}$, we have
$\sum_{j\in S_{h}}F_{j}(\textup{{NSJF}}(S\cup U^{\prime}\cup B^{\prime}))=\sum_{j\in S_{h}}F_{j}(\textup{{NSJF}}(S_{h}\cup B^{\prime}_{h}))\leq F(\textup{{NSJF}}(S_{h}\cup B^{\prime}_{h})).$
An important property of our algorithm is that $|B^{\prime}_{h}|\leq\sqrt{nm}$; otherwise, we would have killed all unproxied jobs before time $t_{h}$. We will use this property to bound $F(\textup{{NSJF}}(S_{h}\cup B^{\prime}_{h})$.
Now consider an optimal preemptive solution $\textup{{OPT}}(S_{h})$. We insert the jobs in $B^{\prime}_{h}$ into $\textup{{OPT}}(S_{h})$ one by one, in a migratory fashion: for each job $j\in B^{\prime}_{h}$, we scan from $r_{j}$ and process it at the first available idle slot, without delaying the original schedule. If $j$’s processing overlaps a busy period, we interrupt it and resume at the next idle slot, possibly on a different machine.
Let $F^{\prime}_{j}$ be the flow time of such a job $j$. It consists of $p_{j}$ plus some waiting time. This waiting time falls entirely in busy periods and is thus upper bounded by $\sum_{i\in S_{h}\cup B^{\prime}_{h}}p_{i}/m$. Since there are at most $\sqrt{nm}$ jobs in $B^{\prime}_{h}$ (by our restart rule), we get:
$F(\textup{{OPT}}^{m}(S_{h}\cup B^{\prime}_{h}))\leq F(\textup{{OPT}}(S_{h}))+O(\sqrt{n/m})\sum_{j\in S_{h}\cup B^{\prime}_{h}}p_{j}.$
Here, we use $\textup{{OPT}}^{m}$ to denote the optimal preemptive solution with migration. Note that migration is required in the above construction, and NSJF is compared against $\textup{{OPT}}^{m}$ in Lemma 2. Summing over all $h$, we obtain:
$\sum_{j\in S}F_{j}(\textup{{NSJF}}(S\cup U^{\prime}\cup B^{\prime}))$ $\leq\sum_{h}F(\textup{{NSJF}}(S_{h}\cup B^{\prime}_{h}))$
$\leq\sum_{h}F(\textup{{OPT}}^{m}(S_{h}\cup B^{\prime}_{h}))+O\left(\frac{|S_{h}|+|B^{\prime}_{h}|}{\sqrt{nm}}\right)\cdot\textup{{OPT}}$
$\leq\sum_{h}F(\textup{{OPT}}(S_{h}))+O\left(\frac{|S_{h}|+|B^{\prime}_{h}|}{\sqrt{nm}}\right)\cdot\textup{{OPT}}+O(\sqrt{n/m}\sum_{j\in S_{h}\cup B^{\prime}_{h}})p_{j}$
$\leq F(\textup{{OPT}}(S))+O(\sqrt{n/m})\cdot\textup{{OPT}}+O(\sqrt{n/m})\sum_{j\in S\cup B^{\prime}}p_{j}$
$\leq O(\sqrt{n/m})\cdot\textup{{OPT}}.$
The last inequality holds because all jobs in $S\cup U^{\prime}$ can be charged to a corresponding larger job in the original job set $J$. Notably, a proxy job appears only when a larger job arrives, and thus, we can charge it to that larger job. Finally, since our algorithm schedules $S\cup U\cup B$ exactly the same as $\textup{{NSJF}}(S\cup U^{\prime}\cup B^{\prime})$, the lemma follows. ∎
Lemma 13 (Blocked Small Job).
Let $B$ be the set of blocked small jobs. Then the total flow time of jobs in $B$ is bounded by:
$\sum_{j\in B}F_{j}\leq O\left(\sqrt{\frac{n}{m}}\right)\cdot\textup{{OPT}}.$
Proof.
Consider an arbitrary blocked small job $j$. By our scheduling rule, $j$ can only experience waiting time when all Small-Only machines are busy, since $j$ is classified as a small job. Therefore, during the entire interval $r_{j},s_{j})$, all Small-Only machines must be fully occupied. We refer to this as a busy time of Small-Only machines.
On the other hand, due to the restart rule, at any busy time $t$ of Small-Only machines, the number of blocked small jobs still waiting is at most $\sqrt{nm}$. This is because the number of blocked small jobs increases only when a new large job starts processing. However, when that happens, all previously blocked small jobs are scheduled and no longer waiting.
Since there are $\lfloor m/2\rfloor$ Small-Only machines and they are fully occupied during busy periods, the total length of all busy intervals is at most
$\frac{2\cdot\textup{{OPT}}}{m-1}.$
Therefore, if $m\geq 2$, the total waiting time of all blocked small jobs is bounded by
$O\left(\sqrt{\frac{n}{m}}\right)\cdot\textup{{OPT}}.$
We remark that when $m=1$, there is only one Small-Only machine, and this argument no longer holds. Finally, the total flow time of all blocked small jobs is:
$\sum_{j\in B}F_{j}=\sum_{j\in B}(s_{j}-r_{j}+p_{j})\leq O\left(\sqrt{\frac{n}{m}}\right)\cdot\textup{{OPT}}.$
For each large job $j$, consider its completion time $C_{j}$, which is either its own completion time or the completion time of its proxy job $j^{\prime}$. If it is proxied, the delay from $r_{j^{\prime}}$ to $C_{j^{\prime}}=C_{j}$ is already charged to the small job flow time in Lemma 12 and Lemma 13. Thus, we only need to analyze the modified flow time where
[…]
Lemma 14 (Large Job).
For large job set $L$. We have $\sum_{j\in L}\hat{F}_{j}\leq O(\sqrt{n/m})\cdot\textup{{OPT}}$.
Proof.
Ignoring the final $p_{j}$ units of processing time, we focus on the delay interval $r_{j},s_{j})$ (or $r_{j},r_{j^{\prime}})$ if $j$ is proxied by $j^{\prime}$). Define $t_{j}=s_{j}$ (or $t_{j}=r_{j^{\prime}}$ if proxied), and aim to bound $t_{j}-r_{j}$.
We begin by showing that shifting the release time $r_{j}$ by $\sqrt{n/m}\cdot p_{j}$ is acceptable. Let $t^{\prime}_{j}=r_{j}+\sqrt{n/m}\cdot p_{j}$. Then:
$\sum_{j\in L}(t^{\prime}_{j}-r_{j})=\sqrt{\frac{n}{m}}\cdot\sum_{j\in L}p_{j}\leq\sqrt{\frac{n}{m}}\cdot\textup{{OPT}}.$ (4)
We now analyze the remaining delay interval $t^{\prime}_{j},t_{j})$. For each time $t\in[t^{\prime}_{j},t_{j})$, we classify it into one of the following two categories:
  1. 1.
    Genuine busy time: All Mixed machines are busy at $t$ originally, and at least $\lceil m/4\rceil$ of the running jobs are not restarted later.
  2. 2.
    Fake busy time: All Mixed machines are busy at $t$ originally, but more than $\lceil m/4\rceil$ of the running jobs are restarted.
We decompose the delay $t_{j}-t^{\prime}_{j}$ into two parts: $F_{j}^{(1)}$ and $F_{j}^{(2)}$, corresponding to time spent in genuine busy and fake busy intervals, respectively:
$t_{j}-t^{\prime}_{j}=F_{j}^{(1)}+F_{j}^{(2)}.$
Bounding $F_{j}^{(1)}$: At any genuine busy time, at least $\lceil m/4\rceil$ machines are processing jobs that are not restarted. The total duration of such intervals is at most $O(\textup{{OPT}}/m)$. Furthermore, each such interval can contribute to the delay of at most $\sqrt{nm}$ different large jobs (by Lemma 11). Hence:
$\sum_{j\in L}F_{j}^{(1)}\leq\sqrt{nm}\cdot O\left(\frac{\textup{{OPT}}}{m}\right)=O\left(\sqrt{\frac{n}{m}}\cdot\textup{{OPT}}\right).$ (5)
Bounding $F_{j}^{(2)}$: Next, we analyze the contribution of fake busy intervals, which are induced by kill events. Let the $i$-th kill event occur at time $\beta_{i}$. For a time $t$ to contribute to $F_{j}^{(2)}$, the following conditions must hold:
  1. 1.
    $t>\beta_{i-1}$ and $t>t^{\prime}_{j}$,
  2. 2.
    $t$ is not a genuine busy time.
We now prove that such $t$ satisfies:
$\beta_{i}-t\leq p_{j}.$ (6)
Proof of the claim: Since $t$ is a fake busy time, more than $\lceil m/4\rceil$ jobs are restarted at $\beta_{i}$. Suppose, for contradiction, that $\beta_{i}-t>p_{j}$. Then these jobs must have been running continuously since $t$, implying that their processing times are at least $p_{j}$. By our algorithm, it is impossible that job $j$ is already released when we decide to start one of these larger jobs. Therefore, letting $U$ denote this set of jobs, we have:
$\beta_{i}<r_{j}+\min_{u\in U}p_{u}.$
However, by definition $t^{\prime}_{j}=r_{j}+\sqrt{n/m}\cdot p_{j}$, and assuming the total workload released up to $r_{j}$ is $P(r_{j})$, we have:
$t^{\prime}_{j}\geq r_{j}+\frac{4P(r_{j})}{m}\geq r_{j}+\frac{4\sum_{u\in U}p_{u}}{m}\geq r_{j}+\min_{u\in U}p_{u}>\beta_{i}.$
This proves (6), by $\beta_{i}-t\leq\beta_{i}-t^{\prime}_{j}<0$.
Thus, for each kill event, the total contribution to $\sum_{j}F_{j}^{(2)}$ is at most:
$\sum_{j\in L}p_{j}\leq\textup{{OPT}}.$
Finally, we bound the number of kill events. Since each kill is triggered only after at least $\lfloor\sqrt{nm}\rfloor$ new small jobs are released, becoming blocked, the total number of restart events is at most $O(\sqrt{n/m})$. Therefore:
$\sum_{j\in L}F_{j}^{(2)}\leq\sqrt{\frac{n}{m}}\cdot\textup{{OPT}}.$ (7)
Combining (4), (5), and (7), we obtain:
$\sum_{j\in L}\hat{F_{j}}\leq\sum_{j\in L}\left(p_{j}+t_{j}-t^{\prime}_{j}+t^{\prime}_{j}-r_{j}\right)\leq O\left(\sqrt{\frac{n}{m}}\right)\cdot\textup{{OPT}}.$
Theorem 3.
There exists an online polynomial-time deterministic algorithm with kill-and-restart for total flow time minimization, e.g., Algorithm 4, that is $O(\sqrt{n/m})$-competitive against the preemptive offline solution, provided that $m\geq 2$.
Proof.
By Lemma 12, Lemma 13, and Lemma 14.
$\sum_{j\in J}F_{j}\leq\sum_{j\in S}F_{j}+\sum_{j\in B}F_{j}+\sum_{j\in L}\hat{F}_{j}\leq O\left(\sqrt{\frac{n}{m}}\right)\cdot\textup{{OPT}}.$
7 Deterministic Algorithm with Kill-and-Restart without the Knowledge of $n$
In this section, we extend the $\sqrt{n/m}$-competitive algorithm (which assumes the prior knowledge of $n$) to an $O(n^{\alpha}/\sqrt{m})$-competitive algorithm without such knowledge, where $\alpha=(\sqrt{5}-1)/2$. The overall framework still follows Section 6: we keep the same machine partition (Small-Only machines 11 to $\lfloor m/2\rfloor$ and Mixed machines $\lfloor m/2\rfloor+1$ to $m$), and we still kill a large job once enough small jobs are blocked.
The key difference is that the total number of jobs $n$ is unknown in advance. At any time $t$, we only observe the number of jobs released so far, denoted by $n(t)$, rather than the final value of $n$. Consequently, the rank-based partition can no longer use a fixed threshold; instead, it must classify jobs as large or small using a dynamic threshold $\ell(t)$ that depends on $n(t)$. This introduces a major difficulty: the partition is no longer stable as in the known-$n$ setting.
Previously, when $n$ was known, the threshold $\ell$ was fixed. A small job would remain small forever, and the only source of instability was when a large job retired, which we handled by introducing a proxy job. In contrast, in the unknown-$n$ setting, $\ell(t)$ increases monotonically as $n(t)$ grows, and thus varies over time.
As a result, a job may switch between being large and small multiple times. For instance, when $n(t)$ is small, we may treat the top 33 jobs as large, so a job of rank 44 is small. Later, as more jobs arrive and $\ell(t)$ becomes 44, this job becomes large. If even more jobs arrive, $\ell(t)$ may increase to 55, but that same job may drop out of the top-55 and become small again. This uncontrolled oscillation between the large and small states makes the partition unstable and invalidates the previous analysis.
To this end, we make two main modifications as follow in the new algorithm as follows, and the formal description is presented in Algorithm 5.
More Dynamic Partition Rule.
We use a partition rule that depends on the number of jobs released so far. Instead of using $n(t)$ directly, we let $N=2^{k}$ be the smallest power of two such that $N\geq n(t)$, and use NN to define the large–small threshold $\ell$ at every time $t$, where $\ell=\lfloor N^{\alpha}\sqrt{m}\rfloor$. Let $T_{j}(t)$ denote the type of job $j$ at time $t$. We classify $T_{j}(t)$ as large if job $j$ is among the top $\ell$ jobs in size and $p_{j}\geq\frac{4P(t)}{\ell}$, where $P(t)$ is the total size of all jobs released by time $t$. Otherwise, $T_{j}(t)$ is classified as small.
A New Kill Event When NN Is Updated.
Previously, we triggered a kill event only when too many small jobs were waiting while a large job was being processed on a Mixed machine. However, in this setting, a large job may be processed on a Small-Only machine, since it might have been classified as small earlier. To prevent this from causing inefficiency, we introduce a new kill event that occurs whenever NN is updated (that is, when $n(t)>N$, and we double NN to $2N$). At this moment, we check whether any currently large jobs are running on the Small-Only machines and kill all of them if so. We remark that this is one reason for using a rounded value of $n(t)$ (the nearest power of two) as the benchmark, since it allows us to bound the number of such kill events.
Data: Estimate of $n$: NN initiated as 11;
On job $j$ release at $t$:
 if the number of jobs exceeds NN then
    $N\leftarrow 2N$;
      Mark NN as updated ;
   end if
 // Update the Partition of all jobs again based on the current number of jobs NN and the current sum of processing time $P$.
 $\ell=\lfloor N^{\alpha}\sqrt{m}\rfloor$ where $\alpha=(\sqrt{5}-1)/2$ ;
[…]
 if NN is updated by the release of job $j$ then
      Kill all currently large jobs processing on Small-Only machines ;
   end if
 if there exists a processing job which is currently large then
    $\phi\leftarrow$ the number of waiting jobs that are currently small ;
    if $\phi>\ell$ then
         Kill all processing jobs that are currently large.
      end if
   end if
On any job completion or after any job release at $t$:
 for each idle Small-Only machine $i$ do
    if there exists a waiting currently small job then
       $j\leftarrow$ the smallest currently small job ;
         Schedule $j$ on $i$;
      end if
   end for
 for each idle Mixed machine $i$ do
    if there exists a waiting job then
       $j\leftarrow$ the smallest waiting job ;
         Schedule $j$ on $i$ ;
      end if
   end for
Algorithm 5 Deterministic Algorithm with Kill-and-Restart on $m$ Machines, without the knowledge of $n$
Next, we move on to the analysis. For each job $j$, let $T_{j}(t)\in\{\text{Large},\text{Small}\}$ denote the state of job $j$ at time $t$. Let $C_{n}$ be the completion time of the last completed job, which marks the end of the instance. Define $\theta_{j}$ as the earliest time such that $T_{j}(t)=\text{Small}$ for all $t\in[\theta_{j},C_{n})$. The interval $\theta_{j},C_{n})$ is called the last small period of job $j$.
We divide the flow time of each job $j$ into waiting time and its processing time, i.e., $r_{j},s_{j})\cup[s_{j},C_{j})$. Note that to prove the total flow time is in $O(n^{\alpha}/\sqrt{m})\cdot\textup{{OPT}}$, we only need to discuss the total waiting time, since the total processing time is bounded by OPT. Furthermore, we divide the waiting time of each job into two parts with respect to the last small periods as follows.
  • Large waiting time ($F_{j}^{(1)}$): the waiting time before the last small period, i.e.,
    $F_{j}^{(1)}=|[r_{j},s_{j})\cap[r_{j},\theta_{j})|.$
  • Small waiting time ($F_{j}^{(2)}$): the waiting time during the last small period, i.e.,
    $F_{j}^{(2)}=|[r_{j},s_{j})\cap[\theta_{j},C_{n})|.$
By this definition, the total flow time of job $j$ is
$F_{j}=F_{j}^{(1)}+F_{j}^{(2)}+p_{j}.$
If job $j$ is a final large job (i.e., $T_{j}(t)=\text{Large}$ at the end of the instance), then it has no last small period, and thus its entire waiting time equals its large waiting time:
$F_{j}=F_{j}^{(1)},\quad F_{j}^{(2)}=0.$
Note that in the version where $n$ is known, a large job can change its state to small at most once, and a small job can never become large. In that setting, we introduced the notion of a proxy job to decompose the flow time. A similar idea applies here: if a large job becomes small exactly once, then the proxy job’s waiting time corresponds precisely to the small waiting time $F_{j}^{(2)}$ defined above, while the remaining part of the original job’s waiting time corresponds to the large waiting time $F_{j}^{(1)}$.
Then we bound $F_{j}^{(1)}$ and $F_{j}^{(2)}$ separately. The analysis for $F_{j}^{(1)}$ is similar to the case where $n$ is known. Note that we have two types of kill events. One occurs when NN is updated, which we call a type-A event; the other occurs when $\phi$ becomes large, which we call a type-B event. The number of type-B kill events can be bounded similarly to the known-$n$ case, with the help of the rounded benchmark value NN.
Lemma 15.
The number of type-B kill events in the algorithm is at most $O(n^{1-\alpha}/\sqrt{m})$.
Proof.
We know that if a type-B kill event occurs at time $t$, there must be at least $\lfloor N(t)^{\alpha}\sqrt{m}\rfloor$ small jobs waiting, where $N(t)$ denotes the value of NN at time $t$. We interpret these small jobs as the creators of this kill event.
Note that when a small job $j$ creates a type-B kill event at time $t$, it remains small until NN is updated. Hence, after the kill event at $t$ and before the next update of NN, no large job will be processed on any machine unless we start job $j$. On the other hand, if we start $j$, it will not be killed again until NN is updated, since $j$ stays small during this period. Therefore, job $j$ cannot trigger another type-B kill event before NN changes.
Consequently, the number of kill events that occur while $N=x$ is at most $x^{1-\alpha}/\sqrt{m}$. Since NN increases geometrically by a factor of 22, the total number of kill events is bounded by $O\!\left(\frac{n^{1-\alpha}}{\sqrt{m}}\right),$ as $N\leq 2n$. ∎
Using the upper bounds of the number of kill events, we are able to provide an upper bound for the large waiting time of all jobs (i.e., $F_{j}^{(1)}$).
Lemma 16.
For the large waiting time, we have
$\sum_{j\in J}F^{(1)}_{j}\leq O\!\left(\frac{n^{\alpha}}{\sqrt{m}}\right)\textup{{OPT}}.$
Proof.
We first show that, for any time $t$, the number of released jobs that are not in their last small period is at most $O(n^{\alpha}\sqrt{m})$. Indeed, if a job $j$ is not in its last small period at time $t$, it must be large at some later time $t^{\prime}>t$. This means that at time $t^{\prime}$, $j$ ranks among the top-$\lfloor(2n(t^{\prime}))^{\alpha}\sqrt{m}\rfloor$ largest jobs within a larger set of released jobs. Tracing back to the earlier time $t$, since the released job set is smaller, $j$ must also be among the top-$\lfloor(2n(t^{\prime}))^{\alpha}\sqrt{m}\rfloor$ largest jobs at time $t$, and therefore also among the top-$\lfloor(2n)^{\alpha}\sqrt{m}\rfloor$ largest jobs because $n>n(t^{\prime})$. Hence, at any time $t$, the number of released jobs not in their last small period is bounded by $O(n^{\alpha}\sqrt{m})$.
Next, we follow the same argument as in Lemma 14. We shift each job’s release time to
$t^{\prime}_{j}=r_{j}+\frac{(2n)^{\alpha}}{\sqrt{m}}\cdot p_{j},$
and bound the waiting time starting from $t^{\prime}_{j}$. It is enough because we have
$\sum_{j\in J}(t^{\prime}_{j}-r_{j})=\frac{(2n)^{\alpha}}{\sqrt{m}}\sum_{j\in J}p_{j}\leq\frac{(2n)^{\alpha}}{\sqrt{m}}\cdot\textup{{OPT}}.$ (8)
If the total waiting time after $t^{\prime}$ is already bounded by $O(n^{\alpha}/\sqrt{m})\textup{{OPT}}$, then $\sum_{j\in J}F^{(1)}_{j}$ is also bounded by $O(n^{\alpha}/\sqrt{m})\textup{{OPT}}$.
For each job $j$, let its large waiting interval be $t^{\prime}_{j},t_{j})$. For every time $t\in[t^{\prime}_{j},t_{j})$, we classify it into two types:
  1. 1.
    Genuine busy time: All Mixed machines are busy at $t$, and at least $\lceil m/4\rceil$ of the running jobs are not restarted later.
  2. 2.
    Fake busy time: All Mixed machines are busy at $t$, but more than $\lceil m/4\rceil$ of the running jobs will be restarted later.
We decompose the delay $t_{j}-t^{\prime}_{j}$ into two parts, $F_{j}^{(A)}$ and $F_{j}^{(B)}$, corresponding to genuine and fake busy intervals:
$t_{j}-t^{\prime}_{j}=F_{j}^{(A)}+F_{j}^{(B)}.$
Bounding $F_{j}^{(A)}$.
During genuine busy intervals, at least $\lceil m/4\rceil$ machines are processing jobs that are never restarted. The total duration of such intervals is at most $O(\textup{{OPT}}/m)$. Each interval can contribute to the waiting time of at most $O(n^{\alpha}\sqrt{m})$ different large jobs. Hence,
$\sum_{j\in J}F_{j}^{(A)}\leq O\!\left(\frac{n^{\alpha}}{\sqrt{m}}\cdot\textup{{OPT}}\right).$ (9)
Bounding $F_{j}^{(B)}$.
Now consider fake busy intervals caused by type-B kill events. Let the $i$-th type-B kill event occur at time $\beta_{i}$. For a time $t$ to contribute to $F_{j}^{(B)}$, it must satisfy:
  1. 1.
    $t>\beta_{i-1}$ and $t>t^{\prime}_{j}$,
  2. 2.
    $t$ is not a genuine busy time.
We claim that
$\beta_{i}-t\leq p_{j}.$ (10)
Proof of the claim. Since $t$ is a fake busy time, more than $\lceil m/4\rceil$ jobs are restarted at $\beta_{i}$. Assume for contradiction that $\beta_{i}-t>p_{j}$. Then these jobs must have been running continuously since $t$, so their processing times are at least $p_{j}$. By the algorithm’s rule, job $j$ cannot have been released before starting any of these larger jobs. Let $U$ denote this set of running jobs. Then
$\beta_{i}<r_{j}+\min_{u\in U}p_{u}.$
However, by definition,
$t^{\prime}_{j}=r_{j}+\frac{(2n)^{\alpha}}{\sqrt{m}}\cdot p_{j},$
and let $P(r_{j})$ denote the total workload released up to $r_{j}$, then
$t^{\prime}_{j}\geq r_{j}+\frac{4P(r_{j})}{m}\geq r_{j}+\frac{4\sum_{u\in U}p_{u}}{m}\geq r_{j}+\min_{u\in U}p_{u}>\beta_{i},$
contradicting $\beta_{i}<t^{\prime}_{j}$. Thus, (10) holds.
Each kill event therefore contributes at most $\sum_{j\in J}p_{j}\leq\textup{{OPT}}$ to $\sum_{j}F_{j}^{(B)}$. Since the total number of type-B kill events is at most $O(n^{1-\alpha}/\sqrt{m})$, we obtain
$\sum_{j\in J}F_{j}^{(B)}\leq O\!\left(\frac{n^{1-\alpha}}{\sqrt{m}}\right)\cdot\textup{{OPT}}\leq O\!\left(\frac{n^{\alpha}}{\sqrt{m}}\right)\cdot\textup{{OPT}}.$ (11)
The last inequality holds because $\alpha$ is set to $(\sqrt{5}-1)/2>0.5$.
Finally, combining (8), (9), and (11), we have
$\sum_{j\in J}F^{(1)}_{j}\leq\sum_{j\in J}\bigl((t^{\prime}_{j}-r_{j})+F_{j}^{(A)}+F_{j}^{(B)}\bigr)\leq O\!\left(\frac{n^{\alpha}}{\sqrt{m}}\right)\cdot\textup{{OPT}}.$
Next, we aim to bound the total small waiting time. We define an increasing sequence of time points $t_{h}$, each corresponding to the start time of a job $j$ that is not in its last small period (i.e. $t_{h}<\theta_{j}$, and it is possible to be killed later). We further define $t_{0}=0$ and $t_{\bar{h}+1}=C_{n}$, where $t_{\bar{h}}$ is the last time we start processing a job that is not in its last small period and $C_{n}$ is the last completion time. These $t_{h}$ allow us to divide the jobs into sets, corresponding to $\theta_{j}$.
$S_{h}=\{j\in J\mid\theta_{j}\in[t_{h},t_{h+1})\text{ and }F^{(2)}_{j}>0\}.$
Some basic properties hold for every $j\in S_{h}$:
  1. 1.
    $j$ is started within the time interval $(\theta_{j},t_{h+1})$.
  2. 2.
    $j$ will not be killed after this start.
We now justify these two properties in order. For the first one, if $h=\bar{h}$, the property holds trivially; otherwise, since $t_{h+1}$ is the start time of another job $j^{\prime}$, where $\theta_{j^{\prime}}>t_{h+1}$. By definition, $T_{j^{\prime}}(t)$ is still large for some $t>t_{h+1}$, while $T_{j}(t)$ is small, implying $p_{j^{\prime}}>p_{j}$. Given that our algorithm prioritizes scheduling smaller waiting jobs, job $j$ must be started before $j^{\prime}$ starts at $t_{h+1}$. On the other hand, since $j$ is small after $\theta_{j}$, if it is started in $t_{h},t_{h+1})$ after $\theta_{j}$, it will not be killed.
To bound jobs’ small waiting time in every $S_{h}$, we further divide the jobs into two types.
  1. 1.
    Blocked: $j\in S_{h}$, and some machine is processing a job $j^{\prime}$ at $\theta_{j}$, where $\theta_{j}<\theta_{j^{\prime}}$. We use $B_{h}$ to denote the set of blocked small jobs in $S_{h}$.
  2. 2.
    Unblocked: otherwise, and we use $U_{h}$ to denote the set of unblocked jobs in $S_{h}$.
Let $t^{*}$ be the last time in $t_{h},t_{h+1})$ during which some machine processes a job $j$ with $\theta_{j}>t^{*}$. If no such time exists, we set $t^{*}=t_{h+1}$, which implies $U_{h}=\emptyset$.
Lemma 17.
If there exists a time in $t_{h},t_{h+1})$ at which the number of waiting blocked jobs exceeds $n^{\alpha}\sqrt{m}$, then
$\sum_{j\in S_{h}}F^{(2)}_{j}\leq F(\textup{{OPT}}(S_{h}))+O\!\left(\frac{|S_{h}|}{n^{1-\alpha}\sqrt{m}}\right)\cdot\textup{{OPT}}.$
Proof.
Since after $t^{*}$ the number of waiting blocked jobs can only decrease, there must exist a time $t\in[t_{h},t^{*})$ such that the number of waiting blocked jobs exceeds $n^{\alpha}\sqrt{m}$. All blocked jobs belong to $S_{h}$, hence they are in their last small period and are guaranteed to be small at time $t$. The only reason our algorithm does not kill the processing job $j$ that completes at $t^{*}$ with $\theta_{j}>t^{*}$ is that $T_{j}(t)$ is still small. Therefore, its size is bounded by $\frac{\textup{{OPT}}}{N(t)^{\alpha}\sqrt{m}}$.
Although NN may be smaller than $n$ since $n$ is unknown at that time, it is lower bounded by $n^{\alpha}\sqrt{m}/2$, because at least $n^{\alpha}\sqrt{m}$ small blocked jobs are waiting, implying $n(t)\geq n^{\alpha}\sqrt{m}$. Hence, the size of $j$ satisfies
$p_{j}\leq\frac{\textup{{OPT}}}{N(t)^{\alpha}\sqrt{m}}\leq\frac{2\textup{{OPT}}}{n^{\alpha^{2}}m^{\alpha/2}\sqrt{m}}\leq\frac{2\textup{{OPT}}}{n^{\alpha^{2}}\sqrt{m}}.$
We now view the schedule of $S_{h}$ in our algorithm as running NSJF on a shifted instance $S^{\prime}_{h}$, with an initial blocking time on each machine, where $S^{\prime}_{h}$ is obtained from $S_{h}$ by shifting all release times by $t_{h}$ (it is the same as viewing $t_{h}$ as time 0).
Each machine may be initially blocked because it is processing some job at time $t_{h}$, either by a job not in its last small period (for at most $t^{*}-t\leq\frac{2\textup{{OPT}}}{n^{\alpha^{2}}\sqrt{m}}$) or by a job already in its last small period (for at most $\frac{\textup{{OPT}}}{n^{\alpha}\sqrt{m}}\leq\frac{2\textup{{OPT}}}{n^{\alpha^{2}}\sqrt{m}}$ since $\alpha<1$). Hence every component of the blocking vector $\vec{b}$ is at most $\frac{2\textup{{OPT}}}{n^{\alpha^{2}}\sqrt{m}}$. It follows that the schedule of $S_{h}$ in our algorithm coincides with $\textup{{NSJF}}(S^{\prime}_{h},\vec{b})$.
By Lemma 29, writing $\vec{b}$ for the initial blocking vector,
$\sum_{j\in S_{h}}F^{(2)}_{j}\;\leq\;F(\textup{{NSJF}}(S^{\prime}_{h},\vec{b}))$ $\leq\;F(\textup{{OPT}}(S_{h}))\;+\;2|S_{h}|\cdot\frac{\textup{{OPT}}}{n^{\alpha}\sqrt{m}}\;+\;|S_{h}|\cdot\frac{2\textup{{OPT}}}{n^{\alpha^{2}}\sqrt{m}}$
$\leq\;F(\textup{{OPT}}(S_{h}))+O\!\left(\frac{|S_{h}|}{n^{\alpha^{2}}\sqrt{m}}\right)\!\cdot\textup{{OPT}}$
$=\;F(\textup{{OPT}}(S_{h}))+O\!\left(\frac{|S_{h}|}{n^{1-\alpha}\sqrt{m}}\right)\!\cdot\textup{{OPT}},$
where the last equality uses $\alpha^{2}=1-\alpha$ for $\alpha=(\sqrt{5}-1)/2$. ∎
Lemma 18.
If for all $t\in[t_{h},t_{h+1})$ the number of waiting blocked jobs is at most $n^{\alpha}\sqrt{m}$, then
$\sum_{j\in U_{h}}F^{(2)}_{j}\leq F(\textup{{OPT}}(S_{h}))+O\!\left(\frac{n^{\alpha}}{\sqrt{m}}\right)\sum_{j\in S_{h}}p_{j}+O\!\left(\frac{|S_{h}|}{n^{\alpha}\sqrt{m}}\right)\cdot\textup{{OPT}}.$
Proof.
We focus on the number of waiting blocked jobs at time $t^{*}$, which is at most $n^{\alpha}\sqrt{m}$. Let $B^{*}_{h}$ denote this set. View $t^{*}$ as time 0 and form two shifted instance $U^{\prime}_{h}$ and $B^{\prime}_{h}$ by moving every release time in $U_{h}$ and $B^{*}_{h}$ by $t^{*}$, repectively. The schedule of $U_{h}$ in our algorithm is a shifting to the schedule of $U^{\prime}_{h}$ under $\textup{{NSJF}}(U^{\prime}_{h}\cup B^{\prime}_{h},\vec{b})$, where $\vec{b}$ is the initial blocking vector from the jobs processing at $t^{*}$. By the definition of $t^{*}$, these processing jobs are already in their last small period, so each component of $\vec{b}$ is at most $O\!\left(\frac{\textup{{OPT}}}{n^{\alpha}\sqrt{m}}\right)$.
By Lemma 29,
$\sum_{j\in U_{h}}F^{(2)}_{j}\;\leq\;F(\textup{{NSJF}}(U^{\prime}_{h}\cup B^{\prime}_{h},\vec{b}))\;\leq\;F(\textup{{OPT}}^{m}(U^{\prime}_{h}\cup B^{\prime}_{h}))\;+\;O\!\left(\frac{|U^{\prime}_{h}\cup B^{\prime}_{h}|}{n^{\alpha}\sqrt{m}}\right)\textup{{OPT}},$ (12)
where $\textup{{OPT}}^{m}$ denotes the optimal preemptive, migratory schedule.
The next step is to bound the gap between $\textup{{OPT}}^{m}(U^{\prime}_{h}\cup B^{\prime}_{h})$ and $\textup{{OPT}}^{m}(U_{h})$ using the fact that $|B^{\prime}_{h}|$ is small. Consider an optimal preemptive, migratory schedule $\textup{{OPT}}^{m}(U^{\prime}_{h})$. Insert the jobs of $B^{\prime}_{h}$ one by one with migration: for each $j\in B^{\prime}_{h}$, we process it in the earliest idle slots without delaying already scheduled work; if the machine is busy, pause and resume at the next idle slot (possibly on another machine).
Let $F^{\prime}_{j}$ be the flow time of such an inserted job $j$. Then
$F^{\prime}_{j}\;=\;p_{j}+\text{wait}_{j},\qquad\text{wait}_{j}\;\leq\;\frac{1}{m}\sum_{i\in U^{\prime}_{h}\cup B^{\prime}_{h}}p_{i}.$
By the lemma’s condition, $|B^{\prime}_{h}|=|B^{*}_{h}|\leq n^{\alpha}\sqrt{m}$, hence
$F(\textup{{OPT}}^{m}(U^{\prime}_{h}\cup B^{\prime}_{h}))\;\leq\;F(\textup{{OPT}}^{m}(U^{\prime}_{h}))\;+\;O\!\left(\frac{n^{\alpha}}{\sqrt{m}}\right)\sum_{i\in U^{\prime}_{h}\cup B^{\prime}_{h}}p_{i}.$
Combining with Equation 12,
$\sum_{j\in U_{h}}F^{(2)}_{j}$ $\leq\;F(\textup{{OPT}}^{m}(U^{\prime}_{h}\cup B^{\prime}_{h}))\;+\;O\!\left(\frac{|U^{\prime}_{h}\cup B^{\prime}_{h}|}{n^{\alpha}\sqrt{m}}\right)\textup{{OPT}}$
$\leq\;F(\textup{{OPT}}^{m}(U^{\prime}_{h}))\;+\;O\!\left(\frac{n^{\alpha}}{\sqrt{m}}\right)\sum_{i\in U^{\prime}_{h}\cup B^{\prime}_{h}}p_{i}\;+\;O\!\left(\frac{|S_{h}|}{n^{\alpha}\sqrt{m}}\right)\textup{{OPT}}$
$\leq\;F(\textup{{OPT}}(S_{h}))\;+\;O\!\left(\frac{n^{\alpha}}{\sqrt{m}}\right)\sum_{i\in S_{h}}p_{i}\;+\;O\!\left(\frac{|S_{h}|}{n^{\alpha}\sqrt{m}}\right)\textup{{OPT}}.$
we call an interval $t_{h},t_{h+1})$ crowded if the number of waiting blocked jobs exceeds $n^{\alpha}\sqrt{m}$ (corresponding to Lemma 17), and uncrowded otherwise (corresponding to Lemma 18). Define $H_{1}:=\{h:[t_{h},t_{h+1})\text{ is crowded}\}$ and $H_{2}:=\{h:[t_{h},t_{h+1})\text{ is uncrowded}\}$. Then
$\sum_{h\in H_{1}}\sum_{j\in S_{h}}F^{(2)}_{j}+\sum_{h\in H_{2}}\sum_{j\in U_{h}}F^{(2)}_{j}$
$\leq\sum_{h\in H_{1}\cup H_{2}}\Big(F(\textup{{OPT}}(S_{h}))+O\!\left(\frac{|S_{h}|}{n^{1-\alpha}\sqrt{m}}\right)\textup{{OPT}}+F(\textup{{OPT}}(S_{h}))$
$\qquad\qquad+O\!\left(\frac{n^{\alpha}}{\sqrt{m}}\right)\sum_{j\in S_{h}}p_{j}+O\!\left(\frac{|S_{h}|}{n^{\alpha}\sqrt{m}}\right)\textup{{OPT}}\Big)$
$\leq 2\,F\!\left(\textup{{OPT}}\!\left(\bigcup_{h}S_{h}\right)\right)+O\!\left(\frac{\sum_{h}|S_{h}|}{n^{1-\alpha}\sqrt{m}}\right)\textup{{OPT}}+O\!\left(\frac{n^{\alpha}}{\sqrt{m}}\right)\sum_{j\in J}p_{j}$
$\leq O\!\left(\frac{n^{\alpha}}{\sqrt{m}}\right)\textup{{OPT}}.$
To bound $\sum_{j\in J}F^{(2)}_{j}$, it remains to control
$\sum_{h\in H_{2}}\sum_{j\in B_{h}}F^{(2)}_{j}$.
Lemma 19.
For the waiting time of blocked jobs during uncrowded periods,
$\sum_{h\in H_{2}}\sum_{j\in B_{h}}F^{(2)}_{j}\;\leq\;O\!\left(\frac{n^{1-\alpha}}{\sqrt{m}}\right)\textup{{OPT}}.$
Proof.
During uncrowded periods, at any time $t$ the number of waiting blocked jobs is at most $n^{\alpha}\sqrt{m}$. Two types of intervals contribute to their waiting time.
Case 1. All Small-Only machines are busy at the end. We have that the total length is at most $\frac{\sum_{j\in J}p_{j}}{m}$. Then the total waiting-time contribution is at most
$O\!\left(\frac{\sum_{j\in J}p_{j}}{m}\right)\cdot n^{\alpha}\sqrt{m}\;\leq\;O\!\left(\frac{n^{\alpha}}{\sqrt{m}}\right)\textup{{OPT}}.$
Case 2. Some Small-Only machines are not busy at the end. This only happens at a type-A kill event. If a blocked job is waiting, it is in its last small period, so it is small and all machines are busy at that moment. Hence, if some Small-Only machines are idle at the end, a type-A kill event must have occurred when NN was updated to $2N$ — the only time a small job can become large. At such an update, at most NN jobs have been released, so there are at most NN blocked jobs. Each killed job was small just before the update, hence its size is at most $\textup{{OPT}}/(N^{\alpha}\sqrt{m})$. Thus each such event contributes to $\sum_{h\in H_{2}}\sum_{j\in B_{h}}F^{(2)}_{j}$ at most
$O\!\left(\frac{\textup{{OPT}}\cdot N^{1-\alpha}}{\sqrt{m}}\right).$
As NN doubles through $1,2,4,\ldots$ up to $n<2^{k}\leq 2n$, the total contribution sums to
$O\!\left(\frac{n^{1-\alpha}}{\sqrt{m}}\right)\textup{{OPT}}.$
Combining all the lemmas above, we can prove the main theorem.
Theorem 4.
There exists an online polynomial-time deterministic algorithm with kill-and-restart for total flow time minimization, e.g., Algorithm 5, that is $O(n^{\alpha}/\sqrt{m})$-competitive against the preemptive offline solution without the prior knowledge of $n$, where $\alpha=(\sqrt{5}-1)/2$, provided that $m\geq 2$.
Proof.
By Lemma 16,
$\sum_{j\in J}F_{j}^{(1)}\;\leq\;O\!\left(\frac{n^{\alpha}}{\sqrt{m}}\right)\textup{{OPT}}.$
By Lemmas 17, 18 and 19,
$\sum_{j\in J}F_{j}^{(2)}\;\leq\;O\!\left(\frac{n^{\alpha}}{\sqrt{m}}\right)\textup{{OPT}}.$
Combining,
$\sum_{j\in J}F_{j}\;=\;\sum_{j\in J}\!\bigl(F_{j}^{(1)}+F_{j}^{(2)}+p_{j}\bigr)\;\leq\;O\!\left(\frac{n^{\alpha}}{\sqrt{m}}\right)\textup{{OPT}}.$
8 Lower Bounds for Algorithms with the Knowledge of $n$
In this section, we present our lower bound results for algorithms that have prior knowledge of $n$. Specifically, for every fixed pair $(n,m)$, we construct an adversarial instance that establishes a lower bound. The results are organized into two parts.
The first part concerns randomized algorithms. We construct hard instances via distributions, following Yao’s Minimax Principle. We begin with the single-machine setting as a warm-up, then generalize this construction to multiple machines, and finally extend it to the setting that allows kill-and-restart.
In the second part, we turn to deterministic algorithms. We first establish an $\Omega(n/\log n)$ lower bound for deterministic algorithms with kill-and-restart on a single machine. Finally, we present a lower bound of $\Omega(n/m^{2}+\sqrt{n/m})$ for deterministic non-preemptive algorithms.
8.1 Warm-Up: Randomized Lower Bound on a Single Machine
We start with the lower bound that any randomized non-preemptive online algorithm for a single machine is $\Omega(\sqrt{n})$ as a warm-up. The counterexample is constructed as Algorithm 6, which is a distribution of two possibilities: the only difference is whether the first $k$ $\varepsilon$-jobs are released at time 11 or 22.
Input: Number of jobs $n$
$k\leftarrow\lfloor\sqrt{n-2}\rfloor,\varepsilon\leftarrow 0.5n^{-2}$;
Release job 11 at time 0 with processing time 22 ;
Flip an independent fair coin: $r\sim\text{Bernoulli}(1/2)$ ;
Release $k$ jobs at time $1+r$ with processing time $\varepsilon$ ;
// released at 11 or 22
for $t=2$ to $k$ do
   Release $k$ jobs at time $1+t$ with processing time $\varepsilon$ ;
 // released at $3,\dots,k+1$
end for
Algorithm 6 Counterexample Construction for $\Omega(\sqrt{n})$ Lower Bound
Lemma 20.
The competitive ratio in flow time of any randomized non-preemptive online kill-and-restart algorithm for a single machine is $\Omega(\sqrt{n})$.
Proof.
We analyze the counterexample described in Algorithm 6. For each of the two possibilities, the optimal algorithm can arrange job 11 at time 0 or 11 correctly and delay each $\varepsilon$-job by at most $O(k^{2}\varepsilon)$. Therefore, the optimal flow time is always $O(1)$. However, for any deterministic online algorithm, it has to decide whether to start job 11 at the beginning or not:
  • If the algorithm starts job 11 at time 0, then with probability $1/2$, the algorithm will delay the first $k$ $\varepsilon$-jobs by 11 for the case the randomized $\varepsilon$-jobs are released at time 11;
  • Otherwise, the algorithm cannot start job 11 before time $k+1$ for the case the randomized $\varepsilon$-jobs are released at time 22, or it will delay at least $k$ $\varepsilon$-jobs by 11.
Therefore, the optimal deterministic algorithm has an expected flow time of $\Omega(\sqrt{n})$ against the adversary, and hence the randomized algorithm has the lower bound $\Omega(\sqrt{n})$ by Yao’s minimax theorem. ∎
8.2 Generalization to Multiple Machines
The main idea is to summarize the single-machine hard instance into a busy gadget and then replicate it $m$ times to form a busy batch. To extend this construction to multiple machines, we must address key differences related to small jobs. In particular, small jobs must have significant (non-negligible) processing time; otherwise, a deterministic algorithm could process a batch of small jobs — originally intended for $m$ machines — on a single machine with little loss. Since a batch of small jobs increases OPT by only a constant factor, a single large job is not sufficient to create a gap in the competitive ratio. Therefore, we release one large job in each busy gadget. Let $k=\Theta(\sqrt{n/m})$ be the target competitive ratio. We release $k$ batches of $m$ gadgets each. Each gadget, parameterized by a start time $t$, contains $2k+1$ jobs structured as follows and is illustrated in Figure 1:
Job Type Indices ($j$) $r_{j}$ $p_{j}$
  Large 0 0 0 22
  Fixed $\{1,\dots,k\}$ 4+(i1)/k4+(i-1)/{k} $1/k$
  Left (when $c=0$) $\{k+2,\dots,2k+1\}$ 1+(j(k+2))/k1+(j-(k+2))/{k} $1/k$
  Right (when $c=1$) $\{k+2,\dots,2k+1\}$ 2+(j(k+2))/k2+(j-(k+2))/{k} $1/k$
Figure 1: A gadget of $2k+1$ jobs defined in Algorithm 7, with two random options.
  • A Large Job: Job 0, with processing time 22 and release time $t$.
  • Fixed Small Jobs: $k$ jobs, each with processing time $1/k$, released sequentially in the interval $[t+4,t+5)$.
  • Random Small Jobs: $k$ jobs, each with processing time $1/k$, released sequentially in either $[t+1,t+2)$ (if a random coin $c=0$) or $[t+2,t+3)$ (if $c=1$).
Finally, we provide a detailed description of how the busy gadget is replicated across all $m$ machines and repeated $k$ times over time in Algorithm 7. A notable aspect is that all $m$ gadgets forming a batch $b$ (which starts at time $t=5b$) share the same random coin flip $c_{b}$.
Input: Number of jobs $n$, number of machines $m$
$k\leftarrow\lfloor\sqrt{n/(2m)+1/16}-1/4\rfloor$ ;
// $k=\Theta(\sqrt{n/m})$ and $km\cdot(2k+1)\leq n$
1ex Function Gadget($t,c$):
 $\mathcal{J}\leftarrow\varnothing$ ;
 // Gadget job (multi-)set
 $\mathcal{J}\leftarrow\mathcal{J}\cup\{(r_{0}=t,p_{0}=2)\}$;
 $\mathcal{J}\leftarrow\mathcal{J}\cup\{(r_{i}=t+4+(i-1)/k,p_{i}=1/k):i=1\text{\penalty 10000\ to\penalty 10000\ }k\}$;
 $\mathcal{J}\leftarrow\mathcal{J}\cup\{(r_{k+i}=t+1+c+(i-1)/k,p_{k+i}=1/k):i=1\text{\penalty 10000\ to\penalty 10000\ }k\}$;
 return $\mathcal{J}$;
1exfor $b=0$ to k1k-1 do
   Flip an independent fair coin: $c_{b}\sim\text{Bernoulli}(1/2)$ ;
 // 0: Left, 11: Right
   Release the $b$-th batch consisting of $m$ copies of $\textnormal{{Gadget}}(5b,c_{b})$ ;
end for
Algorithm 7 Counterexample Construction for $\Omega(\sqrt{n/m})$ Lower Bound
Lemma 21.
$\textup{{OPT}}\leq 6mk$ for the instance constructed in Algorithm 7.
Proof.
In each busy gadget, we can schedule the large job either at $5b+0$ or $5b+2$ depending on $c$, and every small job upon arrival. As a result, the total flow time in this gadget is at most 44 (the large job) +1+1 (fixed small jobs) +1+1 (random small jobs). Therefore $\textup{{OPT}}\leq 6mk$ by summing up all gadgets. ∎
Moving to the analysis of an arbitrary online algorithm, we aim to prove that the algorithm’s flow time is at least $\Omega(mk^{2})$. First, we will show that when a large job is placed such that it conflicts with the small jobs (having a intersection of at least $0.5$), it significantly increases the flow time of those small jobs.
Definition 1.
In a given batch, the set of mkmk jobs released in a time interval $[t,t+1)$ constitutes a small job period. A large job conflicts with this period if its execution starts within the interval [t1.5,t+0.5][t-1.5,t+0.5]. This definition guarantees that the large job’s execution (of length 2) overlaps with the period $[t,t+1)$ for a duration of at least $0.5$.
Lemma 22.
If $x$ large jobs conflicts with a small job period at $[t,t+1)$, then the total flow time of these small jobs within $[t,t+1)$ is at least $xk/8$.
Proof.
Define $f(u)=m\cdot\lceil ku\rceil$ as the number of small jobs released at time $t+u$, and let $g(u)$ be the number of machines not processing large jobs at time $u$. The number of small jobs that can be completed by time $u$ is at most $\lfloor k\int_{0}^{u}g(y)dy\rfloor$. The total flow time of the kmkm small jobs in the batch is no less than
$F$ $\geq\int_{0}^{1}f(u)-\lfloor k\int_{0}^{u}g(y)\mathrm{d}y\rfloor\,\mathrm{d}u$
$\geq k\cdot\int_{0}^{1}mu-\int_{0}^{u}g(y)\mathrm{d}y\,\mathrm{d}u$
$=k\cdot\int_{0}^{1}(m-g(u))\cdot(1-u)\,\mathrm{d}u,$
If no large jobs conflict with this period, $g(u)$ is defined as $m$ for every $u\in[0,1)$. By the condition of the lemma, we have $x$ jobs that decrease $g(u)$ by 11 with a duration of at least $0.5$. The minimum contribution of the lower bound occurs when each large job conflicts the period at $[0.5,1)$, yielding
$F$ $\geq xk\cdot\min_{s\in[0,0.5]}\int_{s}^{s+0.5}1\cdot(1-u)\,\mathrm{d}u$
$=xk\cdot\int_{0.5}^{1}(1-u)\,\mathrm{d}u=xk/8.$
Lemma 22 establishes that an incorrectly placed large job incurs a total flow time of $\Omega(k)$. In the next lemma, we aim to prove that a deterministic algorithm must either suffer a large flow time because of incorrectly placed large jobs or defer the large jobs to the next gadget. We capture the total flow time using a potential function $\Phi(b,d)$ that represents a lower bound on the expected total flow time in the time interval $5b-0.5,\infty)$, conditioned on the algorithm having $d$ unscheduled large jobs at time 5b0.55b-0.5.
Theorem 5.
Any randomized online algorithm for minimizing the total flow time on parallel machines has a competitive ratio of at least $\Gamma(n,m)$, for every fixed pair $(n,m)$, where $\Gamma(n,m)=\Omega(\sqrt{n/m})$.
Proof.
For a fixed busy batch $b$, we analyze any fixed deterministic algorithm’s behavior on large jobs during the interval [5b0.5, 5(b+1)0.5)[5b-0.5,\;5(b+1)-0.5). [^5] Note that when $b=0$, the negative part doesn’t exist, and it will not affect the discussion when the negative part come into consideration. Since $b$ is fixed in context, we simplify the interval to [0.5, 4.5)[-0.5,\;4.5).
For a fixed arbitrary realization of $c_{0},\dots,c_{b-1}$, consider two possible realizations of $c_{b}$. Let $N_{b}$ be the random variable (w.r.t. $c_{b}$) representing the number of large jobs the algorithm starts within [0.5, 4.5)[-0.5,\;4.5), and $I_{b}$ be the random variable (w.r.t. $c_{b}$) representing the number of large jobs scheduled in this batch that conflict with small job periods.
Note that before time 11, the algorithm does not know the realization of $c_{b}$ this round. Therefore, its behavior before time 11 is not related to the result of $c_{b}$.
  • Let $x$ be the number of large jobs started in [0.5, 1)[-0.5,\;1). If small jobs are Left, all these $x$ jobs contribute to II.
After time 11, the algorithm’s scheduling decisions may depend on $c_{b}$. We consider the two cases separately:
  • Right case (w.p. $0.5$): Small jobs occupy [2, 3)[2,\;3) and [4, 5)[4,\;5). Any large job started in [1, 4.5)[1,\;4.5) must conflict with a small job. Let $y$ be the number of such large jobs, which will contribute $y$ to $I_{b}$.
  • Left case (w.p. $0.5$): Small jobs occupy [1, 2)[1,\;2) and [4, 5)[4,\;5). The algorithm can safely schedule up to $m$ large jobs in the gap [1.5, 2.5)[1.5,\;2.5) without conflicting any small job period. Let $z$ be the number of large jobs scheduled in [1, 1.5),[2.5, 4.5)[1,\;1.5),[2.5,\;4.5); then at least $z$ jobs conflict with small jobs.
Combining all cases, we compute the expectations:
$\mathbb{E}_{c_{b}}(N_{b})\leq x+\frac{y}{2}+\frac{z+m}{2},\qquad\mathbb{E}_{c_{b}}(I_{b})\geq\frac{x}{2}+\frac{y}{2}+\frac{z}{2}.$
Let $F^{(S)}_{b}$ denote the total flow time of small jobs in batch $b$. By Lemma 22,
$\mathbb{E}_{c_{b}}(F^{(S)}_{b})\geq\frac{k}{8}\cdot\mathbb{E}_{c_{b}}(I_{b})\geq\frac{k}{8}\cdot\left(\frac{\mathbb{E}_{c_{b}}(N_{b})-\frac{m}{2}}{2}\right)\geq\frac{k}{16}\cdot\left(\mathbb{E}_{c_{b}}(N_{b})-\frac{m}{2}\right).$
The above argument applies to each $b$ with all possible $c_{0},\dots,c_{b-1}$, by summing them up we have
$\mathbb{E}\left(F^{(S)}\right)\geq\sum_{b=0}^{k-1}\mathbb{E}\left(F^{(S)}_{b}\right)\geq\frac{k}{16}\cdot\sum_{b=0}^{k-1}\left(\mathbb{E}(N_{b})-\frac{m}{2}\right).$
For the total flow time $F^{(L)}$ incurred by large jobs, we count the number of unfinished and newly released large jobs at time 0 for each batch. By the fact that unfinished jobs will be completed later than $t+1$ for a batch starting at $t$, they must contribute to the total flow time in $[t,t+1)$. Therefore,
$\mathbb{E}(F^{(L)})$ $\geq\sum_{i=0}^{k-1}\left((i+1)\cdot m-\sum_{j=0}^{i-1}\mathbb{E}(N_{b})\right)\geq\mathbb{E}\left(\sum_{i=0}^{k-1}\sum_{b=0}^{i}\left(m-\mathbb{E}(N_{b})\right)\right)$
$=\sum_{b=0}^{k-1}\left(m-\mathbb{E}(N_{b})\right)\cdot(k-b)$
$\geq\frac{k}{16}\cdot\sum_{b=0}^{\lfloor 15k/16\rfloor}\left(m-\mathbb{E}(N_{b})\right).$
Combining the two bounds, we have
$\mathbb{E}(F^{(L)})+\mathbb{E}(F^{(S)})\geq\frac{k}{16}\cdot\sum_{b=0}^{\lfloor 15k/16\rfloor}\frac{m}{2}=\Omega(k^{2}m).$
By Lemma 21 $\textup{{OPT}}=O(mk)$, any deterministic algorithm is then $\Omega(k)=\Omega(\sqrt{n/m})$ competitive on the counterexample, and we can conclude the theorem by Yao’s minimax principle. ∎
8.3 Generalization to Multiple Machines with Kill-and-Restart
The details of the construction are presented in Figure 2 and Algorithm 8. The overall structure of this hard instance follows that of Section 8.2, with one key difference: Each gadget now contains two large jobs of different sizes. This modification strengthens the lower bound instance, as algorithms must now not only decide whether to schedule a large job, but also determine which type of large job to schedule. Intuitively, in the previous hard instance, the algorithm could start a large job early and decide whether to kill it after the randomness is revealed. However, in the current construction, the algorithm must decide which large job to start in advance; a wrong decision can lead to a substantial cost.
Job Type Indices (ii) $r_{i}$ $p_{i}$
  Large 0 0 0 22
  Large 1 11 0 33
  Fixed $\{2,\dots,k+1\}$ 6+(i2)/k6+(i-2)/{k} $1/k$
  Left (when $c=0$) $\{k+2,\dots,2k+1\}$ 2+(i(k+2))/k2+(i-(k+2))/{k} $1/k$
  Right (when $c=1$) $\{k+2,\dots,2k+1\}$ 3+(i(k+2))/k3+(i-(k+2))/{k} $1/k$
Figure 2: A gadget of $2k+2$ jobs defined in Algorithm 8, with two random options.
Input: Number of jobs $n$, number of machines $m$
$k\leftarrow\lfloor\sqrt{n/(2m)+1/4}-1/2\rfloor$ ;
// $k=\Theta(\sqrt{n/m})$ and $km\cdot(2k+2)\leq n$
1ex Function Gadget($t,c$):
 $\mathcal{J}\leftarrow\varnothing$ ;
 // Gadget job (multi-)set
 $\mathcal{J}\leftarrow\mathcal{J}\cup\{(r_{0}=t,p_{0}=2),\ (r_{1}=t,p_{1}=3)\}$;
 $\mathcal{J}\leftarrow\mathcal{J}\cup\{(r_{i}=t+6+(i-2)/k,p_{i}=1/k):i=2\text{\penalty 10000\ to\penalty 10000\ }k+1\}$;
 $\mathcal{J}\leftarrow\mathcal{J}\cup\{(r_{k+i}=t+2+c+(i-2)/k,p_{k+i}=1/k):i=2\text{\penalty 10000\ to\penalty 10000\ }k+1\}$;
 return $\mathcal{J}$;
1exfor $b=0$ to k1k-1 do
   Flip an independent fair coin: $c_{b}\sim\text{Bernoulli}(1/2)$ ;
 // 0: Left, 1: Right
   Release the $b$-th batch consisting of $m$ copies of $\textsc{Gadget}(7b,c_{b})$ ;
end for
Algorithm 8 Counterexample Construction for $\Omega(\sqrt{n/m})$ Lower Bound
Lemma 23.
$\textup{{OPT}}\leq 11mk$ for the instance constructed in Algorithm 8.
Proof.
In each busy gadget, if $c=0$, then schedule large job 0 at 0 and large job 11 at 33; if $c=1$, then schedule large job 11 at 0 and large job 0 at 44. In this way, no large jobs intersect with any small jobs. As a result, the total flow time in this gadget is at most 99 (large jobs) +1+1 (fixed small jobs) +1+1 (random small jobs). Therefore, $\textup{{OPT}}\leq 11mk$ by summing up all gadgets. ∎
Theorem 6.
Any randomized online algorithm for minimizing the total flow time on parallel machines has a competitive ratio of at least $\Gamma(n,m)$, for every fixed pair $(n,m)$, where $\Gamma(n,m)=\Omega(\sqrt{n/m})$, even with the capability of kill-and-restart.
Proof.
For a fixed busy batch $b$, we analyze any fixed deterministic algorithm’s behavior on large jobs during the interval [7b0.5, 7(b+1)0.5)[7b-0.5,\;7(b+1)-0.5). Since $b$ is fixed in context, we simplify the interval to [0.5, 6.5)[-0.5,\;6.5).
For a fixed realization of $c_{0},\dots,c_{b-1}$, consider two possible realizations of $c_{b}$. Let $N_{b,2},N_{b,3}$ be the random variable (w.r.t. $c_{b}$) representing the number of jobs with processing time $2,3$ the algorithm start within [0.5, 6.5)[-0.5,\;6.5) and don’t restart, and $I_{b,2},I_{b,3}$ defined similarly that intersect with small job periods. Let $N_{b}=N_{b,2}+N_{b,3},I_{b}=I_{b,2}+I_{b,3}$.
Note that before time 22, the algorithm does not know the realization of $c_{b}$ this round. Therefore, its behavior before time 22 is not related to the result of $c_{b}$.
  • Let $x_{2}$ be the number of jobs with processing time 22 started in [0.5, 2)[-0.5,\;2).
  • Let $x_{3}$ be the number of jobs with processing time 33 started in [0.5, 2)[-0.5,\;2).
After time 22, the algorithm scheduling decisions may depend on $c_{b}$. We consider the two cases separately:
  • Left case (w.p. $0.5$): None of the $x_{3}$ jobs can finish without intersection, and at most $x_{2}^{(L)}=\min\{m,x_{2}\}$ out of the $x_{2}$ jobs can finish without intersection. Let $y_{2}^{(L)}$ be the number of jobs with processing time 22 starting in $[2,6.5)$ that finish without intersection, and $y_{3}^{(L)}$ be the number for processing time 33. At most $m$ jobs out of $y_{2}^{(L)}+y_{3}^{(L)}$ can finish without intersection.
  • Right case (w.p. $0.5$): At most $m$ out of these $x_{2}+x_{3}$ jobs could end up without intersection with small jobs; let $x_{2}^{(R)}$ and $x_{3}^{(R)}$ be the number of such jobs, respectively. At most $m$ jobs with processing time 22 released in $[2,6.5)$ can finish without intersection; let $y_{2}^{(R)}$ be the number of such jobs. None of the jobs with processing time 33 can finish without intersection.
Then,
$\mathbb{E}_{c_{b}}(N_{b}-I_{b})$ $\leq\frac{1}{2}\left(x_{2}^{(L)}+y_{2}^{(L)}+y_{3}^{(L)}\right)+\frac{1}{2}\left(x_{2}^{(R)}+x_{3}^{(R)}+y_{2}^{(R)}\right)$ (13)
$\leq\frac{1}{2}\left(x_{2}^{(L)}+m\right)+\frac{1}{2}\left(x_{2}^{(R)}+x_{3}^{(R)}+y_{2}^{(R)}\right)$ (14)
$\leq\frac{3}{2}m+\frac{1}{2}x_{2}^{(L)}\leq\frac{3}{2}m+\frac{1}{2}x_{2}$ (15)
and
$\mathbb{E}_{c_{b}}(N_{b,3}-I_{b,3})$ $\leq\frac{1}{2}y_{3}^{(L)}+\frac{1}{2}x_{3}^{(R)}\leq\frac{1}{2}m+\frac{1}{2}\left(m-x_{2}\right)\leq m-\frac{1}{2}x_{2}.$ (16)
[…]
$\mathbb{E}_{c_{b}}(N_{b}-I_{b})+\mathbb{E}_{c_{b}}(N_{b,3}-I_{b,3})$ $\leq\frac{3}{2}m+\frac{1}{2}x_{2}+m-\frac{1}{2}x_{2}=\frac{5}{2}m,$
and hence $\mathbb{E}(N_{b}-I_{b})+\mathbb{E}(N_{b,3}-I_{b,3})\leq\frac{5}{2}m$ since it doesn’t rely on the specific values of $c_{0},\ldots,c_{b-1}$.
Consider the flow time by the large jobs. If we consider the jobs with processing time 22 or 33 and their intersections only, we have
$\mathbb{E}(F^{(2,3)})$ $\geq\sum_{b=0}^{k-1}\sum_{i=0}^{b}(2m-E(N_{b}))+\sum_{b=0}^{k-1}E(I_{b})\cdot k/8$
$\geq\sum_{b=0}^{k-1}(2m-E(N_{b}))\cdot(b-k)+\sum_{b=0}^{k-1}E(I_{b})\cdot k/8$
$\geq\sum_{b=0}^{\lfloor\frac{7}{8}k\rfloor}\left(2m-E(N_{b}-I_{b})\right)\cdot k/8.$
Similarly, consider the jobs with processing time 33 and their intersections only, we have
$\mathbb{E}(F^{(3)})$ $\geq\sum_{b=0}^{\lfloor\frac{7}{8}k\rfloor}(m-\mathbb{E}(N_{b,3}-I_{b,3}))\cdot k/8.$
Therefore, the total expectation of total flow time is at least
$\mathbb{E}(F)=\frac{1}{2}\cdot\left(2\cdot\mathbb{E}(F)\right)$ $\geq\frac{1}{2}\cdot\left(\mathbb{E}(F^{(2,3)})+\mathbb{E}(F^{(3)})\right)$
$\geq\frac{k}{16}\cdot\left(\sum_{b=0}^{\lfloor\frac{7}{8}k\rfloor}3m-\left(\mathbb{E}(N_{b}-I_{b})+\mathbb{E}(N_{b,3}-I_{b,3})\right)\right)$
$\geq\frac{k}{16}\cdot\left(\sum_{b=0}^{\lfloor\frac{7}{8}k\rfloor}m/2\right)=\Omega(k^{2}m).$
By Lemma 23, $\textup{{OPT}}=O(km)$, hence any deterministic algorithm is $\Omega(k)=\Omega(n/m)$ competitive and the theorem is proved. ∎
8.4 Deterministic Algorithms: Single Machine with Kill-and-Restart
In this section, we show that the competitive ratio of any deterministic online algorithm with kill-and-restart is at least $\Omega(n/\log n)$ for the single-machine case. The adversarial instance, detailed in Algorithm 9, uses a two-phase construction.
  • In Phase 1, the adversary uses two large jobs and carefully timed small jobs to force the online algorithm, ALG, to end the phase with at least one large job still unstarted. If ALG manages to complete both large jobs, it is forced to incur an $\Omega(n)$ flow time, ending the game.
  • In Phase 2, the adversary exploits the presence of the unstarted large job. It releases a stream of tiny $\varepsilon$-jobs at logarithmically increasing intervals before each restart of the large job. As we will show, this construction forces ALG to incur a total flow time of $\Omega(n/\log n)$, regardless of its specific kill-and-restart strategy.
Input: Deterministic online algorithm ALG, number of jobs $n$
$c\leftarrow 0.5,\varepsilon\leftarrow 0.5n^{-2}$;
Phase 1: Construct an unsolved job
   Release job 11 at 0 with processing time 44 ;
   Release job 22 at 22 with processing time 11 ;
 if ALG is working on job 11 before 33 then
      Release $c/2\cdot n$ jobs at 33 with processing time $\varepsilon$ ;
      Release $c/2\cdot n$ jobs at 77 with processing time $\varepsilon$ ;
    $t^{*}\leftarrow 7$;
 else
      Release $c\cdot n$ jobs at 55 with processing time $\varepsilon$ ;
    $t^{*}\leftarrow 5$;
   end if
  Wait until ALG starts any job released at $t^{*}$, let $t$ be the current time ;
 if jobs $1,2$ are both completed or $t-t^{*}\geq 1$ then
      Release n2cnn-2-cn jobs at time $t$ with processing time $\varepsilon$ ;
    // Ensure $n$ jobs
    return ;
    // $\textup{{ALG}}=\Omega(n),\textup{{OPT}}=O(1)$
 else
      Wait $c\cdot n\cdot\varepsilon$ units of time ;
    // Ensure OPT finishes all jobs
    continue to Phase 2 with any unsolved job $u$ among jobs $1,2$ ;
   end if
Phase 2: Exploit the unsolved job $u$
 $T=\{H_{i}/H_{n}:i=1\dots n,H_{i}=\sum_{k=1}^{i}1/k\}$ ;
 // Set of thresholds
 $l\leftarrow n-2-cn$ ;
 // Number of jobs left
 while $l>0$ do
      Wait until ALG is working on job $u$, let $t$ be the current time ;
      Let $t^{\prime}$ be the time when ALG starts job $u$ ;
    $\tau\leftarrow\min\{x\in T\mid x>t-t^{\prime}\}$ ;
    if ALG hasn’t kill job $u$ before time $t^{\prime}+\tau$ then
         Release a job at time $t^{\prime}+\tau$ with processing time $\varepsilon$ ;
       $l\leftarrow l-1$ ;
      end if
   end while
Algorithm 9 Adversary for Deterministic Lower Bound $\Omega(n/\log n)$
Lemma 24.
Against the adversary in Phase 11 of Algorithm 9, any deterministic online algorithm either incurs a flow time of $\Omega(n)$ or proceeds to Phase 2 with at least one of job 1 or job 2 unstarted. In all cases, OPT remains $O(1)$ and all jobs are solved.
Proof.
See Figure 3 which enumerates all possible strategies for ALG that (1) only make new decisions at new job arrivals, and (2) do not let the machine idle since ALG has the ability to restart. Though it only enumerates these well-formed algorithmic decisions and does not detail other strategies that do not satisfy these conditions, we argue that the description of the adaptive adversary is sufficiently valid for them.
Figure 3: The cases in Phase 1 for ALG.
  1. 1.
    If ALG is working on job 11 at time 33, then OPT could solve job 22 by 33 and job 11 by $7+c/2\cdot n\varepsilon$. For ALG, however:
    1. (a)
      If ALG keeps working on job 11 until finished, the flow time is at least $c/2\cdot n$;
    2. (b)
      If ALG kills at time 33 to solve $\varepsilon$-jobs, switch to job 22 and then switch to job 11:
      1. i.
        If it doesn’t switch to $\varepsilon$-jobs at time 77, the flow time is at least $c/2\cdot n$;
      2. ii.
        Otherwise ALG must switch to $\varepsilon$-jobs at time 77, leaving job 11 unsolved.
  2. 2.
    Otherwise, ALG cannot finish job $1,2$ by time 55 while OPT can. In this case ALG would switch to job 22 at time 22 and then to solve job 11, or it would fall back to case 1.
    1. (a)
      If ALG keeps working on job 11 until finished, the flow time is at least $2c\cdot n$;
    2. (b)
      Otherwise ALG must switch to $\varepsilon$-jobs at time 55, leaving job 11 unsolved.
Lemma 25.
In Phase 2 of Algorithm 9, an online algorithm ALG must schedule the jobs released in this phase along with an unstarted large job $u$ (with $p_{u}\geq 1$) carried over from Phase 1. The total flow time for ALG on the Phase 2 jobs is at least $\Omega(n/\log n)$. In contrast, an optimal offline solution for an instance containing only the Phase 2 jobs (with no initial unstarted job) has a flow time of $O(1)$.
Proof.
Let N=n2cnN=n-2-cn be the total number of $\varepsilon$-jobs released in this phase. Consider any deterministic strategy employed by ALG. This strategy will result in some number of restarts of the large job $u$, say $r\geq 0$. These restarts partition the NN small jobs into $r+1$ groups. Let $a_{i}$ be the number of $\varepsilon$-jobs released during the $i$-th run of job $u$ (for $i=1,\dots,r+1$), where a run is a period of continuous work on $u$. We have $\sum_{i=1}^{r+1}a_{i}=N$.
The key insight is that the total flow time of the small jobs is lower-bounded by the same quantity, regardless of how the $a_{i}$ are partitioned. By construction of the adversary, we can count total flow time before the next kill (or completion when $i=r+1$) of the large job by summing up the flow time of $j$-th job in the $i$-th run, where we use the $j=0$ to describe the large job and $j=1,\dots,a_{i}$ to describe the $\varepsilon$-jobs in release order:
$\sum_{i=1}^{r+1}\sum_{j=0}^{a_{i}}(H_{a_{i}}-H_{j})/H_{n}$ $=\sum_{i=1}^{r+1}\sum_{j=0}^{a_{i}}\left(\sum_{k=j+1}^{a_{i}}\frac{1}{k}\right)/H_{n}$
$=\sum_{i=1}^{r+1}\sum_{k=1}^{a_{i}}\frac{1}{k}\sum_{j=0}^{k-1}1/H_{n}$
$=\sum_{i=1}^{r+1}\frac{a_{i}}{H_{n}}=\frac{N}{H_{n}}.$
Since $N=\Theta(n)$ and $H_{n}=\Theta(\log n)$, the flow time is $\Omega(n/\log n)$. In contrast, an offline OPT can serve every $\varepsilon$-job upon its arrival, leading to a total flow time of $N\cdot\varepsilon=O(1)$. ∎
Theorem 7.
All deterministic algorithms with kill-and-restart on single-machine total flow time minimization have a competitive ratio of at least $\Omega(n/\log n)$.
Proof.
From Lemma 24 and Lemma 25, the flow time of any deterministic algorithm is at least $\Omega(n/\log n)$. Meanwhile, the flow time of OPT is $O(1)$ in both phases. Therefore, the competitive ratio is $\Omega(n/\log n)$. ∎
8.5 Deterministic Algorithms: Multiple Machines without Kill-and-Restart
The $\Omega(n/m^{2}+\sqrt{n/m})$ follows by combining our randomized lower bound of $\Omega(\sqrt{n/m})$ with the deterministic $\Omega(n/m^{2})$ lower bound from DBLP:journals/dam/EpsteinS06, as presented below.
Theorem 8.
Any deterministic online algorithm for minimizing the total flow time on parallel machines has a competitive ratio of at least $\Gamma(n,m)$, for every fixed $n$ and $m$, where $\Gamma(n,m)=\Omega(n/m^{2}+\sqrt{n/m})$.
Proof.
We first show that for every $n$, $m$, we can construct an adversary to prove $\Omega(n/m^{2})$, the same as DBLP:journals/dam/EpsteinS06. We still present it for completeness. Consider the following case for any deterministic algorithm ALG:
  • A job with processing time 11 is released at time 0.
  • Once ALG starts a job at $t$, the adversary releases $n$ jobs with processing time $1/n$ in $n/m$ batches, where each batch contains $m$ jobs released at time $t+1/n,t+2/n,\dots,t+1/m$ respectively.
OPT can start all small jobs at their release time immediately; it can also solve the first job at 0 if $t>1$, or solve the first job at 22 otherwise. Hence, OPT is $O(1)$ in both cases. For ALG, as it can only use m1m-1 machines to solve the small jobs, at most m1m-1 jobs can be solved between the releases of two small job batchs. The total flow time for ALG is at least $\sum_{i=1}^{n/m}i/n=\Omega(n/m^{2})$. Therefore, the competitive ratio is $\Omega(n/m^{2})$.
Finally, combining with our $\Omega(\sqrt{n/m})$ lower bound for all $n$, $m$, it proves the lemma. ∎
9 Lower Bounds for Algorithms without the Knowledge of $n$
In this section, we establish strong lower bounds for the “unknown nn” setting by the construction of a powerful oblivious adversary. Unlike an adaptive adversary, this adversary does not react to an algorithm’s random choices. Instead, it operates by analyzing an algorithm’s description to derive its probability distribution for a single, initial job released at time 0. Based on this distribution, the adversary classifies the algorithm into one of two categories: those that tend to start the job early, and those that tend to delay it. For each category, the adversary then commits to a specific instance to exploit that particular behavior.
Theorem 9.
No randomized algorithm can achieve a competitive ratio of $o(n)$ without prior knowledge of $n$ in the single-machine setting.
Proof.
Recall that if a randomized algorithm ALG could achieve a competitive ratio of $o(n)$, then for every constant $C>0$, there would exist a positive integer $n_{0}$ such that
$\frac{\mathbb{E}[\textup{{ALG}}]}{\textup{{OPT}}}\leq C\cdot n,\quad\text{for all }n\geq n_{0}.$
We now prove a contradiction: for any randomized algorithm ALG, there exists a constant $C>0$ such that for every positive integer $n_{0}$, one can construct an instance II with $|I|=n\geq n_{0}$ satisfying
$\frac{\mathbb{E}[\textup{{ALG}}]}{\textup{{OPT}}}>C\cdot n.$
Hence, no randomized algorithm can achieve a competitive ratio of $o(n)$.
Adversary construction. We construct an oblivious adversary as follows. At time 0, the adversary releases a single job of size 22. For any algorithm, let $p(t)$ denote the probability density function that ALG starts this job at time $t$ if only the single job is present, which is known by the oblivious adversary since the probability distribution solely depend on the description of the algorithm and does not depend on algorithm’s randomized results. The algorithm must fall into one of the following categories (the parameters $x$ and $y$ will be fixed later):
  1. A.
    There exists an interval $[t,t+1)$, where $t\in\{1,2,\dots,x\}$, such that
    $\int_{t}^{t+1}p(t)\,dt\geq y.$
    In this case, the adversary releases $n_{1}$ dummy jobs of size 0 at time $t+1$. With probability at least $y$, ALG eventually starts the size-22 job within $[t,t+1)$, all these $n_{1}$ jobs complete no earlier than time $t+2$, so each contributes at least 11 to flow time and the algorithm’s total flow time is at least $n_{1}$, while an optimal solution completes all jobs by time 44, i.e., $\textup{{OPT}}\leq 4$. Therefore,
    $\frac{\mathbb{E}[\textup{{ALG}}]}{n\cdot\textup{{OPT}}}>\frac{yn_{1}}{4(n_{1}+1)}\to\frac{y}{4},\quad\text{as }n_{1}\to\infty.$
  2. B.
    For all $t\in\{1,2,\dots,x\}$, it holds that $\int_{t}^{t+1}p(t)\,dt<y$. In this case, the probability that the size-22 job has not been started by time $x$ is at least 1xy1-xy. The adversary releases $n_{0}$ dummy jobs of size 0 at time $x+1$. If ALG delays the size-22 job until after time $x$, each dummy job has flow time at least $x$, while $\textup{{OPT}}\leq 2$. Therefore,
    $\frac{\mathbb{E}[\textup{{ALG}}]}{n\cdot\textup{{OPT}}}>\frac{(1-xy)x}{2(n_{0}+1)}.$
Parameter setting.
Let $x=n_{0}$ and $y=\frac{1}{2n_{0}}$, so that $xy=\frac{1}{2}$. Then, in the Type A case, there exists an instance II with $|I|=n\gg n_{0}$ such that $\mathbb{E}[\textup{{ALG}}]/\textup{{OPT}}>(1/(8n_{0}))\cdot n$; in the Type B case, there exists an instance II with $|I|=n=n_{0}+1>n_{0}$ such that $\mathbb{E}[\textup{{ALG}}]/\textup{{OPT}}>(1/8)\cdot n$.
Therefore, for every constant $0<C\leq 1/8$, there exists a corresponding integer $n_{0}=1/(8C)$ such that
$\frac{\mathbb{E}[\textup{{ALG}}]}{\textup{{OPT}}}>C\cdot n,$
for some instance with $n\geq n_{0}$. Consequently, no randomized algorithm can achieve a competitive ratio of $o(n)$. ∎
Corollary 26.
No randomized algorithm can achieve a competitive ratio of $o(n/m^{2}+\sqrt{n/m})$ without prior knowledge of $n$ in the multiple-machine setting.
Proof.
It is known that even when $n$ is known in advance, there exists a lower bound of $\Omega(\sqrt{n/m})$. Therefore, it suffices to generalize Theorem 9 to obtain a hard instance with ratio $\Omega(n/m^{2})$, from which the theorem follows.
Adversary construction. We use the same setup as in the single-machine case: the adversary first releases a size-2 job at time 0 and observes the algorithm’s random behavior. The only difference lies in how we construct the hard instance for Type A algorithms in the multi-machine setting.
Hard instance for Type A.
If ALG is of Type A, the adversary releases $n_{1}$ additional jobs of size $1/k$, where $k=n_{1}/m$. The release pattern is as follows:
$t+1,\;t+1+\frac{1}{k},\;t+1+\frac{2}{k},\;\dots,\;t+1+\frac{k-1}{k},$
and at each of these $k$ release times, $m$ new jobs (one per machine) are released. Hence, the total number of small jobs is $n_{1}=m\cdot k$.
The optimal scheduler can avoid any conflict between the size-2 job and the small jobs, so that all small jobs start immediately at their release times. The total flow time is therefore at most $m+4$. However, with probability at least $y$, the algorithm starts the size-2 job in $[t,t+1)$, leaving only m1m-1 machines for small jobs. In this case, the total flow time is at least
$m+k\cdot 1=m+\frac{n_{1}}{m}\geq\frac{n_{1}}{m}.$
Therefore,
$\frac{\mathbb{E}[\textup{{ALG}}]}{\textup{{OPT}}}\geq\frac{y\cdot(n_{1}/m)}{m+4}\geq\frac{yn_{1}}{5m^{2}},$
where the last inequality uses $m\geq 1$. As $n_{1}\to\infty$, this ratio approaches $yn/(5m^{2})$.
Hard instance for Type B.
The construction and argument are identical to the single-machine case: if the algorithm delays the size-2 job beyond time $x$, the expected ratio satisfies
$\frac{\mathbb{E}[\textup{{ALG}}]}{\textup{{OPT}}}>\frac{(1-xy)x}{2(n_{0}+1)}.$
Parameter setting.
We use the same parameters as before, setting $x=n_{0}$ and $y=1/(2n_{0})$ so that $xy=1/2$. Under this choice, the ratio for Type A algorithms is at least $(1/(10n_{0}))\cdot n/m^{2}$, while for Type B algorithms it is at least $(1/8)\cdot n\geq(1/8)\cdot n/m^{2}$.
Therefore, for every $0<C\leq 1/10$, there exists a corresponding integer $n_{0}=1/(10C)\geq 1$ such that
$\frac{\mathbb{E}[\textup{{ALG}}]}{\textup{{OPT}}}>C\cdot\frac{n}{m^{2}},$
for some instance with $n>n_{0}$. Consequently, no randomized algorithm can achieve a competitive ratio of $o(n/m^{2})$. ∎
Appendix A Schedule Small Jobs: Non-Preemptive Shortest Job First
In this section, we analyze the performance of the Non-Preemptive Shortest Job First (NSJF) algorithm in a generalized version, with the concept of blocking period: for blocking vector $\vec{b}$, where each $b_{i}$ denotes the initial blocking time of machine $i$ — that is, machine $i$ is unavailable for processing during the interval $0,b_{i})$, we aim to prove the following bound:
$F(\textup{{NSJF}}(\vec{b}))\leq F(\textup{{OPT}}(\vec{0}))+\frac{nB}{m}+2n\tau.$
where $B=\sum_{i=1}^{m}b_{i}$.
Since the context is clear, we will refer to $\textup{{NSJF}}(\vec{b})$ simply as NSJF and $\textup{{OPT}}(\vec{0})$ as OPT. Due to the initial blocking times, the number of active machines may differ between NSJF and OPT. To capture this, we define a function $a(t)$ to be the number of active machines at time $t$ under NSJF. We also define the total active power of NSJF up to time $t$ as:
$A(t)=\int_{0}^{t}a(y)\,dy.$
Next, for any threshold $p$ and time $t$, we define $J_{\leq p}(t)$ as the set of jobs with $p_{j}\leq p$ and release time $r_{j}\leq t$. We are interested in how NSJF and OPT handle the jobs in $J_{\leq p}(t)$, though we measure progress differently for each:
  • For OPT, let $V_{\leq p}(t)$ denote the total workload of jobs in $J_{\leq p}(t)$ that has been processed by time $t$.
  • For NSJF, let $V^{\prime}_{\leq p}(t)$ denote the total processing time of jobs in $J_{\leq p}(t)$ that have started under NSJF by time $t$; that is, the total processing time of job $j$ in the set $\{j\in J_{\leq p}(t)\mid s_{j}(\textup{{NSJF}})\leq t\}$.
Lemma 27.
For every $t\geq 0$, define $t^{\prime}=A^{-1}(m(t+\tau))$. Then for every upper bound $p$, we have: $V^{\prime}_{\leq p}(t^{\prime})\geq V_{\leq p}(t).$
Proof.
Fix an arbitrary $p$, let $t_{0}$ be the last time some machine is idle by NSJF before $t^{\prime}$. Define $s_{j}$ as the latest start time, within the interval $(t_{0},\,t+\tau)$, at which NSJF schedules a job with $p_{j}>p$. If no such job exists, we set $s_{j}=t_{0}$ and conceptually assume that an infinitely large job starts at time $s_{j}$.
Since NSJF starts job $j$ with $p_{j}>p$ at time $s_{j}$, it must have already started all jobs with size at most $p$ released up to $s_{j}$ (due to its SJF policy). Thus, we have:
$V^{\prime}_{\leq p}(s_{j})=\sum_{h\in J_{\leq p}(s_{j})}p_{h}\geq V_{\leq p}(s_{j}).$
We now analyze two cases:
Case 1: $s_{j}\geq t$. We straightforwardly have:
$V^{\prime}_{\leq p}(t^{\prime})\geq V^{\prime}_{\leq p}(s_{j})\geq V_{\leq p}(s_{j})\geq V_{\leq p}(t).$
Case 2: $s_{j}<t$. Then the time interval $s_{j}+\tau,t^{\prime})$ must be fully used by NSJF to process jobs with size at most $p$ (if any).
$V^{\prime}_{\leq p}(t^{\prime})$ $\geq V^{\prime}_{\leq p}(s_{j})+\int_{s_{j}+\tau}^{t^{\prime}}a(y)\,dy$
$\geq V_{\leq p}(s_{j})+A(t^{\prime})-A(s_{j}+\tau)$
$\geq V_{\leq p}(s_{j})+m(t+\tau)-m(s_{j}+\tau)$
$\geq V_{\leq p}(s_{j})+m(t-s_{j})$
$\geq V_{\leq p}(t).$
In both cases, we conclude:
$V^{\prime}_{\leq p}(t^{\prime})\geq V_{\leq p}(t),$
which proves the lemma. ∎
Denote $N_{\leq p}(t)$ (or $N^{\prime}_{\leq p}(t)$ below) as the number of jobs in $J_{\leq p}(t)$ completed by time $t$ by OPT (or NSJF, respectively). Note that without the subscription $\leq p$, we remove the condition on job size. We prove:
Lemma 28.
For every $t\geq 0$, define $t^{\prime}=A^{-1}(m(t+2\tau))$. we have $N^{\prime}_{\leq p}(t^{\prime})\geq N_{\leq p}(t)$.
Proof.
Define $t_{1}=A^{-1}(m(t+\tau))$ and let $t_{2}=t^{\prime}$. By Lemma 27, we have the volume bound: $V^{\prime}_{\leq p}(t_{1})\geq V_{\leq p}(t)$ for every pp.
We claim the following: the $i$-th smallest job (in terms of size) that has started (either in processing or completed) under NSJF by time $t_{1}$, denoted by size $p^{\prime}_{i}$, must not exceed the size of the $i$-th smallest completed job under OPT by time $t$, denoted $p_{i}$. That is, we aim to prove: $p^{\prime}_{i}\leq p_{i}$ for all ii.
To prove this, assume for contradiction that there exists the smallest index $i$ such that $p^{\prime}_{i}>p_{i}$. Under this assumption, we have the following bound for OPT:
$V_{\leq p_{i}}(t)\geq\sum_{j=1}^{i}p_{j},$
since at least $i$ jobs of size at most $p_{i}$ are completed by time $t$ in OPT.
Now consider NSJF. Since the $i$-th smallest started job under NSJF has a size strictly greater than $p_{i}$, all started jobs of size $\leq p_{i}$ must be among the first i1i-1 jobs. Therefore, we have:
$V^{\prime}_{\leq p_{i}}(t_{1})\leq\sum_{j=1}^{i-1}p^{\prime}_{j}\leq\sum_{j=1}^{i-1}p_{j},$
where the second inequality uses the assumption that $p^{\prime}_{j}\leq p_{j}$ for all $j<i$. Combining these two, we get: $V^{\prime}_{\leq p_{i}}(t_{1})<V_{\leq p_{i}}(t),$ which contradicts the earlier volume bound from Lemma 27. Hence, our assumption must be false, and we conclude: $p^{\prime}_{i}\leq p_{i}$ for all ii.
Finally, this implies that the number of started jobs with $p_{j}\leq p$ by time $t_{1}$ under NSJF is at least the number of completed jobs with $p_{j}\leq p$ by time $t$ under OPT. Moreover, since the number of active machines in NSJF is non-decreasing over time, all jobs started by time $t_{1}$ must be completed by time $t_{2}$. This completes the proof of the lemma.
Lemma 29.
Consider running NSJF on jobs with processing times bounded by $\tau$, even in the presence of an initial blocking vector $\vec{b}$ on machines, where the total blocking time is $B=\sum_{i=1}^{m}b_{i}$. Then, compared to an offline preemptive optimal solution, even with migration, and without any blocking, the flow time of NSJF is bounded by:
$F(\textup{{NSJF}}(\vec{b}))\leq F(\textup{{OPT}}(\vec{0}))+2n\tau+\frac{nB}{m}.$
Proof.
Let $u$ denote the maximum completion time under NSJF. The flow times of NSJF and OPT can be expressed as:
$F(\textup{{OPT}})=\int_{0}^{\infty}\left(|J(t)|-N(t)\right)\,dt,\quad F(\textup{{NSJF}})=\int_{0}^{u}\left(|J(t)|-N^{\prime}(t)\right)\,dt,$
where $|J(t)|$ is the number of jobs released by time $t$.
By Lemma 28, NSJF dominates OPT after a shifted time point, specifically after $A^{-1}(m(t+2\tau))$. To bound the shift, observe that for any $\theta$,
$A(\theta)=\int_{0}^{\theta}a(y)\,dy\geq m\theta-B,$
which implies:
$A^{-1}(m(t+2\tau))\leq t+2\tau+\frac{B}{m}.$
Using this, we bound the flow time of NSJF by shifting the completion curve by at most $2\tau+\frac{B}{m}$:
$F(\textup{{NSJF}})$ $=\int_{0}^{u}\left(|J(t)|-N^{\prime}(t)\right)\,dt$
$\leq\int_{0}^{u-2\tau-\frac{B}{m}}\left(|J(t)|-N^{\prime}(t+2\tau+\frac{B}{m})\right)\,dt+\int_{u-2\tau-\frac{B}{m}}^{u}|J(t)|\,dt$
$\leq\int_{0}^{u-2\tau-\frac{B}{m}}\left(|J(t)|-N(t)\right)\,dt+2n\tau+\frac{nB}{m}$
$\leq F(\textup{{OPT}})+2n\tau+\frac{nB}{m}.$
Appendix B Randomized Non-Preemptive Algorithm: Multiple Machines
In this section, we extend Algorithm 2 from single-machine case to multi-machine case. The intuition is already introduced. We detail the dynamic but online stable algorithm in Algorithm 10.
Include : Algorithm 1
Data: Every job’s current states as determined by Algorithm 1.
On job $i$ release:
   Classify job $i$ large or small based on Partition($i,\ell$) with $\ell=\lfloor\sqrt{nm}\rfloor$;
 if $i$ $i$ is classified as large then
      Sample $w_{i}\sim\mathrm{Unif}\{1,2,\dotsc,\lfloor\sqrt{n/m}\rfloor\}$ ;
      Sample $m_{i}\sim\mathrm{Unif}\{1,2,\dotsc,m\}$ ;
   end if
  Run NSJF on all small jobs to get schedule $\S_{1}$ ;
 // Including proxy jobs.
 $\S_{2}\leftarrow\S_{1}$ ;
 for each unproxied (active or committed) large job $j$ in order of release time $r_{j}$ do
    $t\leftarrow$ the first time in $\S_{2}$ where the cumulative idle time on machine $m_{j}$ since $r_{j}$ is at least $w_{j}p_{j}$;
      Insert $j$ into $\S_{2}$ to start at $t$ on $m_{j}$ ;
   end for
  Schedule jobs according to $\S_{2}$;
Algorithm 10 Dynamic (Online Stable) Non-Preemptive Randomized Algorithm for Multi-Machine
Theorem 10.
There exists an online polynomial-time randomized non-preemptive algorithm for total flow time minimization, e.g., Algorithm 10, that is $O(\sqrt{n/m})$-competitive against the preemptive offline solution.
Proof.
The algorithm is online simulatable and runs in polynomial time, for the same reason as in Lemma 5. We use $S$ and $L$ to denote the sets of small and large jobs, where $S$ also includes proxy jobs. We use $J$ to denote the real job set, which includes only actual jobs and excludes proxy jobs.
For small jobs, by Lemma 1, and Lemma 2, we have:
$F(\S_{1})=O\left(n\cdot\frac{1}{\sqrt{nm}}\right)\cdot F(\textup{{OPT}}(S))\leq O(\sqrt{n/m})\cdot\textup{{OPT}}.$
For large jobs, we bound both the external delay and the self-delay. The external delay $\delta_{j}$ of a large job $j$ (i.e., the flow time increase caused to earlier jobs by inserting $j$) satisfies $\mathbb{E}[\delta_{j}]=O(\sqrt{n/m})\cdot p_{j}$. The proof is as follows: Let $D_{k,l}$ be the set of jobs delayed by $j$ when $w_{j}=k$, $m_{j}=l$, where $k\in\{1,\dots,\lfloor\sqrt{n/m}\rfloor\}$ and $l\in\{1,\dots,m\}$. Since there is $p_{j}$ idle time between two possible adjacent insertion locations, all $D_{k,l}$ are disjoint. Then:
$\mathbb{E}[|D_{k,l}|]=\sum_{k=1}^{\lfloor\sqrt{n/m}\rfloor}\sum_{l=1}^{m}|D_{k,l}|\cdot\Pr[w_{j}=k]\cdot\Pr[m_{j}=l]\leq O(\sqrt{n/m}),$
so $\mathbb{E}[\delta_{j}]=O(\sqrt{n/m})\cdot p_{j}$.
For the self-delay, if $j$ is proxied by $j^{\prime}$, we count $r_{j^{\prime}}-r_{j}$; otherwise, we count $C^{\prime}_{j}-r_{j}$, where $C^{\prime}_{j}$ is the completion time of $j$ immediately after insertion. This self-delay is denoted by $\hat{F}_{j}$. Ignoring the final $p_{j}$ processing time, the delay term only comes from collecting idle time or passing through busy time.
Each job collects at most $\sqrt{n/m}\cdot p_{j}$ units of idle time, so the total contribution from this is:
$\sqrt{n/m}\cdot\sum_{j\in L}p_{j}\leq\sqrt{n/m}\cdot\textup{{OPT}}.$
For busy time, each unit of busy period can be charged to at most $\sqrt{nm}$ jobs (by Lemma 1), and the total busy time is at most $\textup{{OPT}}/m$, so (putting back $p_{j}$ for each $j$):
$\sum_{j\in L}\hat{F}_{j}\leq\sum_{j\in L}p_{j}+\sqrt{nm}\cdot\frac{\textup{{OPT}}}{m}=O(\sqrt{n/m})\cdot\textup{{OPT}}.$
Putting everything together:
$F(\S_{2})\leq F(\S_{1})+\sum_{j\in L}\mathbb{E}[\delta_{j}]+\sum_{j\in L}\hat{F}_{j}=O(\sqrt{n/m})\cdot\textup{{OPT}}.$

### v2

@@
   • The processing time of any job classified as small is bounded by $\textup{{OPT}}/\ell$.
-To achieve this, we maintain a set of active large jobs, denoted AA, containing the $\ell$ jobs with the largest processing times seen so far.
+To achieve this, we maintain a set of active large jobs, denoted $A$, containing the $\ell$ jobs with the largest processing times seen so far.
 The logic of our framework defines a clear life cycle for every job based on its interaction with this set.
@@
 If it is not large enough to enter the top-$\ell$ group, it is permanently classified as a small job.
-Otherwise, it is designated as an active large job and added to the set AA.
+Otherwise, it is designated as an active large job and added to the set $A$.
 Retirement and the Proxy Mechanism.
-The main complexity arises when a job $j\in A$ is forced out (or retired) at time $r_{i}$ due to the arrival of a new, larger job ii. The state of $j$ at this moment determines its new classification:
+The main complexity arises when a job $j\in A$ is forced out (or retired) at time $r_{i}$ due to the arrival of a new, larger job $i$.
+The state of $j$ at this moment determines its new classification:
   • If job $j$ is still waiting, it becomes proxied.
@@
 We call this a committed large job.
-Although no longer in the active set AA, it must still be tracked and scheduled as a large job.
+Although no longer in the active set $A$, it must still be tracked and scheduled as a large job.
 This is a subtle point in the analysis and can be safely ignored for a high-level understanding, but it is an important case required for the correctness of our online algorithm.
-Data: Active large job set AA, and every job’s other current states
+Data: Active large job set $A$, and every job’s other current states
 Function Partition(ii, $\ell$):
  if $|A|<\ell$ then
-      Add $i$ to AA;
+      Add $i$ to $A$;
     return $i$ is large
@@
       Let $j=\arg\min_{j\in A}p_{j}$;
-      Remove $j$ from AA and add $i$ to AA;
+      Remove $j$ from $A$ and add $i$ to $A$;
     if job $j$ is still waiting then
@@
 We now present the pseudocode for this partitioning step in Algorithm 1.
-For simplicity, the function only explicitly manages the active large job set AA; the other sets (small, proxied, and committed jobs) are implicitly defined by the procedure.
+For simplicity, the function only explicitly manages the active large job set $A$; the other sets (small, proxied, and committed jobs) are implicitly defined by the procedure.
 This partitioning mechanism guarantees our desired invariants along with a key monotonicity property: the set of all small jobs (including proxy jobs) forms an “easier” scheduling instance than the original job set.
@@
   1. 1.
-    The number of active large jobs (the set AA) is at most $\ell$ at any time.
+    The number of active large jobs (the set $A$) is at most $\ell$ at any time.
   2. 2.
@@
 Proof.
-We discuss two timepoints $t+\tau$ and $t+2\tau$. By Lemma 3, we have the volume bound: $V^{\prime}_{\leq p}(t+\tau)\geq V_{\leq p}(t)$ for every pp.
+We discuss two timepoints $t+\tau$ and $t+2\tau$. By Lemma 3, we have the volume bound: $V^{\prime}_{\leq p}(t+\tau)\geq V_{\leq p}(t)$ for every $p$.
 We claim the following: the $i$-th smallest job (in terms of size) that has started (either in processing or completed) under NSJF by time $t+\tau$, denoted by size $p^{\prime}_{i}$, must not exceed the size of the $i$-th smallest completed job under OPT by time $t$, denoted $p_{i}$.
-That is, we aim to prove: $p^{\prime}_{i}\leq p_{i}$ for all rank ii.
+That is, we aim to prove: $p^{\prime}_{i}\leq p_{i}$ for all rank $i$.
 To prove this, assume for contradiction that there exists the smallest rank $i$ such that $p^{\prime}_{i}>p_{i}$. Under this assumption, we have the following bound for OPT:
@@
 Combining these two, we get: $V^{\prime}_{\leq p_{i}}(t+\tau)<V_{\leq p_{i}}(t),$ which contradicts the earlier volume bound from Lemma 3.
-Hence, our assumption must be false, and we conclude: $p^{\prime}_{i}\leq p_{i}$ for all ii.
+Hence, our assumption must be false, and we conclude: $p^{\prime}_{i}\leq p_{i}$ for all $i$.
 Finally, this implies that the number of started jobs with $p_{j}\leq p$ by time $t+\tau$ under NSJF is at least the number of completed jobs with $p_{j}\leq p$ by time $t$ under OPT.
@@
   • Machine 1: Reserved for large jobs.
-  • Mixed machines (Machines 22 to $\lceil m/2\rceil+1$): Accept both small and large jobs.
+  • Mixed machines (Machines $2$ to $\lceil m/2\rceil+1$): Accept both small and large jobs.
   • Small-Only machines (Machines $\lceil m/2\rceil+2$ to $m$): Restricted to small jobs only.
@@
 These changes are mild and preserve all the original properties of the online partitioning framework.
-Data: Active large job set AA, and every job’s other current states
+Data: Active large job set $A$, and every job’s other current states
 Function Partition($i,\ell$):
@@
  else if $|A|<\ell$ then
-      Add $i$ to AA;
+      Add $i$ to $A$;
     return $i$ is large
@@
       Let $j=\arg\min_{j\in A}p_{j}$;
-      Remove $j$ from AA and add $i$ to AA;
+      Remove $j$ from $A$ and add $i$ to $A$;
     if job $j$ is still waiting then
@@
 After introducing the fine-grained partitioning framework, we now describe the basic design of our scheduling policy, which divides the $m$ machines into two roles:
-  • Small-Only machines (Machines 11 to $\lfloor m/2\rfloor$): These machines accept only small jobs.
+  • Small-Only machines (Machines $1$ to $\lfloor m/2\rfloor$): These machines accept only small jobs.
   • Mixed machines (Machines $\lfloor m/2\rfloor+1$ to $m$): These machines can process both small and large jobs.
@@
 In this section, we extend the $\sqrt{n/m}$-competitive algorithm (which assumes the prior knowledge of $n$) to an $O(n^{\alpha}/\sqrt{m})$-competitive algorithm without such knowledge, where $\alpha=(\sqrt{5}-1)/2$.
-The overall framework still follows Section 6: we keep the same machine partition (Small-Only machines 11 to $\lfloor m/2\rfloor$ and Mixed machines $\lfloor m/2\rfloor+1$ to $m$), and we still kill a large job once enough small jobs are blocked.
+The overall framework still follows Section 6: we keep the same machine partition (Small-Only machines $1$ to $\lfloor m/2\rfloor$ and Mixed machines $\lfloor m/2\rfloor+1$ to $m$), and we still kill a large job once enough small jobs are blocked.
 The key difference is that the total number of jobs $n$ is unknown in advance.
@@
 As a result, a job may switch between being large and small multiple times.
-For instance, when $n(t)$ is small, we may treat the top 33 jobs as large, so a job of rank 44 is small.
-Later, as more jobs arrive and $\ell(t)$ becomes 44, this job becomes large.
-If even more jobs arrive, $\ell(t)$ may increase to 55, but that same job may drop out of the top-55 and become small again.
+For instance, when $n(t)$ is small, we may treat the top $3$ jobs as large, so a job of rank $4$ is small.
+Later, as more jobs arrive and $\ell(t)$ becomes $4$, this job becomes large.
+If even more jobs arrive, $\ell(t)$ may increase to $5$, but that same job may drop out of the top-$5$ and become small again.
 This uncontrolled oscillation between the large and small states makes the partition unstable and invalidates the previous analysis.
@@
 We use a partition rule that depends on the number of jobs released so far.
-Instead of using $n(t)$ directly, we let $N=2^{k}$ be the smallest power of two such that $N\geq n(t)$, and use NN to define the large–small threshold $\ell$ at every time $t$, where $\ell=\lfloor N^{\alpha}\sqrt{m}\rfloor$.
+Instead of using $n(t)$ directly, we let $N=2^{k}$ be the smallest power of two such that $N\geq n(t)$, and use $N$ to define the large–small threshold $\ell$ at every time $t$, where $\ell=\lfloor N^{\alpha}\sqrt{m}\rfloor$.
 Let $T_{j}(t)$ denote the type of job $j$ at time $t$.
@@
 Otherwise, $T_{j}(t)$ is classified as small.
-A New Kill Event When NN Is Updated.
+A New Kill Event When $N$ Is Updated.
 Previously, we triggered a kill event only when too many small jobs were waiting while a large job was being processed on a Mixed machine.
 However, in this setting, a large job may be processed on a Small-Only machine, since it might have been classified as small earlier.
-To prevent this from causing inefficiency, we introduce a new kill event that occurs whenever NN is updated (that is, when $n(t)>N$, and we double NN to $2N$).
+To prevent this from causing inefficiency, we introduce a new kill event that occurs whenever $N$ is updated (that is, when $n(t)>N$, and we double $N$ to $2N$).
 At this moment, we check whether any currently large jobs are running on the Small-Only machines and kill all of them if so.
@@
    end if
- // Update the Partition of all jobs again based on the current number of jobs NN and the current sum of processing time $P$.
+ // Update the Partition of all jobs again based on the current number of jobs $N$ and the current sum of processing time $P$.
  $\ell=\lfloor N^{\alpha}\sqrt{m}\rfloor$ where $\alpha=(\sqrt{5}-1)/2$ ;
 […]
- if NN is updated by the release of job $j$ then
+ if $N$ is updated by the release of job $j$ then
       Kill all currently large jobs processing on Small-Only machines ;
@@
 Note that we have two types of kill events.
-One occurs when NN is updated, which we call a type-A event; the other occurs when $\phi$ becomes large, which we call a type-B event.
-The number of type-B kill events can be bounded similarly to the known-$n$ case, with the help of the rounded benchmark value NN.
+One occurs when $N$ is updated, which we call a type-A event; the other occurs when $\phi$ becomes large, which we call a type-B event.
+The number of type-B kill events can be bounded similarly to the known-$n$ case, with the help of the rounded benchmark value $N$.
 Lemma 15.
@@
 Proof.
-We know that if a type-B kill event occurs at time $t$, there must be at least $\lfloor N(t)^{\alpha}\sqrt{m}\rfloor$ small jobs waiting, where $N(t)$ denotes the value of NN at time $t$.
+We know that if a type-B kill event occurs at time $t$, there must be at least $\lfloor N(t)^{\alpha}\sqrt{m}\rfloor$ small jobs waiting, where $N(t)$ denotes the value of $N$ at time $t$.
 We interpret these small jobs as the creators of this kill event.
-Note that when a small job $j$ creates a type-B kill event at time $t$, it remains small until NN is updated.
-Hence, after the kill event at $t$ and before the next update of NN, no large job will be processed on any machine unless we start job $j$.
-On the other hand, if we start $j$, it will not be killed again until NN is updated, since $j$ stays small during this period.
-Therefore, job $j$ cannot trigger another type-B kill event before NN changes.
+Note that when a small job $j$ creates a type-B kill event at time $t$, it remains small until $N$ is updated.
+Hence, after the kill event at $t$ and before the next update of $N$, no large job will be processed on any machine unless we start job $j$.
+On the other hand, if we start $j$, it will not be killed again until $N$ is updated, since $j$ stays small during this period.
+Therefore, job $j$ cannot trigger another type-B kill event before $N$ changes.
 Consequently, the number of kill events that occur while $N=x$ is at most $x^{1-\alpha}/\sqrt{m}$.
-Since NN increases geometrically by a factor of 22, the total number of kill events is bounded by $O\!\left(\frac{n^{1-\alpha}}{\sqrt{m}}\right),$ as $N\leq 2n$.
+Since $N$ increases geometrically by a factor of $2$, the total number of kill events is bounded by $O\!\left(\frac{n^{1-\alpha}}{\sqrt{m}}\right),$ as $N\leq 2n$.
 ∎
@@
 Therefore, its size is bounded by $\frac{\textup{{OPT}}}{N(t)^{\alpha}\sqrt{m}}$.
-Although NN may be smaller than $n$ since $n$ is unknown at that time, it is lower bounded by $n^{\alpha}\sqrt{m}/2$, because at least $n^{\alpha}\sqrt{m}$ small blocked jobs are waiting, implying $n(t)\geq n^{\alpha}\sqrt{m}$.
+Although $N$ may be smaller than $n$ since $n$ is unknown at that time, it is lower bounded by $n^{\alpha}\sqrt{m}/2$, because at least $n^{\alpha}\sqrt{m}$ small blocked jobs are waiting, implying $n(t)\geq n^{\alpha}\sqrt{m}$.
 Hence, the size of $j$ satisfies
@@
 If a blocked job is waiting, it is in its last small period, so it is small and all machines are busy at that moment.
-Hence, if some Small-Only machines are idle at the end, a type-A kill event must have occurred when NN was updated to $2N$ — the only time a small job can become large.
-At such an update, at most NN jobs have been released, so there are at most NN blocked jobs.
+Hence, if some Small-Only machines are idle at the end, a type-A kill event must have occurred when $N$ was updated to $2N$ — the only time a small job can become large.
+At such an update, at most $N$ jobs have been released, so there are at most $N$ blocked jobs.
 Each killed job was small just before the update, hence its size is at most $\textup{{OPT}}/(N^{\alpha}\sqrt{m})$.
@@
 $O\!\left(\frac{\textup{{OPT}}\cdot N^{1-\alpha}}{\sqrt{m}}\right).$
-As NN doubles through $1,2,4,\ldots$ up to $n<2^{k}\leq 2n$, the total contribution sums to
+As $N$ doubles through $1,2,4,\ldots$ up to $n<2^{k}\leq 2n$, the total contribution sums to
 $O\!\left(\frac{n^{1-\alpha}}{\sqrt{m}}\right)\textup{{OPT}}.$
@@
 We start with the lower bound that any randomized non-preemptive online algorithm for a single machine is $\Omega(\sqrt{n})$ as a warm-up.
-The counterexample is constructed as Algorithm 6, which is a distribution of two possibilities: the only difference is whether the first $k$ $\varepsilon$-jobs are released at time 11 or 22.
+The counterexample is constructed as Algorithm 6, which is a distribution of two possibilities: the only difference is whether the first $k$ $\varepsilon$-jobs are released at time $1$ or $2$.
 Input: Number of jobs $n$
 $k\leftarrow\lfloor\sqrt{n-2}\rfloor,\varepsilon\leftarrow 0.5n^{-2}$;
-Release job 11 at time 0 with processing time 22 ;
+Release job $1$ at time 0 with processing time $2$ ;
 Flip an independent fair coin: $r\sim\text{Bernoulli}(1/2)$ ;
 Release $k$ jobs at time $1+r$ with processing time $\varepsilon$ ;
-// released at 11 or 22
+// released at $1$ or 22
 for $t=2$ to $k$ do
@@
 We analyze the counterexample described in Algorithm 6.
-For each of the two possibilities, the optimal algorithm can arrange job 11 at time 0 or 11 correctly and delay each $\varepsilon$-job by at most $O(k^{2}\varepsilon)$.
+For each of the two possibilities, the optimal algorithm can arrange job $1$ at time 0 or $1$ correctly and delay each $\varepsilon$-job by at most $O(k^{2}\varepsilon)$.
 Therefore, the optimal flow time is always $O(1)$.
-However, for any deterministic online algorithm, it has to decide whether to start job 11 at the beginning or not:
-  • If the algorithm starts job 11 at time 0, then with probability $1/2$, the algorithm will delay the first $k$ $\varepsilon$-jobs by 11 for the case the randomized $\varepsilon$-jobs are released at time 11;
-  • Otherwise, the algorithm cannot start job 11 before time $k+1$ for the case the randomized $\varepsilon$-jobs are released at time 22, or it will delay at least $k$ $\varepsilon$-jobs by 11.
+However, for any deterministic online algorithm, it has to decide whether to start job $1$ at the beginning or not:
+  • If the algorithm starts job $1$ at time 0, then with probability $1/2$, the algorithm will delay the first $k$ $\varepsilon$-jobs by $1$ for the case the randomized $\varepsilon$-jobs are released at time $1$;
+  • Otherwise, the algorithm cannot start job $1$ before time $k+1$ for the case the randomized $\varepsilon$-jobs are released at time $2$, or it will delay at least $k$ $\varepsilon$-jobs by $1$.
 Therefore, the optimal deterministic algorithm has an expected flow time of $\Omega(\sqrt{n})$ against the adversary, and hence the randomized algorithm has the lower bound $\Omega(\sqrt{n})$ by Yao’s minimax theorem.
@@
 Each gadget, parameterized by a start time $t$, contains $2k+1$ jobs structured as follows and is illustrated in Figure 1:
-Job Type Indices ($j$) $r_{j}$ $p_{j}$
+Job Type Indices ($i$) $r_{i}$ $p_{i}$
   Large 0 0 0 22
   Fixed $\{1,\dots,k\}$ 4+(i1)/k4+(i-1)/{k} $1/k$
-  Left (when $c=0$) $\{k+2,\dots,2k+1\}$ 1+(j(k+2))/k1+(j-(k+2))/{k} $1/k$
-  Right (when $c=1$) $\{k+2,\dots,2k+1\}$ 2+(j(k+2))/k2+(j-(k+2))/{k} $1/k$
+  Left (when $c=0$) $\{k+2,\dots,2k+1\}$ 1+(i(k+2))/k1+(i-(k+2))/{k} $1/k$
+  Right (when $c=1$) $\{k+2,\dots,2k+1\}$ 2+(i(k+2))/k2+(i-(k+2))/{k} $1/k$
 Figure 1: A gadget of $2k+1$ jobs defined in Algorithm 7, with two random options.
-  • A Large Job: Job 0, with processing time 22 and release time $t$.
+  • A Large Job: Job 0, with processing time $2$ and release time $t$.
   • Fixed Small Jobs: $k$ jobs, each with processing time $1/k$, released sequentially in the interval $[t+4,t+5)$.
@@
 In each busy gadget, we can schedule the large job either at $5b+0$ or $5b+2$ depending on $c$, and every small job upon arrival.
-As a result, the total flow time in this gadget is at most 44 (the large job) +1+1 (fixed small jobs) +1+1 (random small jobs).
+As a result, the total flow time in this gadget is at most $4$ (the large job) +1+1 (fixed small jobs) +1+1 (random small jobs).
 Therefore $\textup{{OPT}}\leq 6mk$ by summing up all gadgets.
@@
 Definition 1.
-In a given batch, the set of mkmk jobs released in a time interval $[t,t+1)$ constitutes a small job period.
+In a given batch, the set of $mk$ jobs released in a time interval $[t,t+1)$ constitutes a small job period.
 A large job conflicts with this period if its execution starts within the interval [t1.5,t+0.5][t-1.5,t+0.5].
@@
 The number of small jobs that can be completed by time $u$ is at most $\lfloor k\int_{0}^{u}g(y)dy\rfloor$.
-The total flow time of the kmkm small jobs in the batch is no less than
+The total flow time of the $km$ small jobs in the batch is no less than
 $F$ $\geq\int_{0}^{1}f(u)-\lfloor k\int_{0}^{u}g(y)\mathrm{d}y\rfloor\,\mathrm{d}u$
@@
 If no large jobs conflict with this period, $g(u)$ is defined as $m$ for every $u\in[0,1)$.
-By the condition of the lemma, we have $x$ jobs that decrease $g(u)$ by 11 with a duration of at least $0.5$.
+By the condition of the lemma, we have $x$ jobs that decrease $g(u)$ by $1$ with a duration of at least $0.5$.
 The minimum contribution of the lower bound occurs when each large job conflicts the period at $[0.5,1)$, yielding
@@
 $c_{b}$) representing the number of large jobs scheduled in this batch that conflict with small job periods.
-Note that before time 11, the algorithm does not know the realization of $c_{b}$ this round. Therefore, its behavior before time 11 is not related to the result of $c_{b}$.
-  • Let $x$ be the number of large jobs started in [0.5, 1)[-0.5,\;1). If small jobs are Left, all these $x$ jobs contribute to II.
-After time 11, the algorithm’s scheduling decisions may depend on $c_{b}$. We consider the two cases separately:
+Note that before time $1$, the algorithm does not know the realization of $c_{b}$ this round. Therefore, its behavior before time $1$ is not related to the result of $c_{b}$.
+  • Let $x$ be the number of large jobs started in [0.5, 1)[-0.5,\;1). If small jobs are Left, all these $x$ jobs contribute to $I$.
+After time $1$, the algorithm’s scheduling decisions may depend on $c_{b}$. We consider the two cases separately:
   • Right case (w.p.
@@
 However, in the current construction, the algorithm must decide which large job to start in advance; a wrong decision can lead to a substantial cost.
-Job Type Indices (ii) $r_{i}$ $p_{i}$
+Job Type Indices ($i$) $r_{i}$ $p_{i}$
   Large 0 0 0 22
@@
 Proof.
-In each busy gadget, if $c=0$, then schedule large job 0 at 0 and large job 11 at 33; if $c=1$, then schedule large job 11 at 0 and large job 0 at 44.
+In each busy gadget, if $c=0$, then schedule large job 0 at 0 and large job $1$ at $3$; if $c=1$, then schedule large job $1$ at 0 and large job 0 at $4$.
 In this way, no large jobs intersect with any small jobs.
-As a result, the total flow time in this gadget is at most 99 (large jobs) +1+1 (fixed small jobs) +1+1 (random small jobs).
+As a result, the total flow time in this gadget is at most $9$ (large jobs) +1+1 (fixed small jobs) +1+1 (random small jobs).
 Therefore, $\textup{{OPT}}\leq 11mk$ by summing up all gadgets.
@@
 Let $N_{b}=N_{b,2}+N_{b,3},I_{b}=I_{b,2}+I_{b,3}$.
-Note that before time 22, the algorithm does not know the realization of $c_{b}$ this round. Therefore, its behavior before time 22 is not related to the result of $c_{b}$.
-  • Let $x_{2}$ be the number of jobs with processing time 22 started in [0.5, 2)[-0.5,\;2).
-  • Let $x_{3}$ be the number of jobs with processing time 33 started in [0.5, 2)[-0.5,\;2).
-After time 22, the algorithm scheduling decisions may depend on $c_{b}$. We consider the two cases separately:
+Note that before time $2$, the algorithm does not know the realization of $c_{b}$ this round. Therefore, its behavior before time $2$ is not related to the result of $c_{b}$.
+  • Let $x_{2}$ be the number of jobs with processing time $2$ started in [0.5, 2)[-0.5,\;2).
+  • Let $x_{3}$ be the number of jobs with processing time $3$ started in [0.5, 2)[-0.5,\;2).
+After time $2$, the algorithm scheduling decisions may depend on $c_{b}$. We consider the two cases separately:
   • Left case (w.p.
 $0.5$): None of the $x_{3}$ jobs can finish without intersection, and at most $x_{2}^{(L)}=\min\{m,x_{2}\}$ out of the $x_{2}$ jobs can finish without intersection.
-Let $y_{2}^{(L)}$ be the number of jobs with processing time 22 starting in $[2,6.5)$ that finish without intersection, and $y_{3}^{(L)}$ be the number for processing time 33.
+Let $y_{2}^{(L)}$ be the number of jobs with processing time $2$ starting in $[2,6.5)$ that finish without intersection, and $y_{3}^{(L)}$ be the number for processing time $3$.
 At most $m$ jobs out of $y_{2}^{(L)}+y_{3}^{(L)}$ can finish without intersection.
@@
 $0.5$): At most $m$ out of these $x_{2}+x_{3}$ jobs could end up without intersection with small jobs; let $x_{2}^{(R)}$ and $x_{3}^{(R)}$ be the number of such jobs, respectively.
-At most $m$ jobs with processing time 22 released in $[2,6.5)$ can finish without intersection; let $y_{2}^{(R)}$ be the number of such jobs.
-None of the jobs with processing time 33 can finish without intersection.
+At most $m$ jobs with processing time $2$ released in $[2,6.5)$ can finish without intersection; let $y_{2}^{(R)}$ be the number of such jobs.
+None of the jobs with processing time $3$ can finish without intersection.
 Then,
@@
 and hence $\mathbb{E}(N_{b}-I_{b})+\mathbb{E}(N_{b,3}-I_{b,3})\leq\frac{5}{2}m$ since it doesn’t rely on the specific values of $c_{0},\ldots,c_{b-1}$.
-Consider the flow time by the large jobs. If we consider the jobs with processing time 22 or 33 and their intersections only, we have
+Consider the flow time by the large jobs. If we consider the jobs with processing time $2$ or $3$ and their intersections only, we have
 $\mathbb{E}(F^{(2,3)})$ $\geq\sum_{b=0}^{k-1}\sum_{i=0}^{b}(2m-E(N_{b}))+\sum_{b=0}^{k-1}E(I_{b})\cdot k/8$
@@
 $\geq\sum_{b=0}^{\lfloor\frac{7}{8}k\rfloor}\left(2m-E(N_{b}-I_{b})\right)\cdot k/8.$
-Similarly, consider the jobs with processing time 33 and their intersections only, we have
+Similarly, consider the jobs with processing time $3$ and their intersections only, we have
 $\mathbb{E}(F^{(3)})$ $\geq\sum_{b=0}^{\lfloor\frac{7}{8}k\rfloor}(m-\mathbb{E}(N_{b,3}-I_{b,3}))\cdot k/8.$
@@
 Phase 1: Construct an unsolved job
-   Release job 11 at 0 with processing time 44 ;
-   Release job 22 at 22 with processing time 11 ;
- if ALG is working on job 11 before 33 then
-      Release $c/2\cdot n$ jobs at 33 with processing time $\varepsilon$ ;
-      Release $c/2\cdot n$ jobs at 77 with processing time $\varepsilon$ ;
+   Release job $1$ at 0 with processing time $4$ ;
+   Release job $2$ at $2$ with processing time $1$ ;
+ if ALG is working on job $1$ before 33 then
+      Release $c/2\cdot n$ jobs at $3$ with processing time $\varepsilon$ ;
+      Release $c/2\cdot n$ jobs at $7$ with processing time $\varepsilon$ ;
     $t^{*}\leftarrow 7$;
  else
-      Release $c\cdot n$ jobs at 55 with processing time $\varepsilon$ ;
+      Release $c\cdot n$ jobs at $5$ with processing time $\varepsilon$ ;
     $t^{*}\leftarrow 5$;
@@
 Lemma 24.
-Against the adversary in Phase 11 of Algorithm 9, any deterministic online algorithm either incurs a flow time of $\Omega(n)$ or proceeds to Phase 2 with at least one of job 1 or job 2 unstarted.
+Against the adversary in Phase $1$ of Algorithm 9, any deterministic online algorithm either incurs a flow time of $\Omega(n)$ or proceeds to Phase 2 with at least one of job 1 or job 2 unstarted.
 In all cases, OPT remains $O(1)$ and all jobs are solved.
@@
   1. 1.
-    If ALG is working on job 11 at time 33, then OPT could solve job 22 by 33 and job 11 by $7+c/2\cdot n\varepsilon$. For ALG, however:
+    If ALG is working on job $1$ at time $3$, then OPT could solve job $2$ by $3$ and job $1$ by $7+c/2\cdot n\varepsilon$. For ALG, however:
     1. (a)
-      If ALG keeps working on job 11 until finished, the flow time is at least $c/2\cdot n$;
+      If ALG keeps working on job $1$ until finished, the flow time is at least $c/2\cdot n$;
     2. (b)
-      If ALG kills at time 33 to solve $\varepsilon$-jobs, switch to job 22 and then switch to job 11:
+      If ALG kills at time $3$ to solve $\varepsilon$-jobs, switch to job $2$ and then switch to job $1$:
       1. i.
-        If it doesn’t switch to $\varepsilon$-jobs at time 77, the flow time is at least $c/2\cdot n$;
+        If it doesn’t switch to $\varepsilon$-jobs at time $7$, the flow time is at least $c/2\cdot n$;
       2. ii.
-        Otherwise ALG must switch to $\varepsilon$-jobs at time 77, leaving job 11 unsolved.
+        Otherwise ALG must switch to $\varepsilon$-jobs at time $7$, leaving job $1$ unsolved.
   2. 2.
-    Otherwise, ALG cannot finish job $1,2$ by time 55 while OPT can. In this case ALG would switch to job 22 at time 22 and then to solve job 11, or it would fall back to case 1.
+    Otherwise, ALG cannot finish job $1,2$ by time $5$ while OPT can. In this case ALG would switch to job $2$ at time $2$ and then to solve job $1$, or it would fall back to case 1.
     1. (a)
-      If ALG keeps working on job 11 until finished, the flow time is at least $2c\cdot n$;
+      If ALG keeps working on job $1$ until finished, the flow time is at least $2c\cdot n$;
     2. (b)
-      Otherwise ALG must switch to $\varepsilon$-jobs at time 55, leaving job 11 unsolved.
+      Otherwise ALG must switch to $\varepsilon$-jobs at time $5$, leaving job $1$ unsolved.
 Lemma 25.
@@
 This strategy will result in some number of restarts of the large job $u$, say $r\geq 0$.
-These restarts partition the NN small jobs into $r+1$ groups.
+These restarts partition the $N$ small jobs into $r+1$ groups.
 Let $a_{i}$ be the number of $\varepsilon$-jobs released during the $i$-th run of job $u$ (for $i=1,\dots,r+1$), where a run is a period of continuous work on $u$.
@@
 Consider the following case for any deterministic algorithm ALG:
-  • A job with processing time 11 is released at time 0.
+  • A job with processing time $1$ is released at time 0.
   • Once ALG starts a job at $t$, the adversary releases $n$ jobs with processing time $1/n$ in $n/m$ batches, where each batch contains $m$ jobs released at time $t+1/n,t+2/n,\dots,t+1/m$ respectively.
-OPT can start all small jobs at their release time immediately; it can also solve the first job at 0 if $t>1$, or solve the first job at 22 otherwise.
+OPT can start all small jobs at their release time immediately; it can also solve the first job at 0 if $t>1$, or solve the first job at $2$ otherwise.
 Hence, OPT is $O(1)$ in both cases.
@@
 $\frac{\mathbb{E}[\textup{{ALG}}]}{\textup{{OPT}}}\leq C\cdot n,\quad\text{for all }n\geq n_{0}.$
-We now prove a contradiction: for any randomized algorithm ALG, there exists a constant $C>0$ such that for every positive integer $n_{0}$, one can construct an instance II with $|I|=n\geq n_{0}$ satisfying
+We now prove a contradiction: for any randomized algorithm ALG, there exists a constant $C>0$ such that for every positive integer $n_{0}$, one can construct an instance $I$ with $|I|=n\geq n_{0}$ satisfying
 $\frac{\mathbb{E}[\textup{{ALG}}]}{\textup{{OPT}}}>C\cdot n.$
@@
 We construct an oblivious adversary as follows.
-At time 0, the adversary releases a single job of size 22.
+At time 0, the adversary releases a single job of size $2$.
 For any algorithm, let $p(t)$ denote the probability density function that ALG starts this job at time $t$ if only the single job is present, which is known by the oblivious adversary since the probability distribution solely depend on the description of the algorithm and does not depend on algorithm’s randomized results.
@@
     In this case, the adversary releases $n_{1}$ dummy jobs of size 0 at time $t+1$.
-With probability at least $y$, ALG eventually starts the size-22 job within $[t,t+1)$, all these $n_{1}$ jobs complete no earlier than time $t+2$, so each contributes at least 11 to flow time and the algorithm’s total flow time is at least $n_{1}$, while an optimal solution completes all jobs by time 44, i.e., $\textup{{OPT}}\leq 4$.
+With probability at least $y$, ALG eventually starts the size-$2$ job within $[t,t+1)$, all these $n_{1}$ jobs complete no earlier than time $t+2$, so each contributes at least $1$ to flow time and the algorithm’s total flow time is at least $n_{1}$, while an optimal solution completes all jobs by time $4$, i.e., $\textup{{OPT}}\leq 4$.
 Therefore,
@@
     For all $t\in\{1,2,\dots,x\}$, it holds that $\int_{t}^{t+1}p(t)\,dt<y$.
-In this case, the probability that the size-22 job has not been started by time $x$ is at least 1xy1-xy.
+In this case, the probability that the size-$2$ job has not been started by time $x$ is at least 1xy1-xy.
 The adversary releases $n_{0}$ dummy jobs of size 0 at time $x+1$.
-If ALG delays the size-22 job until after time $x$, each dummy job has flow time at least $x$, while $\textup{{OPT}}\leq 2$.
+If ALG delays the size-$2$ job until after time $x$, each dummy job has flow time at least $x$, while $\textup{{OPT}}\leq 2$.
 Therefore,
@@
 Let $x=n_{0}$ and $y=\frac{1}{2n_{0}}$, so that $xy=\frac{1}{2}$.
-Then, in the Type A case, there exists an instance II with $|I|=n\gg n_{0}$ such that $\mathbb{E}[\textup{{ALG}}]/\textup{{OPT}}>(1/(8n_{0}))\cdot n$; in the Type B case, there exists an instance II with $|I|=n=n_{0}+1>n_{0}$ such that $\mathbb{E}[\textup{{ALG}}]/\textup{{OPT}}>(1/8)\cdot n$.
+Then, in the Type A case, there exists an instance $I$ with $|I|=n\gg n_{0}$ such that $\mathbb{E}[\textup{{ALG}}]/\textup{{OPT}}>(1/(8n_{0}))\cdot n$; in the Type B case, there exists an instance $I$ with $|I|=n=n_{0}+1>n_{0}$ such that $\mathbb{E}[\textup{{ALG}}]/\textup{{OPT}}>(1/8)\cdot n$.
 Therefore, for every constant $0<C\leq 1/8$, there exists a corresponding integer $n_{0}=1/(8C)$ such that
@@
 Proof.
-Define $t_{1}=A^{-1}(m(t+\tau))$ and let $t_{2}=t^{\prime}$. By Lemma 27, we have the volume bound: $V^{\prime}_{\leq p}(t_{1})\geq V_{\leq p}(t)$ for every pp.
+Define $t_{1}=A^{-1}(m(t+\tau))$ and let $t_{2}=t^{\prime}$. By Lemma 27, we have the volume bound: $V^{\prime}_{\leq p}(t_{1})\geq V_{\leq p}(t)$ for every $p$.
 We claim the following: the $i$-th smallest job (in terms of size) that has started (either in processing or completed) under NSJF by time $t_{1}$, denoted by size $p^{\prime}_{i}$, must not exceed the size of the $i$-th smallest completed job under OPT by time $t$, denoted $p_{i}$.
-That is, we aim to prove: $p^{\prime}_{i}\leq p_{i}$ for all ii.
+That is, we aim to prove: $p^{\prime}_{i}\leq p_{i}$ for all $i$.
 To prove this, assume for contradiction that there exists the smallest index $i$ such that $p^{\prime}_{i}>p_{i}$. Under this assumption, we have the following bound for OPT:
@@
 Combining these two, we get: $V^{\prime}_{\leq p_{i}}(t_{1})<V_{\leq p_{i}}(t),$ which contradicts the earlier volume bound from Lemma 27.
-Hence, our assumption must be false, and we conclude: $p^{\prime}_{i}\leq p_{i}$ for all ii.
+Hence, our assumption must be false, and we conclude: $p^{\prime}_{i}\leq p_{i}$ for all $i$.
 Finally, this implies that the number of started jobs with $p_{j}\leq p$ by time $t_{1}$ under NSJF is at least the number of completed jobs with $p_{j}\leq p$ by time $t$ under OPT.

### v3

@@
 Abstract
-This paper studies the classical online scheduling problem of minimizing total flow time for $n$ jobs on $m$ identical machines.
-Prior work often cites the $\Omega(n)$ lower bound for non-preemptive algorithms to argue for the necessity of preemption or resource augmentation, which shows the trivial $O(n)$-competitive greedy algorithm is tight.
-However, this lower bound applies only to deterministic algorithms in the single-machine case, leaving several fundamental questions unanswered.
-Can randomness help in the non-preemptive setting, and what is the optimal online deterministic algorithm when $m\geq 2$?
-We resolve both questions.
-We present a polynomial-time randomized algorithm with competitive ratio $\Theta(\sqrt{n/m})$ and prove a matching randomized lower bound, settling the randomized non-preemptive setting for every $m$.
-This also improves the best-known offline approximation ratio from $O(\sqrt{n/m}\log(n/m))$ to $O(\sqrt{n/m})$.
-On the deterministic side, we present a non-preemptive algorithm with competitive ratio $O(n/m^{2}+\sqrt{n/m}\log m)$ and prove a nearly matching lower bound.
-Our framework also extends to the kill-and-restart model, where we reveal a sharp transition of deterministic algorithms: we design an asymptotically optimal algorithm with the competitive ratio $O(\sqrt{n/m})$ for $m\geq 2$, yet establish a strong $\Omega(n/\log n)$ lower bound for $m=1$.
-Moreover, we show that randomization provides no further advantage, as the lower bound coincides with that of the non-preemptive setting.
-While our main results assume prior knowledge of $n$, we also investigate the setting where $n$ is unknown.
-We show kill-and-restart is powerful enough to break the $O(n)$ barrier for $m\geq 2$ even without knowing $n$.
-Conversely, we prove randomization alone is insufficient, as no algorithm can achieve an $o(n)$ competitive ratio in this setting.
+This paper studies the online scheduling problem of minimizing total flow time for $n$ jobs on $m$ identical machines.
+A classical $\Omega(n)$ lower bound shows that no deterministic single-machine algorithm can beat the trivial greedy, even when $n$ is known in advance.
+However, this barrier is specific to deterministic algorithms on a single machine, leaving open what randomization, multiple machines, or the kill-and-restart capability can achieve.
+We give a nearly complete answer.
+For randomized non-preemptive algorithms, we establish a tight $\Theta(\sqrt{n/m})$ competitive ratio, which also improves the best offline approximation to $O(\sqrt{n/m})$.
+For deterministic non-preemptive algorithms on multiple machines, we prove an $O(n/m^{2}+\sqrt{n/m}\log m)$ upper bound and an $\Omega(n/m^{2}+\sqrt{n/m})$ lower bound.
+In the kill-and-restart model, we reveal a sharp transition for deterministic algorithms: $\Omega(n/\log n)$ for $m=1$ versus $\Theta(\sqrt{n/m})$ for $m\geq 2$; the latter matches the optimal randomized ratio, and we further show that randomization provides no additional power in this model.
+We also investigate the setting where $n$ is unknown.
+We prove that no randomized non-preemptive algorithm achieves $o(n)$ on one machine or $o(n/m^{2}+\sqrt{n/m})$ on $m$ machines.
+In contrast, our kill-and-restart algorithm achieves $O(n^{\alpha}/\sqrt{m})$ for $m\geq 2$, where $\alpha=(\sqrt{5}-1)/2$, breaking the trivial bound without knowledge of $n$.
 1 Introduction
-The scheduling problem is a core combinatorial optimization problem across various disciplines in computer science.
-Its significance stems not only from its wide range of applications but also from its simple yet representative formulation.
-Numerous variants of scheduling problems exist, motivated by diverse real-world scenarios, such as different machine types (identical, related, or unrelated) and varying objectives (makespan, completion time, flow time).
-A prominent line of research in scheduling focuses on making irrevocable decisions for jobs that arrive online over time — a framework known as online scheduling, which was initiated in the 1960s by graham1966bounds.
-Consider an online service platform that must dispatch jobs submitted by multiple users as they arrive; this naturally leads to the objective of minimizing the total flow time, the sum over jobs of their completion time minus release time.
+Scheduling is one of the most fundamental problems in combinatorial optimization, with deep roots in both theory and practice.
+A rich landscape of scheduling variants spans different machine environments (identical, related, and unrelated [8, 24, 13]) and optimization objectives (makespan [15], completion time [18], and flow time [2, 23, 25]).
+A prominent line of research focuses on the online setting [16, 3], where jobs arrive over time and the scheduler must make irrevocable decisions without knowledge of future arrivals.
+In modern applications such as cloud resource management and data center scheduling [21, 14, 12], an online service platform must dispatch jobs submitted by multiple users as they arrive; this naturally leads to the objective of minimizing the total flow time (or equivalently, the average flow time), defined as the sum over all jobs of the difference between completion time and release time.
 In this paper, we focus on the flow time minimization problem on $m$ identical machines.
-Most studies on the flow time objective allow additional capabilities for algorithms, such as preemption, resource augmentation, and rejection.
-Many papers (e.g., [DBLP:conf/focs/GargK07, DBLP:conf/esa/LucarelliTST16, DBLP:conf/stoc/ChadhaGKM09, DBLP:conf/soda/ChenIP25, DBLP:conf/icalp/GuptaKL18, DBLP:journals/dam/EpsteinS06]) have argued that they are necessary to surpass the trivial greedy algorithm [^1] The trivial greedy algorithm simply schedules any waiting job whenever any machine becomes available, achieving a competitive ratio of $\Theta(n/m)$.
-because of the following single-machine lower bound consisting of $n$ jobs: the adversary releases a unit job at time 0 and, upon scheduling, reveals n1n-1 $\varepsilon$-jobs, forcing any deterministic non-preemptive algorithm to incur $\Omega(n)$ flow time while the offline optimum remains $O(1)$ [DBLP:journals/siamcomp/KellererTW99, DBLP:conf/stoc/ChekuriKZ01].
+Most positive results (e.g., [27, 4, 6, 17, 10]) on the flow time objective require additional capabilities — such as preemption, resource augmentation, and rejection — to surpass the trivial greedy algorithm [^1] The trivial greedy algorithm simply schedules any waiting job whenever any machine becomes available, achieving a competitive ratio of $\Theta(n/m)$..
+This is due to the following single-machine lower bound, which holds even when $n$ is known in advance:
+The adversary first releases a unit job at time 0. As soon as the algorithm starts processing it or after waiting for $n$ units of time, the adversary releases n1n-1 $\varepsilon$-jobs.
+Any deterministic non-preemptive algorithm is thus forced to incur $\Omega(n)$ flow time, whereas the offline optimum remains $O(1)$ [23, 5].
 However, this lower bound applies only to deterministic algorithms in the single-machine setting.
@@
 Randomization is widely employed in the design of online algorithms.
-Against an oblivious adversary who does not adapt to the randomness used by the algorithm, randomization often breaks deterministic barriers, including the classic caching problem and matching problem [DBLP:journals/jal/FiatKLMSY91, DBLP:conf/stoc/KarpVV90].
-In our problem, the previously known lower bound $\Omega(\sqrt{n})$ for oblivious adversaries by DBLP:journals/tcs/EpsteinS03 applies only to single-machine; no information-theoretic lower bound stronger than $\Omega(1)$ was previously known when $m\geq 2$.
-A gap remains between the current best-known $\Theta(n/m)$-competitive trivial greedy algorithm and the largely unanswered questions about lower bounds.
-Question for Randomized Algorithms Can we use randomization to beat the trivial greedy algorithm?
-Is it possible to find an algorithm much better than $\Theta(\sqrt{n})$ (perhaps even a constant) when $m\geq 2$?
+Against an oblivious adversary who does not adapt to the randomness used by the algorithm, randomization often breaks deterministic barriers, including the classic caching problem and matching problem [11, 22].
+In our problem, the previously known lower bound $\Omega(\sqrt{n})$ for oblivious adversaries by [9] applies only to the single-machine case; for $m\geq 2$, no information-theoretic lower bound stronger than $\Omega(1)$ was previously known, leaving a substantial gap from the $O(n/m)$ competitive ratio of the trivial greedy algorithm.
+Open Question 1: Randomized Algorithms Can we use randomization to beat the trivial greedy algorithm?
+Is it possible to find a randomized algorithm much better than $\Theta(\sqrt{n})$ (perhaps even a constant) when $m\geq 2$?
 Deterministic Multi-Machine Case.
-In fact, even without randomization, the problem remains unsolved when $m$ is large: the deterministic lower bound $\Omega(n/m^{2})$ established by DBLP:journals/dam/EpsteinS06 decreases much faster than the trivial upper bound $\Theta(n/m)$ achieved by the greedy algorithm, leaving a substantial gap when $m$ is comparable to $n$.
-Question for Deterministic Algorithms Is greedy indeed the best deterministic algorithm for multi-machine case? What is the optimal competitive ratio when $m$ is large?
+In fact, even without randomization, the problem remains unsolved when $m$ is large: the deterministic lower bound $\Omega(n/m^{2})$ established by [10] decreases much faster than the trivial upper bound $\Theta(n/m)$ achieved by the greedy algorithm, leaving a substantial gap when $m$ is comparable to $n$.
+Open Question 2: Deterministic Algorithms Is greedy indeed the best deterministic algorithm for the multi-machine case? What is the optimal competitive ratio when $m$ is large?
 Kill-and-Restart Model.
-Another direction that is closely related to the non-preemptive setting is the kill-and-restart capability, first introduced by DBLP:journals/siamcomp/ShmoysWW95, where the scheduler is allowed to kill a job and restart it from scratch, losing all processed work.
-Kill-and-restart lies between preemption and non-preemption.
-This model is particularly suitable in settings where partially processed work cannot be resumed due to system or application constraints (e.g., data corruption or non-checkpointable tasks).
-It also has a practical advantage over preemption, as it avoids the need to maintain and store potentially unbounded intermediate states of suspended jobs.
-This notion has been studied under objectives such as makespan and total completion time, where it has been shown to surpass previously known lower bounds [DBLP:conf/approx/0002KTWZ18, DBLP:journals/mp/JagerSWW25, DBLP:journals/jal/SteeP05].
-Notably, DBLP:journals/mp/JagerSWW25 considered online non-clairvoyant scheduling with the objective of minimizing total completion time and demonstrated that kill-and-restart significantly improves the competitive ratio from $\Omega(n)$ to $O(1)$.
-Unlike the completion time objective where kill-and-restart has proven powerful, its potential for flow time minimization has remained largely unexplored.
-No algorithms were previously known to utilize the kill-and-restart capability in the flow time minimization problem, and the only known lower bound $\Omega(\sqrt{n})$ by DBLP:journals/tcs/EpsteinS03 for the kill-and-restart model is also restricted to the single-machine case.
-Since the only known algorithm without preemption is the $O(n/m)$-competitive greedy algorithm, it remained unclear whether the kill-and-restart capability can improve competitiveness in flow-time minimization, and to what extent such improvement is possible.
-Question for Kill-and-Restart Can we use kill-and-restart to beat the trivial greedy algorithm?
-Is it possible to find an algorithm much better than $\Theta(\sqrt{n})$ (perhaps even a constant) when $m\geq 2$?
-1.1 Our Results
-We give a complete set of answers to the above questions, summarized in Table 1 (the single-machine case) and Table 2 (the multi-machine case).
+A closely related direction is the kill-and-restart model, first introduced by [29], where the scheduler may kill a running job and restart it from scratch, losing all processed work.
+This model is particularly well-suited to applications where partially processed work cannot be resumed (e.g., data corruption or non-checkpointable tasks).
+Unlike preemption, it avoids the need to maintain and store potentially unbounded intermediate states of suspended jobs, preserving the non-preemptive nature of job execution while giving the scheduler additional flexibility to correct past decisions.
+Kill-and-restart has proven powerful for other scheduling objectives.
+For makespan and total completion time, it surpasses known non-preemptive barriers [19, 30]; most notably, [20] showed that for non-clairvoyant scheduling under total completion time, kill-and-restart improves the competitive ratio from $\Omega(n)$ to $O(1)$.
+Despite these successes, the potential of kill-and-restart for flow time minimization has remained largely unexplored.
+No algorithm was previously known to exploit this capability, and the only known lower bound $\Omega(\sqrt{n})$ [9] is restricted to the single-machine case.
+It thus remained unclear whether, and to what extent, kill-and-restart can improve upon the $O(n/m)$-competitive greedy algorithm.
+Open Question 3: Kill-and-Restart Can we use kill-and-restart capability to beat the trivial greedy algorithm?
+Is it possible to find an algorithm under the kill-and-restart model that is much better than $\Theta(\sqrt{n})$ (perhaps even a constant) when $m\geq 2$?
+1.1 Our Results with Known $n$
+We give a complete set of answers to Open Questions 1–3 when $n$ is known in advance, summarized in Table 1 (the single-machine case) and Table 2 (the multi-machine case).
 A key technical feature of our lower bounds in the multi-machine case is their robustness: unlike bounds that only apply to specific choices of these parameters, they hold for every fixed pair of $(n,m)$.
-The results presented in this section assume prior knowledge of $n$.
-This assumption is common when we study the non-preemptive version of the flow time minimization problem [DBLP:journals/dam/EpsteinS06, DBLP:journals/tcs/EpsteinS03], and we will further discuss this in Section 1.2 and present results on the algorithms without prior knowledge of $n$.
+Throughout this paper, we assume $n\geq m$ when presenting competitive ratio; when $n<m$, all jobs can be processed immediately without waiting, yielding optimal schedules with a competitive ratio of $1$.
 Randomized Algorithms: Beating the Deterministic Barrier
-We first address whether randomization can overcome the $\Omega(n)$ deterministic barrier for non-preemptive algorithms.
-We answer this affirmatively by presenting an $O(\sqrt{n})$-competitive randomized algorithm for the single-machine case, matching the established lower bound by DBLP:journals/tcs/EpsteinS03.
-We then extend this to the multi-machine setting, achieving an $O(\sqrt{n/m})$ competitive ratio.
-This result is optimal, as we prove a matching $\Omega(\sqrt{n/m})$ lower bound that holds for every pair of $(n,m)$.
-This completely characterizes the power of randomization in the non-preemptive setting.
+We show that randomization can overcome the $\Omega(n)$ deterministic barrier.
+For the single-machine case, we present an $O(\sqrt{n})$-competitive randomized algorithm, matching the lower bound of [9].
+For the multi-machine case, we achieve an $O(\sqrt{n/m})$ competitive ratio and prove a matching $\Omega(\sqrt{n/m})$ lower bound that holds for every pair $(n,m)$.
+Together, these results completely characterize the power of randomization in non-preemptive flow time minimization.
 Deterministic Algorithms: Closing the Multi-Machine Gap
@@
 The Power of Kill-and-Restart
-Perhaps our most surprising finding is non-uniform impact of the kill-and-restart capability on deterministic algorithms.
-  • For the single-machine case, we show that kill-and-restart offers little advantage, proving a strong lower bound of $\Omega(n/\log n)$.
-  • However, with just one additional machine, the problem’s competitive ratio collapses.
-We present a deterministic algorithm with kill-and-restart that achieves an $O(\sqrt{n/m})$ competitive ratio for all $m\geq 2$.
-This is a quadratic improvement over the single-machine case and remarkably matches the optimal bound achievable with randomization.
-Conversely, we show that for randomized algorithms, kill-and-restart provides no additional power, as the $\Omega(\sqrt{n/m})$ lower bound holds even in this stronger model.
+Perhaps our most surprising finding is the non-uniform impact of the kill-and-restart capability on deterministic algorithms.
+  • For the single-machine case, kill-and-restart offers little advantage: we prove a lower bound of $\Omega(n/\log n)$.
+  • For $m\geq 2$, the competitive ratio drops dramatically. We present a deterministic kill-and-restart algorithm achieving $O(\sqrt{n/m})$, a quadratic improvement over the single-machine case.
+Remarkably, this deterministic kill-and-restart ratio of $O(\sqrt{n/m})$ matches the optimal randomized bound.
+We further show that this coincidence is tight: the $\Omega(\sqrt{n/m})$ lower bound holds even for randomized algorithms with kill-and-restart.
 Improving the State-of-the-Art for Offline Scheduling
 A direct consequence of our online results is a new state-of-the-art for the classic offline problem.
-We improve the long-standing best approximation ratio from $O(\sqrt{n/m}\log(n/m))$ by DBLP:journals/jcss/LeonardiR07 to $O(\sqrt{n/m})$.
+We improve the long-standing best approximation ratio from $O(\sqrt{n/m}\log(n/m))$ by [25] to $O(\sqrt{n/m})$.
 The improved bound can be realized by either derandomizing our randomized algorithm or applying our deterministic kill-and-restart algorithm in an offline context where restarts can be ignored at no cost.
 Table 1: Summary of competitive ratios for the single-machine setting.
+Shaded cells indicate results from this paper.
 Single machine, $n$ jobs Non-Preemptive Kill-and-Restart
@@
 Lower Bound $\Omega(n)$ $\Omega(n/\log n)$
-[DBLP:journals/siamcomp/KellererTW99, DBLP:conf/stoc/ChekuriKZ01] Theorem 7
+[23, 5] Theorem 7
 Randomized Upper Bound $O(\sqrt{n})$ $O(\sqrt{n})$
@@
 Lower Bound $\Omega(\sqrt{n})$ $\Omega(\sqrt{n})$
-[DBLP:journals/tcs/EpsteinS03] [DBLP:journals/tcs/EpsteinS03]
+[9] [9]
 Table 2: Summary of competitive ratios for the multi-machine setting.
+Shaded cells indicate results from this paper.
 $m$ machines, $n$ jobs Non-Preemptive Kill-and-Restart
@@
 Theorem 5 Theorem 6
-1.2 Further Discussion of the Prior Knowledge of $n$
-In our main results, we follow the common convention in non-preemptive flow-time scheduling by assuming that $n$, the total number of jobs, is known in advance.
-This assumption is common to nearly all significant prior work in this area, including both upper and lower bounds [DBLP:journals/dam/EpsteinS06, DBLP:journals/tcs/EpsteinS03].
-The necessity of this assumption is underscored by strong lower bounds showing that without it, no deterministic algorithm can beat the trivial $O(n)$ competitive ratio, even if the algorithm can use $\ell$ machines but the optimal solution can only use one [DBLP:journals/dam/EpsteinS06].
-However, this barrier does not preclude the possibility that randomization or the kill-and-restart capability could yield improvements in the more challenging “unknown nn” setting.
-We investigate this question and show that these two tools have different capabilities:
-  • Randomization alone is insufficient.
-We prove that for a single machine, no randomized algorithm can achieve a competitive ratio of $o(n)$ without knowing $n$.
-For the multi-machine setting, we further show that no randomized algorithm can achieve a competitive ratio of $o(n/m^{2}+\sqrt{n/m})$ without prior knowledge of $n$.
-This demonstrates that knowing $n$ is necessary for designing non-trivial randomized algorithms in the non-preemptive setting.
+1.2 Our Results without Prior Knowledge of $n$
+All results above assume that the total number of jobs $n$ is known in advance, a convention shared by prior positive results on non-preemptive flow time minimization [10, 9].
+This assumption is critical for deterministic algorithms: [10] showed that without knowledge of $n$, no deterministic algorithm on any constant number of machines can achieve a competitive ratio of $o(n)$, even when compared against a single-machine offline optimum.
+Is this assumption also critical for randomized algorithms and algorithms with kill-and-restart capability?
+Open Question 4: Unknown Number of Jobs Can randomization or kill-and-restart still overcome the deterministic $\Theta(n)$ barrier when $n$ is not known in advance?
+We find a sharp separation between the two.
+Randomization is insufficient.
+We prove that without knowledge of $n$, no randomized non-preemptive algorithm can achieve a competitive ratio of $o(n)$ on a single machine, or $o(n/m^{2}+\sqrt{n/m})$ on $m$ machines.
 Kill-and-restart overcomes the barrier.
-In sharp contrast, we show that the kill-and-restart model is powerful enough to break the $O(n)$ barrier when $m\geq 2$.
-We extend our deterministic kill-and-restart algorithm to operate without knowledge of $n$ and prove that it achieves an $O(n^{\alpha}/\sqrt{m})$-competitive ratio, where $\alpha=(\sqrt{5}-1)/2\approx 0.618$ is the golden ratio.
-This demonstrates that the kill-and-restart capability provides a fundamental advantage over purely randomized approaches in this setting.
-Finally, we note that our algorithms are robust to estimation.
-If only an approximate estimate $\hat{n}\in[n/c,cn]$ is known instead of the accurate value of $n$, as might be provided by heuristics or machine-learned advice, all algorithms’ guarantees hold within a constant factor of $c$, preserving their asymptotic performance.
+In contrast, we extend our deterministic kill-and-restart algorithm to operate without knowledge of $n$, achieving a competitive ratio of $O(n^{\alpha}/\sqrt{m})$ for all $m\geq 2$, where $\alpha=(\sqrt{5}-1)/2\approx 0.618$.
+This significantly improves over the trivial $O(n/m)$ bound and demonstrates that kill-and-restart provides a fundamental advantage that persists even without knowledge of $n$.
+Algorithms with predicted $n$.
+Finally, we note that if an approximate prediction $\hat{n}\in[n/c,cn]$ is available instead of the exact value of $n$, all our known-$n$ algorithms’ guarantees degrade by at most a constant factor of $c$, preserving their asymptotic performance.
 1.3 Our Techniques
@@
 This distinction is key to circumventing the classic $\Omega(n)$ lower bound, where handling a single large job at the wrong time can delay a swarm of subsequent small jobs, leading to high flow time.
-Offline algorithms can address this by partitioning jobs based on global properties of the total job set (e.g., the offline approximation algorithms by DBLP:journals/siamcomp/KellererTW99 and DBLP:journals/jcss/LeonardiR07 relied on a pre-solved preemptive solution over all jobs, and use the flow time of jobs in the preemptive solution to partition jobs), but this is impossible to implement in an online context, since future jobs are unknown.
+Offline algorithms can address this by partitioning jobs based on global properties of the total job set (e.g., the offline approximation algorithms by [23] and [25] relied on a pre-solved preemptive solution over all jobs, and used the flow time of jobs in the preemptive solution to partition jobs), but this is impossible to implement in an online context, since future jobs are unknown.
 Therefore, a fundamentally new approach is required for the online setting where decisions are irrevocable and based on incomplete information.
@@
 However, it also introduces two significant technical challenges that we resolve:
+  • (1)
     Managing dynamic re-classification.
@@
 To handle this transition gracefully, we introduce the concept of proxy jobs.
-Instead of re-label a large job when it should become small, we treat it as a newly arrived small job, called a proxy job.
+Instead of relabeling a large job when it should become small, we treat it as a newly arrived small job, called a proxy job.
 We then decompose its total flow time into two parts: the waiting time during its large phase, and the flow time of the proxy job.
 Note that a proxy job is not an actual job from the original instance, so we must further bound the additional impact introduced by creating such proxy jobs.
+  • (2)
     Scheduling small jobs without a global view.
@@
 A key advantage of this framework is that it no longer requires a preemptive algorithm to exploit global properties of the entire job set.
-This distinction not only enables our algorithm to be implemented online, but also improves the offline approximation ratio by a logarithmic factor over the result of DBLP:journals/jcss/LeonardiR07, since their logarithmic loss arises from relying on a suboptimal preemptive algorithm in the multi-machine setting.
-We believe that our online partitioning framework and the analysis of NSJF provide a general technical insight that can be applied to broader classes of non-preemptive scheduling problems, such as the flow time minimization problem in the related machine setting.
+This distinction not only enables our algorithm to be implemented online, but also improves the offline approximation ratio by a logarithmic factor over the result of [25], since their logarithmic loss arises from relying on a suboptimal preemptive algorithm in the multi-machine setting.
+We remark that our online partitioning framework and the analysis of NSJF do not inherently rely on the machines being identical, and we expect them to be useful for broader settings such as related machines.
 1.4 Paper Organization
 The remainder of this paper is organized as follows.
-In Section 3, we introduce our core technical contribution: the online rank-based partitioning framework that serves as the foundation for all our algorithms.
-The analysis of the NSJF algorithm, used for handling small jobs within this framework, is detailed in Appendix A.
+In Section 3, we introduce our core technical contribution: the online rank-based partitioning framework that serves as the foundation for all our algorithms, and the analysis of the NSJF algorithm for handling small jobs within this framework.
+In the deterministic algorithms, we use a generalized analysis of the NSJF algorithm, which is detailed in Appendix A.
 We then present our main algorithmic results.
 Section 4 develops our $O(\sqrt{n/m})$-competitive randomized non-preemptive algorithm.
-Section 5 presents our deterministic non-preemptive algorithm.
-Section 6 details our deterministic $O(\sqrt{n/m})$-competitive algorithm using kill-and-restart for $m\geq 2$.
-We then address the unknown-$n$ setting in Section 7 with an extension of this algorithm.
+Section 5 presents our deterministic non-preemptive $O(n/m^{2}+\sqrt{n/m}\log m)$-competitive algorithm.
+Section 6 details our deterministic $O(\sqrt{n/m})$-competitive algorithm using kill-and-restart for $m\geq 2$, and then we extend it to the unknown-$n$ setting in Section 7.
 Finally, we establish the tightness of our results in Section 8 and Section 9, where we present a comprehensive set of matching lower bounds for both the known-$n$ and unknown-$n$ settings, respectively.
@@
 In the classic non-preemptive offline setting, the problem is known to be NP-hard.
-A key result by DBLP:journals/jcss/LeonardiR07 provided an $O(\sqrt{n/m}\log(n/m))$-approximation algorithm, which stood as the best-known guarantee for many years.
-On the approximation hardness front, DBLP:journals/siamcomp/KellererTW99 established a polynomial-time lower bound of $\Omega(\sqrt{n})$ for the single-machine case, later complemented by DBLP:journals/jcss/LeonardiR07, presenting an $\Omega(n^{1/3-\varepsilon})$ lower bound for multiple machines when $n\geq m^{4/\varepsilon}$.
+A key result by [25] provided an $O(\sqrt{n/m}\log(n/m))$-approximation algorithm, which stood as the best-known guarantee for many years.
+On the approximation hardness front, [23] established a polynomial-time lower bound of $\Omega(\sqrt{n})$ for the single-machine case, later complemented by [25], who presented an $\Omega(n^{1/3-\varepsilon})$ lower bound for multiple machines when $n\geq m^{4/\varepsilon}$.
 In the online setting, the problem is notoriously difficult, with a simple greedy algorithm achieving a $\Theta(n/m)$ competitive ratio, which for a long time was the best-known result for deterministic algorithms.
@@
 In this setting, the folklore SRPT (Shortest Remaining Processing Time) policy is optimal on a single machine.
-For multiple machines, SRPT achieves a tight competitive ratio of $\Theta(\log(n/m))$ when job migration is allowed [DBLP:journals/jcss/LeonardiR07].
-For the more restrictive model without migration, DBLP:journals/siamcomp/AwerbuchALR02 presented an algorithm with a competitive ratio of $O(\log(\min\{\log n,\log P\}))$.
+For multiple machines, SRPT achieves a tight competitive ratio of $\Theta(\log(n/m))$ when job migration is allowed [25].
+For the more restrictive model without migration, [1] presented an algorithm with a competitive ratio of $O(\log(\min\{\log n,\log P\}))$.
 Models with Algorithmic Relaxations.
@@
 One common approach is resource augmentation, where the algorithm is provided with faster or more machines than the adversary.
-DBLP:journals/algorithmica/PhillipsSTW02 showed that an $O(1)$-competitive ratio is achievable with $m\log P$ machines, where $P$ is the max-to-min job size ratio.
+[28] showed that an $O(1)$-competitive ratio is achievable with $m\log P$ machines, where $P$ is the max-to-min job size ratio.
 Another popular relaxation is the rejection model, where the algorithm can choose to reject a small fraction of jobs.
-DBLP:journals/jcss/ChoudhuryDGK18 first introduced this notion for flow time, with subsequent works [DBLP:conf/icalp/GuptaKL18, DBLP:conf/esa/LucarelliMTST18] achieving an $O(1/\varepsilon^{3})$-competitive algorithm for rejecting an $\varepsilon$-fraction of jobs.
+[7] first introduced this notion for flow time, with subsequent works [17, 26] achieving an $O(1/\varepsilon^{3})$-competitive algorithm for rejecting an $\varepsilon$-fraction of jobs.
 The Kill-and-Restart Model.
-The kill-and-restart capability, introduced by DBLP:journals/siamcomp/ShmoysWW95 in the context of makespan minimization, represents a minimal form of preemption.
+The kill-and-restart capability, introduced by [29] in the context of makespan minimization, represents a minimal form of preemption.
 This model has been successfully applied to other objectives, demonstrating its power to surpass non-preemptive lower bounds.
-For minimizing total completion time, DBLP:journals/jal/SteeP05 designed a $3/2$-competitive algorithm, breaking the non-preemptive barrier of $1.582$.
-Even more dramatically, for non-clairvoyant scheduling, DBLP:journals/mp/JagerSWW25 showed that kill-and-restart reduces the competitive ratio for total completion time from $\Omega(n)$ to $O(1)$.
+For minimizing total completion time, [30] designed a $3/2$-competitive algorithm, breaking the non-preemptive barrier of $1.582$.
+Even more dramatically, for non-clairvoyant scheduling, [20] showed that kill-and-restart reduces the competitive ratio for total completion time from $\Omega(n)$ to $O(1)$.
 Despite these successes, the potential of kill-and-restart for flow time minimization has remained largely unexplored.
-The primary work in this specific area was by DBLP:journals/tcs/EpsteinS03, who established several lower bounds, including an $\Omega(\sqrt{n})$ bound for randomized algorithms on a single machine, but no upper bounds were known.
+The primary work in this specific area was by [9], who established several lower bounds, including an $\Omega(\sqrt{n})$ bound for randomized algorithms on a single machine, but no upper bounds were known.
 2 Preliminaries
-In this section, we formally define the scheduling problem, the algorithmic models under consideration, and the metrics used for performance evaluation.
 2.1 Problem Formulation
@@
 A job becomes known to the algorithm at its release time $r_{j}$ and must be assigned to a machine at or after $r_{j}$.
-The job must be processed on a single machine at a time, and must be processed for $p_{j}$ time in total to complete.
-One machine can serve at most one job at a time.
-We introduce standard notation to describe the status of a job or a machine at a given time $t$.
+Each job must be processed on a single machine at a time and requires $p_{j}$ units of processing to complete.
+Each machine can process at most one job at a time.
+We introduce standard terminology for describing the state of jobs and machines.
 A machine is said to be busy at time $t$ if it is processing a job, and idle otherwise.
 Time $t$ is said to be busy if all machines are busy at time $t$.
-A job $j$ is said to be waiting at time $t$ if it has been released but not yet started, i.e., $r_{j}\leq t<s_{j}$.
-A job is said to be processing at time $t$ if it is being processed by a machine.
+A job $j$ is said to be waiting at time $t$ if it has been released but is not currently being processed by any machine.
 2.2 Scheduling Models
-A schedule $\S$ assigns a start time $s_{j}\geq r_{j}$ to each job $j$ on one of the $m$ machines. This paper focuses on two primary models:
+This paper focuses on two primary models:
   • Non-Preemptive: Once a job begins execution on a machine, it must run to completion without interruption.
-  • Kill-and-Restart: Algorithms are permitted to terminate (kill) a job that is currently running.
-The job then returns to the set of waiting jobs and can be scheduled again (restart) from the beginning at a later time, possibly on a different machine.
+  • Kill-and-Restart: The algorithm may terminate (kill) a job that is currently running.
+The job returns to the pool of available jobs and can be started again from the beginning at a later time, possibly on a different machine.
 All work done prior to the termination is lost.
 For the purpose of analysis, we also refer to the following preemptive models.
-  • Preemptive (without Migration): The algorithm is allowed to interrupt a job and later resume it on the same machine without losing any processed work.
-  • Preemptive with Migration: The algorithm is allowed to interrupt a job and later resume it on any machine without losing any processed work.
-2.3 Objective and Performance Analysis
-Let $\S=\textup{{ALG}}(J)$ be the schedule produced by some algorithm ALG on job set $J$.
-We write $s_{j}(\S)$ for the start time of job $j$ in schedule $\S$, and define its completion time as $C_{j}(\S)=s_{j}(\S)+p_{j}$.
-The flow time of job $j$ is $F_{j}(\S)=C_{j}(\S)-r_{j}$, and the objective is to minimize the total flow time on the job set $J$:
-$F(\S)=\sum_{j\in J}F_{j}(\S).$
-When the schedule $\S$ is clear from context, we omit it from the notation and write $s_{j}$, $C_{j}$, and $F_{j}$ instead.
+  • Preemptive (without Migration): The algorithm may interrupt a job and later resume it on the same machine without losing any processed work.
+  • Preemptive with Migration: The algorithm may interrupt a job and later resume it on any machine without losing any processed work.
+2.3 Objective and Performance Metric
+We define $C_{j}$ as the time when job $j$ finishes its last unit of processing and its flow time as $F_{j}=C_{j}-r_{j}$.
+For the non-preemptive and kill-and-restart models, we write $s_{j}$ for the start time of job $j$’s (final successful) execution, and we have $C_{j}=s_{j}+p_{j}$.
+The objective is to minimize the total flow time:
+$F(\S)=\sum_{j\in J}F_{j}.$
 Competitive Analysis.
 The performance of an online algorithm is measured by its competitive ratio.
-An algorithm ALG is $\Gamma$-competitive [^4] The competitive ratio, $\Gamma$, is often a function $\Gamma(n,m)$ of the number of jobs $n$ and machines $m$ in this paper.
+An algorithm ALG is $\Gamma$-competitive [^4] The competitive ratio $\Gamma$ is often a function $\Gamma(n,m)$ of the number of jobs $n$ and machines $m$ in this paper.
 if for any job instance $J$:
-$F(\textup{{ALG}}(J))\leq\Gamma\cdot F(\textup{{OPT}}(J))$
-where $\textup{{OPT}}(J)$ is the total flow time of an optimal offline solution. For the purpose of analysis, we define OPT differently for our upper and lower bounds:
-  • For our upper bounds, we compare our online algorithm against a more powerful adversary: OPT is the optimal offline preemptive schedule without migration.
-  • For our lower bounds, we use the offline non-preemptive schedule as OPT.
-For convenience, when the context is clear, we will use OPT and ALG to denote the costs $F(\textup{{OPT}}(J))$ and $F(\textup{{ALG}}(J))$, respectively.
+$F(\textup{{ALG}}(J))\leq\Gamma\cdot F(\textup{{OPT}}(J)),$
+where $\textup{{OPT}}(J)$ is the total flow time of an optimal offline solution. We define OPT differently for upper and lower bounds:
+  • For upper bounds, OPT is the optimal offline preemptive schedule without migration.
+  • For lower bounds, OPT is the optimal offline non-preemptive schedule.
+For convenience, we use OPT and ALG to denote the costs $F(\textup{{OPT}}(J))$ and $F(\textup{{ALG}}(J))$, respectively, when the context is clear.
 3 Online Rank-Based Partitioning Framework
@@
   • The processing time of any job classified as small is bounded by $\textup{{OPT}}/\ell$.
-To achieve this, we maintain a set of active large jobs, denoted $A$, containing the $\ell$ jobs with the largest processing times seen so far.
+To achieve this, we maintain a set of active large jobs, denoted $A$, containing up to $\ell$ jobs with the largest processing times seen so far.
 The logic of our framework defines a clear life cycle for every job based on its interaction with this set.
@@
 Data: Active large job set $A$, and every job’s other current states
-Function Partition(ii, $\ell$):
+Procedure Partition(ii, $\ell$):
  if $|A|<\ell$ then
-      Add $i$ to $A$;
-    return $i$ is large
+      Add $i$ to AA;
+      Mark $i$ as large;
  else if $p_{i}>\min_{j\in A}p_{j}$ then
       Let $j=\arg\min_{j\in A}p_{j}$;
-      Remove $j$ from $A$ and add $i$ to $A$;
+    // $j$ will be retired.
+      Remove $j$ from $A$ and add $i$ to AA;
     if job $j$ is still waiting then
          Create a new proxy job $j^{\prime}$ with $r_{j^{\prime}}=r_{i}$ and $p_{j^{\prime}}=p_{j}$;
+         Mark $j^{\prime}$ as small;
          Mark $j$ as proxied by $j^{\prime}$;
-       // The scheduler will actually process $j$ when it processes $j^{\prime}$.
+       // The scheduler will actually process $j$ when processing $j^{\prime}$.
     else
-       // $j$ is a committed large job.
+         Mark $j$ as committed large;
       end if
-    return $i$ is large
+     Mark $i$ as large;
  else
-    return $i$ is small
+      Mark $i$ as small;
    end if
@@
 For clarity, we will first prove the lemma as stated.
-In Appendix A, we prove a more general version of this result (Lemma 29) that is robust to interference from large jobs, where machines may be subject to initial “blocking periods.”, during which they are unavailable for scheduling small jobs.
-Specifically, given a vector of initial blocking times $\vec{b}$ that NSJF cannot use a machine $i$ in its blocking period $0,b_{i})$, we show that NSJF remains competitive with respect to the optimal flow time on an instance without blocking, up to an extra additive factor depending on the total blocking length.
+In Appendix A, we prove a more general version of this result (Lemma 29) that is robust to interference from large jobs, where machines may be subject to initial “blocking periods,” during which they are unavailable for scheduling small jobs.
+Specifically, given a vector of initial blocking times $\vec{b}$ such that NSJF cannot use a machine $i$ in its blocking period $0,b_{i})$, we show that NSJF remains competitive with respect to the optimal flow time on an instance without blocking, up to an extra additive term depending on the total blocking length.
 This general version is a fundamental component in the final analysis of our main algorithms.
@@
 Note that the general version of Lemma 29 follows the same strategy; however, the analysis must additionally account for how the available processing power varies over time in the presence of blocking periods.
-Proof of Lemma 2.
-To begin, we introduce notation to compare the progress of NSJF against OPT.
 For any processing time threshold $p$ and time $t$, let $J_{\leq p}(t)$ be the set of jobs with $p_{j}\leq p$ and release time $r_{j}\leq t$.
-We measure progress on this set differently for each algorithm:
-  • For OPT, let $V_{\leq p}(t)$ denote the total volume of jobs in $J_{\leq p}(t)$ that has been processed by time $t$.
-  • For NSJF, let $V^{\prime}_{\leq p}(t)$ denote the total volume of jobs in $J_{\leq p}(t)$ that have started by time $t$.
-Our first step is to prove the following volume-matching lemma.
-Lemma 3.
-For every $t\geq 0$, and every upper bound $p$, we have: $V^{\prime}_{\leq p}(t+\tau)\geq V_{\leq p}(t).$
+Step 1: Volume Matching.
+For each algorithm, we track the volume of work associated with jobs in $J_{\leq p}(t)$.
+Let $V_{\leq p}(t)$ denote the total volume of work processed on jobs in $J_{\leq p}(t)$ by time $t$ under OPT, and let $V^{\prime}_{\leq p}(t)$ denote the total volume of jobs in $J_{\leq p}(t)$ that have been started by time $t$ under NSJF.
+Lemma 3 (Volume Matching).
+For every $t\geq 0$ and every threshold $p$, we have
+$V^{\prime}_{\leq p}(t+\tau)\geq V_{\leq p}(t).$
 Proof.
-Fix an arbitrary $p$, let $t_{0}$ be the last time some machine is idle by NSJF before $t+\tau$.
-Define $s_{j}$ as the latest start time, within the interval $(t_{0},\,t+\tau)$, at which NSJF schedules a job with $p_{j}>p$.
-If no such job exists, we set $s_{j}=t_{0}$ and conceptually assume that an infinitely large job starts at time $s_{j}$.
+Fix an arbitrary $p$, and let $t_{0}$ be the last time some machine is idle under NSJF before $t+\tau$.
+Let $j$ be the job with $p_{j}>p$ that is started latest by NSJF within the interval $(t_{0},t+\tau)$.
+If no such job exists, we let $s_{j}=t_{0}$ for a fictitious job $j$ of infinite size.
 Since NSJF starts job $j$ with $p_{j}>p$ at time $s_{j}$, it must have already started all jobs with size at most $p$ released up to $s_{j}$ (due to its SJF policy). Thus, we have:
 $V^{\prime}_{\leq p}(s_{j})=\sum_{h\in J_{\leq p}(s_{j})}p_{h}\geq V_{\leq p}(s_{j}).$
-We now analyze two cases to conclude the lemma. If $s_{j}\geq t$. We straightforwardly have:
+We now analyze two cases. If $s_{j}\geq t$, we immediately have:
 $V^{\prime}_{\leq p}(t+\tau)\geq V^{\prime}_{\leq p}(s_{j})\geq V_{\leq p}(s_{j})\geq V_{\leq p}(t).$
-Otherwise, if $s_{j}<t$, the time interval $s_{j}+\tau,t+\tau)$ must be fully used by NSJF to process jobs with size at most $p$ (if any). We have:
+Otherwise, if $s_{j}<t$, then since $s_{j}$ is the latest start of a job with $p_{j}>p$ after the last idle time $t_{0}$, all processing in $s_{j}+\tau,t+\tau)$ is on jobs of size at most $p$, and no machine is idle in this interval.
+Therefore:
 $V^{\prime}_{\leq p}(t+\tau)\geq V_{\leq p}(s_{j})+m(t-s_{j})\geq V_{\leq p}(t).$
-Denote $N_{\leq p}(t)$ (or $N^{\prime}_{\leq p}(t)$ below) as the number of jobs in $J_{\leq p}(t)$ completed by time $t$ by OPT (or NSJF, respectively).
-Note that without the subscription $\leq p$, we remove the condition on job size.
-We prove:
-Lemma 4.
-For every $t\geq 0$, and every upper bound $p$, we have $N^{\prime}_{\leq p}(t+2\tau)\geq N_{\leq p}(t)$.
+Step 2: Completion Matching.
+Let $N_{\leq p}(t)$ and $N^{\prime}_{\leq p}(t)$ denote the number of jobs in $J_{\leq p}(t)$ completed by time $t$ under OPT and NSJF, respectively.
+When the subscript $\leq p$} is omitted, we count all completed jobs regardless of size.
+Lemma 4 (Completion Matching).
+For every $t\geq 0$ and every threshold $p$, we have
+$N^{\prime}_{\leq p}(t+2\tau)\geq N_{\leq p}(t).$
 Proof.
-We discuss two timepoints $t+\tau$ and $t+2\tau$. By Lemma 3, we have the volume bound: $V^{\prime}_{\leq p}(t+\tau)\geq V_{\leq p}(t)$ for every $p$.
-We claim the following: the $i$-th smallest job (in terms of size) that has started (either in processing or completed) under NSJF by time $t+\tau$, denoted by size $p^{\prime}_{i}$, must not exceed the size of the $i$-th smallest completed job under OPT by time $t$, denoted $p_{i}$.
-That is, we aim to prove: $p^{\prime}_{i}\leq p_{i}$ for all rank $i$.
-To prove this, assume for contradiction that there exists the smallest rank $i$ such that $p^{\prime}_{i}>p_{i}$. Under this assumption, we have the following bound for OPT:
-$V_{\leq p_{i}}(t)\geq\sum_{j=1}^{i}p_{j},$
-since at least $i$ jobs of size at most $p_{i}$ are completed by time $t$ in OPT.
-Now consider NSJF.
-Since the $i$-th smallest started job under NSJF has a size strictly greater than $p_{i}$, all started jobs of size $\leq p_{i}$ must be among the first i1i-1 jobs.
-Therefore, we have:
+For each rank $i$, let $p_{i}$ be the size of the $i$-th smallest job completed under OPT by time $t$, and let $p^{\prime}_{i}$ be the size of the $i$-th smallest job started under NSJF by time $t+\tau$.
+Let $k$ and $k^{\prime}$ denote the total number of such jobs, respectively.
+We claim that $k^{\prime}\geq k$ and $p^{\prime}_{i}\leq p_{i}$ for all $i\leq k$.
+Assume for contradiction that $i\leq k$ is the smallest rank such that either $i>k^{\prime}$ or $p^{\prime}_{i}>p_{i}$.
+For OPT, at least $i$ jobs of size at most $p_{i}$ are completed by time $t$, so:
+$V_{\leq p_{i}}(t)\geq\sum_{j=1}^{i}p_{j}.$
+For NSJF, all started jobs of size $\leq p_{i}$ are among the first i1i-1 jobs (either because only i1i-1 jobs have started, or because the $i$-th smallest is strictly larger than $p_{i}$). Therefore:
 $V^{\prime}_{\leq p_{i}}(t+\tau)\leq\sum_{j=1}^{i-1}p^{\prime}_{j}\leq\sum_{j=1}^{i-1}p_{j},$
-where the second inequality uses the assumption that $p^{\prime}_{j}\leq p_{j}$ for all $j<i$.
-Combining these two, we get: $V^{\prime}_{\leq p_{i}}(t+\tau)<V_{\leq p_{i}}(t),$ which contradicts the earlier volume bound from Lemma 3.
-Hence, our assumption must be false, and we conclude: $p^{\prime}_{i}\leq p_{i}$ for all $i$.
-Finally, this implies that the number of started jobs with $p_{j}\leq p$ by time $t+\tau$ under NSJF is at least the number of completed jobs with $p_{j}\leq p$ by time $t$ under OPT.
-Since all jobs started by time $t+\tau$ must be completed by time $t+2\tau$, it completes the proof of the lemma.
+where the second inequality uses $p^{\prime}_{j}\leq p_{j}$ for all $j<i$. This gives $V^{\prime}_{\leq p_{i}}(t+\tau)<V_{\leq p_{i}}(t)$, contradicting Lemma 3.
+Therefore $k^{\prime}\geq k$ and $p^{\prime}_{i}\leq p_{i}$ for all $i\leq k$, which implies that for every threshold $p$, the number of jobs of size $\leq p$ started by NSJF by time $t+\tau$ is at least the number completed by OPT by time $t$.
+Since all jobs started by time $t+\tau$ are completed by time $t+2\tau$, the lemma follows.
+∎
+Step 3: Bounding the Flow Time.
 Proof of lemma 2.
-Let $u$ denote the maximum completion time under NSJF. The flow times of NSJF and OPT can be expressed as:
-$\textup{{OPT}}=\int_{0}^{\infty}\left(|J(t)|-N(t)\right)\,dt,\quad F(\textup{{NSJF}})=\int_{0}^{u}\left(|J(t)|-N^{\prime}(t)\right)\,dt,$
+Let $u$ denote the maximum completion time under NSJF. The total flow times can be expressed as:
+$F(\textup{{OPT}})=\int_{0}^{\infty}\left(|J(t)|-N(t)\right)\,dt,\quad F(\textup{{NSJF}})=\int_{0}^{u}\left(|J(t)|-N^{\prime}(t)\right)\,dt,$
 where $|J(t)|$ is the number of jobs released by time $t$.
@@
 $\leq\int_{0}^{u-2\tau}\left(|J(t)|-N(t)\right)\,dt+2n\tau$
-$\leq\textup{{OPT}}+2n\tau.$
+$\leq F(\textup{{OPT}})+2n\tau.$
 4 Randomized Non-Preemptive Algorithm
@@
 On job $i$ release:
-   Classify job $i$ large or small based on Partition($i,\ell$) with $\ell=\lfloor\sqrt{n}\rfloor$ ;
+   Call Partition($i,\ell$) with $\ell=\lfloor\sqrt{n}\rfloor$ to classify job $i$ and update related jobs;
  if $i$ is classified as large then
@@
       Insert $j$ into $\S_{2}$ to start at $t$ ;
+    // Overlapping jobs are shifted right.
    end for
@@
 Theorem 1.
-There exists an online polynomial time randomized non-preemptive algorithm for total flow time minimization on the single-machine case, e.g., Algorithm 2, is $O(\sqrt{n})$-competitive, against the preemptive offline solution.
+Algorithm 2 is an online polynomial-time randomized non-preemptive algorithm for total flow time minimization on a single machine that is $O(\sqrt{n})$-competitive against the preemptive offline solution.
 Proof.
@@
 First, $\S_{1}$ is competitive with the optimal schedule for the job set $S$. By Lemma 1, all jobs in $S$ have size at most $\textup{{OPT}}/\lfloor\sqrt{n}\rfloor$. Therefore, by Lemma 2, we have:
-$F(\S_{1})\leq O(\sqrt{n})\cdot F(\textup{{OPT}}(S)).$
-Moreover, since $S$ is weaker than $J$ by Lemma 1, we further have:
-$F(\S_{1})\leq O(\sqrt{n})\cdot F(\textup{{OPT}}(S))\leq O(\sqrt{n})\cdot\textup{{OPT}}.$ (1)
+$F(\S_{1})=O(\sqrt{n})\cdot F(\textup{{OPT}}(S)).$
+Moreover, since $S$ is weaker than $J$ by Lemma 1, $F(\textup{{OPT}}(S))\leq\textup{{OPT}}$, and thus:
+$F(\S_{1})=O(\sqrt{n})\cdot\textup{{OPT}}.$ (1)
 It remains to analyze the additional flow time incurred during the transition from $\S_{1}$ to $\S_{2}$ due to the insertion of large jobs.
@@
 Any subsequent increases to this flow time are charged to the external delays ($\delta_{j^{\prime}}$) caused by later jobs.
-Let $\hat{F}_{j}$ denote the self-delay of $j$. We define:
+Let $\hat{F}_{j}$ denote the self-delay of $j$:
 […]
-where $C^{\prime}_{j}$ is the completion time of $j$ in $\S_{2}$ immediately after its insertion (in the last update round of our algorithm). In Lemma 7, we show:
+where $\hat{C}_{j}$ is the completion time of $j$ in $\S_{2}$ immediately after its insertion (in the last update round of our algorithm). In Lemma 7, we show:
 $\sum_{j\in L}\hat{F}_{j}=O(\sqrt{n})\cdot\textup{{OPT}}.$ (3)
@@
 Lemma 6 (External Delay).
-For each large job $j$, the additional flow time arising from its delay of other previous jobs is bounded by $O(\sqrt{n})\cdot p_{j}$ in expectation, i.e., $\mathbb{E}(\delta_{j})\leq O(\sqrt{n})\cdot p_{j}$.
+For each large job $j$, the additional flow time arising from its delay of other previous jobs is bounded by $O(\sqrt{n})\cdot p_{j}$ in expectation, i.e., $\mathbb{E}(\delta_{j})=O(\sqrt{n})\cdot p_{j}$.
 Proof.
-Let $D_{k}$ be the set of jobs delayed by job $j$ when $w_{j}=k$.
-We aim to show that $\mathbb{E}[|D_{k}|]=O(\sqrt{n})$.
-The key claim is that for each job $i\neq j$, it appears in at most one $D_{k}$.
-This is because there is exactly $p_{j}$ idle time between any two adjacent insertion locations, since we collect $p_{j}$ units of idle time between them.
+Let $D_{k}$ be the set of jobs delayed by large job $j$ conditioned on the randomness $w_{j}=k$.
+To show that $\mathbb{E}[|D_{w_{j}}|]=O(\sqrt{n})$, the key claim is that every job $i\penalty 10000\ (i\neq j)$ appears in at most one of $D_{1},D_{2},\dots,D_{\lfloor\sqrt{n}\rfloor}$.
+This is because when inserting job $j$, there is exactly $p_{j}$ idle time between any two adjacent insertion locations.
 This prevents a single job from being delayed by more than one possible insertion position of $j$.
 Therefore, we have:
-$\mathbb{E}[|D_{k}|]=\sum_{k=1}^{\sqrt{n}}|D_{k}|\cdot\Pr[w_{j}=k]\leq n\cdot\frac{1}{\lfloor\sqrt{n}\rfloor}\leq O(\sqrt{n}).$
+$\mathbb{E}[|D_{w_{j}}|]=\sum_{k=1}^{\lfloor\sqrt{n}\rfloor}|D_{k}|\cdot\Pr[w_{j}=k]\leq n\cdot\frac{1}{\lfloor\sqrt{n}\rfloor}=O(\sqrt{n}).$
 Thus, the expected number of jobs delayed by $j$ is $O(\sqrt{n})$, and since each of them is delayed by at most $p_{j}$, this implies the lemma. ∎
@@
 Proof.
-For a fixed large job $j$, the self-delay is either $r_{j^{\prime}}-r_{j}$, where $j^{\prime}$ is the proxied virtual job, or $C^{\prime}_{j}-r_{j}$ if $j$ is unproxied.
-Both types of self-delay arise from only two sources: collecting idle time or passing through busy time.
-First, consider the idle time collected. Since $w_{j}\leq\sqrt{n}$, the total idle time accumulated for each large job $j$ is at most $\sqrt{n}\cdot p_{j}$. Summing over all large jobs, we have:
-$\sum_{j\in L}\text{(idle contribution to self-delay)}\leq\sqrt{n}\cdot\sum_{j\in L}p_{j}\leq\sqrt{n}\cdot\textup{{OPT}}.$
-Next, we analyze the busy time passed through by jobs.
-This may occur during the intervals $r_{j^{\prime}},r_{j})$ (for proxied jobs) or $r_{j},C^{\prime}_{j})$ (for unproxied jobs).
+For a fixed large job $j$, the self-delay is either $r_{j^{\prime}}-r_{j}$ (if proxied by $j^{\prime}$) or $\hat{C}_{j}-r_{j}$ (if unproxied).
+We decompose the self-delay into three sources: collecting idle time, passing through busy time, and the job’s own processing time $p_{j}$.
+First, consider the idle time collected. Since $w_{j}\leq\sqrt{n}$, the total idle time accumulated for each large job $j$ is at most $\sqrt{n}\cdot p_{j}$. Summing over all large jobs:
+$\sum_{j\in L}\text{(idle contribution)}\leq\sqrt{n}\cdot\sum_{j\in L}p_{j}\leq\sqrt{n}\cdot\textup{{OPT}}.$
+Next, we analyze the busy time passed through.
+For each large job $j$, this occurs during the interval $r_{j},r_{j^{\prime}})$ (if proxied) or $r_{j},\hat{s}_{j})$ (if unproxied, where $\hat{s}_{j}$ denotes the start time of $j$ immediately after its insertion).
 In both cases, the large job remains unproxied throughout the interval.
-By Lemma 1, the number of unproxied large jobs at any time $t$ is at most $\sqrt{n}$.
-Since the total length of busy intervals is at most the total processing time of all jobs, we have:
-$\sum_{j\in L}\text{(busy-time contribution to self-delay)}\leq\sqrt{n}\cdot\textup{{OPT}}.$
-Combining both contributions, the total self-delay across all large jobs is bounded by
+By Lemma 1, the number of active unproxied large jobs at any time $t$ is at most $\sqrt{n}$, and there is at most one committed unproxied large job at any time $t$.
+Since the total length of busy intervals is at most OPT:
+$\sum_{j\in L}\text{(busy-time contribution)}\leq(\sqrt{n}+1)\cdot\textup{{OPT}}.$
+Finally, the processing time contribution is $\sum_{j\in L}p_{j}\leq\textup{{OPT}}$. Combining all three:
 $\sum_{j\in L}\hat{F}_{j}=O(\sqrt{n})\cdot\textup{{OPT}}.$
@@
 In the single-machine setting, each job has $\sqrt{n}$ insertion choices, resulting in an expected delay affecting $\sqrt{n}$ jobs.
-In the multi-machine setting, we allocate $\sqrt{n/m}$ choices to $w_{j}$ and $m$ choices to $m_{j}$, resulting in an expected delay affecting only $\sqrt{n/m}$ jobs.
+In the multi-machine setting, we allocate $\sqrt{n/m}$ choices to $w_{j}$ and $m$ choices to $m_{j}$, resulting in an expected delay affecting only $O(\sqrt{n/m})$ jobs.
   • Self-Delay (Idle Time): The self-delay incurred by collecting idle time is at most $\sqrt{n/m}\cdot p_{j}$, due to the smaller upper bound on $w_{j}$.
@@
 5 Deterministic Non-Preemptive Algorithm
-This section introduces a deterministic scheduling algorithm that operates without preemption or restart.
-In the single-machine setting, the trivial $O(n)$-competitive algorithm is known to be tight.
-However, when the number of machines $m$ is no longer constant, this $O(n)$ bound becomes loose.
-Our algorithm achieves a competitive ratio of $O(n/m^{2}+\sqrt{n/m}\log m)$.
-The algorithm builds on the online rank-based partitioning framework introduced in Algorithm 1.
-Upon the arrival of a job $j$, it is classified as either large or small according to the rule specified in Algorithm 1 with $\ell=\lfloor\sqrt{nm}\rfloor$.
-Recall that this partitioning framework not only distinguishes between large and small jobs but may also release proxy jobs and mark certain jobs as proxied.
-Machines are assigned specific roles as follows:
-  • Machine 1: Reserved for large jobs.
-  • Mixed machines (Machines $2$ to $\lceil m/2\rceil+1$): Accept both small and large jobs.
-  • Small-Only machines (Machines $\lceil m/2\rceil+2$ to $m$): Restricted to small jobs only.
-Note that not all Mixed machines are available to large jobs at all times.
-Instead, we adopt a fine-grained policy to dynamically determine the number of active Mixed machines dedicated to large jobs.
-This is generated by a function $\gamma$, which depends on $k$, the number of active large jobs still waiting.
-Roughly speaking, at any given time, we activate $\gamma(k)$ Mixed machines to serve large jobs, where:
-$\gamma(k)=\left\lfloor\frac{k}{\sqrt{n/m}}\right\rfloor.$
-Finally, the scheduling policy follows a generalized version of NSJF, respecting the machine roles defined above:
-  • If Machine 1 is idle, it schedules the smallest waiting large job, if available.
-  • If a Mixed machine is idle, it first schedules the smallest waiting small job, if one is available.
-Otherwise, let $\theta$ be the number of Mixed machines currently processing large jobs.
-If $\theta<\gamma(k)$, where $k$ is the number of active large jobs still waiting, the algorithm schedules the smallest such job.
-  • If a Small-Only machine is idle, it schedules the smallest waiting small job, if available.
+This section presents a deterministic non-preemptive algorithm with a competitive ratio of $O(n/m^{2}+\sqrt{n/m}\log m)$.
+This nearly closes the gap between the trivial $O(n/m)$ upper bound and the $\Omega(n/m^{2})$ lower bound of [10], particularly when $m$ is large relative to $n$.
+The algorithm builds on the online rank-based partitioning framework (Algorithm 1) with $\ell=\lfloor\sqrt{nm}\rfloor$.
+Based on this classification, machines are assigned different roles to handle large and small jobs separately, as described below.
+The first machine, called the Large-Only machine, is dedicated exclusively to large jobs and schedules the smallest waiting large job if available.
+For the remaining Mixed machines $2\dots m$, a dynamic threshold function
+$\gamma(k)=\min\left\{\left\lfloor\frac{k}{\sqrt{n/m}}\right\rfloor,\lceil m/2\rceil\right\}.$
+controls how many of them may process large jobs, where $k$ is the number of active large jobs currently waiting.
+When one of these Mixed machines becomes idle, it schedules the smallest waiting job regardless of its size if the number of Mixed machines currently processing large jobs is less than $\gamma(k)$; otherwise, it schedules small jobs only using NSJF.
+Since there are at most $\lceil m/2\rceil$ Mixed machines running large jobs at any time, there are always $m-1-\lceil m/2\rceil$ machines exclusively dedicated to small jobs.
 We use $J$ to denote the real job set, which includes only real jobs and excludes proxy jobs.
@@
 Lemma 8 (Large Jobs).
-$\sum_{j\in L}\hat{F}_{j}\leq O(\sqrt{n/m})\cdot\textup{{OPT}}.$
+$\sum_{j\in L}\hat{F}_{j}=O(\sqrt{n/m})\cdot\textup{{OPT}}.$
 Proof.
-We first bound the waiting time component $r_{j},s_{j})$ (or $r_{j},r_{j^{\prime}})$ if proxied) of the flow time which does not include the $p_{j}$ part.
-At any time $t$ in this interval, $j$ is not scheduled because all active Mixed machines are busy.
-Let $\theta(t)$ be the number of Mixed machines currently processing jobs, and let $k(t)$ be the number of active large jobs waiting at time $t$.
-By the scheduling rule, we either have
-$\gamma(k(t))\leq\theta(t)+1,\quad\text{or}\quad\theta(t)=\lceil m/2\rceil.$
-In the second case, since $k(t)$ is at most $\lfloor\sqrt{nm}\rfloor$, and using the definition $\gamma(k)=\left\lfloor\frac{k}{\sqrt{n/m}}\right\rfloor$, we get:
-$k(t)\leq O(\sqrt{n/m})\cdot(\theta(t)+1).$
-Also, note that while a large job is waiting, the large-only machine (Machine 1) must be busy. Letting $\mathbb{I}_{k(t)\geq 1}(t)$ denote the indicator function for whether $k(t)\geq 1$, we have:
+Consider the waiting time of a large job $j$ during the interval $r_{j},s_{j})$ (or $r_{j},r_{j^{\prime}})$ if proxied).
+At any time $t$ in this interval, $j$ is an active large job that has not been scheduled, either because all Mixed machines are busy, or because at least $\gamma(k(t))$ of them are already processing large jobs.
+Let $\theta(t)$ denote the number of busy Mixed machines at time $t$ (processing any job, large or small), and let $k(t)$ denote the number of active large jobs waiting at time $t$.
+Since at least $\gamma(k(t))$ of these machines are processing large jobs, we have by the scheduling rule:
+$\theta(t)=m-1\quad\text{or}\quad\gamma(k(t))\leq\theta(t).$
+By the definition of $\gamma$, this gives $k(t)=O(\sqrt{n/m})\cdot(\theta(t)+1)$.
+Whenever $k(t)\geq 1$, the Large-Only machine must also be busy, so the total number of busy machines is at least $\theta(t)+1$. Since each busy machine contributes to the total processing time:
 $\int_{0}^{\infty}(\theta(t)+1)\cdot\mathbb{I}_{k(t)\geq 1}(t)\,dt\leq\sum_{j\in J}p_{j}.$
-Combining all these elements and reincorporating $p_{j}$, we have:
+Combining with the processing time of the jobs themselves:
 $\sum_{j\in L}\hat{F}_{j}$ $\leq\sum_{j\in L}p_{j}+\int_{0}^{\infty}k(t)\,dt$
 $\leq\sum_{j\in L}p_{j}+O(\sqrt{n/m})\cdot\int_{0}^{\infty}(\theta(t)+1)\cdot\mathbb{I}_{k(t)\geq 1}(t)\,dt$
-$\leq\sum_{j\in L}p_{j}+O(\sqrt{n/m})\cdot\sum_{j\in J}p_{j}$
-$=O\left(\sqrt{\frac{n}{m}}\right)\cdot\textup{{OPT}}.$
+$\leq\sum_{j\in L}p_{j}+O(\sqrt{n/m})\cdot\sum_{j\in J}p_{j}=O\left(\sqrt{\frac{n}{m}}\right)\cdot\textup{{OPT}}.$
 Next, we analyze the contribution of small jobs.
-The first loss occurs because of the large-only machine.
-Ideally, even without any large job, our algorithm is a (m1)(m-1)-machine NSJF algorithm, which is competitive to the (m1)(m-1)-machine preemptive optimal solution.
-However, the (m1)(m-1)-machine preemptive optimal may already be larger than the $m$-machine optimal solution.
-To address it, we give the following lemma.
+Since the Large-Only machine is reserved for large jobs, small jobs are effectively scheduled on m1m-1 machines.
+To compare against the $m$-machine optimum, we first bound the cost of losing one machine.
 Lemma 9.
-For every $0\leq k\leq m-1$, letting $\textup{{OPT}}_{m-k}$ be the optimal preemptive solution only using mkm-k machines, we have $\textup{{OPT}}_{m-k}=O\left(\frac{nk}{m(m-k)}\right)\cdot\textup{{OPT}}_{m}$.
-Before proceeding, we recall that the preemptive optimal solutions discussed here, $\textup{{OPT}}_{m}$ and $\textup{{OPT}}_{m-k}$, operate under preemptive model without migration.
+For every $0\leq k\leq m-1$, the optimal preemptive flow time on mkm-k machines satisfies $\textup{{OPT}}_{m-k}\leq\left(1+\frac{nk}{m(m-k)}\right)\cdot\textup{{OPT}}_{m}$, where both optima are under the preemptive model without migration.
 Proof.
-Consider a preemptive optimal solution with $m$ machines.
-We will use it to construct a feasible solution using mkm-k machines and prove that its total flow time increases by at most $\frac{nk}{m(m-k)}\cdot\textup{{OPT}}_{m}$.
-This will complete the proof of the lemma.
-First, we sort the machines in the optimal solution $\textup{{OPT}}_{m}$ by their total workloads.
-Let $J_{i}$ denote the set of jobs scheduled on machine $i$ in $\textup{{OPT}}_{m}$, and let $J_{[i]}=\bigcup_{h=1}^{i}J_{h}$ denote the union of jobs on the first $i$ machines.
-Machines are sorted so that
-$\sum_{j\in J_{1}}p_{j}\leq\sum_{j\in J_{2}}p_{j}\leq\cdots\leq\sum_{j\in J_{m}}p_{j}.$
-We consider the first $k$ machines and reassign all jobs in $J_{[k]}$ to the remaining mkm-k machines in arbitrary order. By construction, we have:
-$\sum_{j\in J_{[k]}}p_{j}\leq\frac{k\sum_{j\in J}p_{j}}{m}.$
-The reassignment proceeds as follows.
-For each job $j\in J_{[k]}$, we preserve its starting time $t$ from $\textup{{OPT}}_{m}$, and insert it at the same time $t$ on one of the machines among machines mk+1m-k+1 to $m$.
-As a result, the existing jobs on that machine are delayed by at most $p_{j}$.
-To minimize the impact, we always place $j$ on the machine with the fewest currently assigned jobs.
-Since there are mkm-k machines and at most $n$ jobs, this ensures the number of jobs on that machine is at most n/(mk)n/(m-k).
-Therefore, the additional cost due to delaying existing jobs is at most $np_{j}/(m-k)$.
-On the other hand, since we insert job $j$ at its original start time, its flow time is at most what it was in $\textup{{OPT}}_{m}$.
-Even though future delays may increase $j$’s flow time, these increases are charged to the delay costs of subsequent jobs.
-Thus, the increase in total flow time due to inserting a single job $j$ compared to $\textup{{OPT}}_{m}$ is at most
-$\frac{np_{j}}{m-k}.$
-Summing over all $j\in J_{[k]}$, the total increase in flow time is at most
-$\frac{n\sum_{j\in J_{[k]}}p_{j}}{m-k}\leq\frac{nk\sum_{j\in J}p_{j}}{m(m-k)}\leq\frac{nk\cdot\textup{{OPT}}_{m}}{m(m-k)}.$
-In conclusion, we obtain the desired bound:
-$\textup{{OPT}}_{m-k}\leq\left(1+\frac{nk}{m(m-k)}\right)\cdot\textup{{OPT}}_{m}.$
-As in Lemma 12, we illustrate the ideas again here for completeness.
-We partition the timeline into intervals based on when large jobs are started.
-Let $t_{h}$ denote the start time of the $h$-th large job, with $t_{0}=0$.
-Let $S_{h}$ be the set of small jobs released during the interval $t_{h-1},t_{h})$.
-Let $\textup{{OPT}}_{m-1}(S_{h})$ denote the optimal total flow time for scheduling $S_{h}$ on m1m-1 machines. Since the intervals are disjoint, we have:
-$\textup{{OPT}}_{m-1}\geq\sum_{h}F(\textup{{OPT}}_{m-1}(S_{h})).$
-We compare our algorithm’s flow time on $S_{h}$ to $\textup{{OPT}}_{m-1}(S_{h})$.
-During $t_{h-1},t_{h})$, no new large jobs are started, and all jobs in $S_{h}$ arrive after $t_{h-1}$.
-The only difference from running $\textup{{NSJF}}(S_{h})$ on a fresh instance is that each machine may still be processing a job that started before $t_{h-1}$.
-We refer to these as initial blocking jobs, and define the corresponding initial blocking period on machine $i$ as $t_{h-1},t_{h-1}+b_{i})$, where $b_{i}$, the initial blocking time is the remaining processing time at $t_{h-1}$.
-Thus, our algorithm behaves like NSJF on m1m-1 machines with such initial blocking times periods.
-This behavior is captured by Lemma 29.
-To apply it, we define a shifted instance $S^{\prime}_{h}$ by translating the time origin to $t_{h-1}$, i.e., replacing each release time $r_{j}$ with $r_{j}-t_{h-1}$.
-Then, the flow time of each job $j\in S_{h}$ under our algorithm matches its flow time in $\textup{{NSJF}}(S^{\prime}_{h},\vec{b})$, where each machine has initial blocking interval $0,b_{i})$.
-Lemma 10.
-For each small-job period $(t_{h-1},t_{h}]$, the total length of initial blocking time satisfies:
-$\sum_{i=1}^{m-1}b_{i}\leq O\left(\sqrt{\frac{m}{n}}\log m\right)\cdot\textup{{OPT}}.$
+Starting from an optimal preemptive schedule on $m$ machines, we construct a feasible schedule on mkm-k machines.
+Sort the machines by total workload and let $J_{[k]}$ denote the jobs on the $k$ lightest machines.
+Their total processing time satisfies $\sum_{j\in J_{[k]}}p_{j}\leq k\sum_{j\in J}p_{j}/m$.
+Each job $j\in J_{[k]}$ is reassigned to one of the remaining mkm-k machines, inserted at its original start time.
+This delays existing jobs on the target machine by at most $p_{j}$.
+To minimize impact, each job is placed on the machine with the fewest currently assigned jobs, ensuring at most n/(mk)n/(m-k) jobs per machine.
+The additional flow time from inserting $j$ is therefore at most $np_{j}/(m-k)$.
+Summing over all reassigned jobs:
+$\textup{{OPT}}_{m-k}\leq\textup{{OPT}}_{m}+\frac{n\sum_{j\in J_{[k]}}p_{j}}{m-k}\leq\textup{{OPT}}_{m}+\frac{nk}{m(m-k)}\cdot\textup{{OPT}}_{m}.$
+Our key tool for analyzing small jobs is the generalized NSJF analysis (Lemma 29, proved in Appendix A), which extends Lemma 2 to handle initial blocking periods.
+If NSJF runs on $m^{\prime}$ machines where machine $i$ is unavailable during $0,b_{i})$, and all job sizes are bounded by $\tau$, then
+$F(\textup{{NSJF}}(\vec{b}))\leq F(\textup{{OPT}}(\vec{0}))+2n\tau+\frac{nB}{m^{\prime}},$
+where $B=\sum_{i}b_{i}$ is the total blocking time and $\textup{{OPT}}(\vec{0})$ is the optimal preemptive schedule on $m^{\prime}$ machines without blocking.
+To apply this, we partition the timeline into intervals $t_{h-1},t_{h})$, where $t_{h}$ is the start time of the $h$-th large job on an Mixed machine (excluding the Large-Only machine) and $t_{0}=0$.
+Let $S_{h}$ denote the small jobs released during $t_{h-1},t_{h})$.
+Since no Mixed machine starts a new large job during $t_{h-1},t_{h})$, all m1m-1 Mixed machines are available for small jobs, and every job in $S_{h}$ is scheduled before $t_{h}$.
+Some machines may be initially blocked by jobs still in progress from the previous interval: machine $i$ is unavailable during $t_{h-1},t_{h-1}+b_{h,i})$, where $b_{h,i}$ is its remaining processing time at $t_{h-1}$.
+Shifting the time origin to $t_{h-1}$, the flow time of $S_{h}$ under our algorithm matches that of $\textup{{NSJF}}(S^{\prime}_{h},\vec{b_{h}})$, where $S^{\prime}_{h}$ is the shifted instance and $\vec{b_{h}}$ is the blocking vector.
+It remains to bound the total blocking time.
+Lemma 10 (Blocking Time).
+For each interval $(t_{h-1},t_{h}]$, the total initial blocking time satisfies:
+$\sum_{i=1}^{m-1}b_{i}=O\left(\sqrt{\frac{m}{n}}\log m\right)\cdot\textup{{OPT}}.$
 Proof.
-We split the analysis based on whether the blocking job is small or large.
-We first consider the blocking caused by small jobs.
-By Lemma 1, each small job has size at most $O\left(\frac{\textup{{OPT}}}{\sqrt{nm}}\right)$.
-Since at most $m$ machines can be blocked, the total blocking time from small jobs is:
-$O\left(\frac{\textup{{OPT}}}{\sqrt{nm}}\right)\cdot m=O\left(\sqrt{\frac{m}{n}}\cdot\textup{{OPT}}\right).$
-For the blocking caused by large jobs, we proceed as follows.
-For large jobs causing blocking, let $p_{i}$ be the size of the $i$-th large job that causes blocking.
-When it starts, at least i1i-1 Mixed machines are already processing large jobs, so it must satisfy $\gamma(k)\geq i$ for the number of active large jobs still waiting $k$.
-Using $\gamma(k)=\lfloor k/(\sqrt{n/m})\rfloor$, we get $k\geq i\sqrt{n/m}$, and thus:
-$p_{i}\leq\frac{\textup{{OPT}}}{k}\leq\frac{\textup{{OPT}}}{i\sqrt{n/m}}\leq\frac{1}{i}\cdot O\left(\sqrt{\frac{m}{n}}\cdot\textup{{OPT}}\right).$
-Summing over $i=1$ to $m$, the total blocking from large jobs is:
-$\sum_{i=1}^{m}p_{i}\leq O\left(\sqrt{\frac{m}{n}}\cdot\textup{{OPT}}\right)\cdot\sum_{i=1}^{m}\frac{1}{i}=O\left(\sqrt{\frac{m}{n}}\log m\cdot\textup{{OPT}}\right).$
-Adding both parts completes the proof. ∎
-With the bound on the initial blocking times $b_{i}$, we can apply the generalized NSJF analysis in Lemma 29.
-This lemma allows us to compare the performance of NSJF with initial blocking times against the optimal schedule without any blocking.
-We now prove the main result.
+At time $t_{h-1}$, each Mixed machine is blocked by at most one job still in progress. We bound the total blocking separately for small and large blocking jobs.
+Small blocking jobs.
+Each small job has size at most $O(\textup{{OPT}}/\sqrt{nm})$ by Lemma 1. Since at most m1m-1 machines can be blocked, the total blocking from small jobs is at most $O(\sqrt{m/n}\cdot\textup{{OPT}})$.
+Large blocking jobs.
+Order the large jobs that are still running on Mixed machines at time $t_{h-1}$ by their start times, and let $p_{i}$ denote the size of the $i$-th such job.
+When the $i$-th job was started, at least i1i-1 other Mixed machines were already processing large jobs, so the scheduling rule required $\gamma(k)\geq i$, where $k$ is the number of active large jobs waiting at that moment.
+By the definition of $\gamma$, this implies $k\geq i\sqrt{n/m}$.
+Since these $k$ waiting jobs are all at least as large as $p_{i}$ (otherwise they would have been scheduled first), we have $k\cdot p_{i}\leq\textup{{OPT}}$, giving:
+$p_{i}\leq\frac{\textup{{OPT}}}{i\sqrt{n/m}}=\frac{1}{i}\cdot O\left(\sqrt{\frac{m}{n}}\right)\cdot\textup{{OPT}}.$
+There are at most $\lceil m/2\rceil$ such jobs, so the total blocking from large jobs is:
+$\sum_{i=1}^{\lceil m/2\rceil}p_{i}\leq O\left(\sqrt{\frac{m}{n}}\right)\cdot\textup{{OPT}}\cdot\sum_{i=1}^{\lceil m/2\rceil}\frac{1}{i}=O\left(\sqrt{\frac{m}{n}}\log m\right)\cdot\textup{{OPT}}.$
+Combining both parts gives the claimed bound. ∎
+With the blocking time bounded, Lemma 29 applies directly. We now prove the main result.
 Theorem 2.
-There exists an online polynomial time deterministic non-preemptive algorithm for online total flow time minimization with competitive ratio $O(n/m^{2}+\sqrt{n/m}\log m)$, against the preemptive offline optimal solution.
+There exists an online polynomial-time deterministic non-preemptive algorithm for total flow time minimization that is $O(n/m^{2}+\sqrt{n/m}\log m)$-competitive against the preemptive offline solution.
 Proof.
@@
 Let $\tau$ be the size upper bound for small jobs, i.e., $\tau=O(\textup{{OPT}}/\sqrt{nm})$.
-For each small-job segment $t_{h-1},t_{h})$, our algorithm performs the same as $\textup{{NSJF}}(S^{\prime}_{h})$ with initial blocking vector $\vec{b}$, denoted by $F_{j}(\textup{{NSJF}}(S^{\prime}_{h},\vec{b}))$.
-By Lemma 29, and letting $B=\sum_{i}b_{i}$, we have:
-$\sum_{j\in S^{\prime}_{h}}F_{j}(\textup{{NSJF}}(S^{\prime}_{h},\vec{b}))\leq\textup{{OPT}}_{m-1}(S_{h})+2|S_{h}|\tau+\frac{|S_{h}|B}{m}.$
+For each small-job segment $t_{h-1},t_{h})$, our algorithm performs the same as $\textup{{NSJF}}(S^{\prime}_{h})$ with initial blocking vector $\vec{b_{h}}$.
+By Lemma 29 on m1m-1 machines, let $B_{h}=\sum_{i}b_{h,i}$:
+$\sum_{j\in S_{h}}F_{j}\leq F(\textup{{OPT}}_{m-1}(S_{h}))+2|S_{h}|\tau+\frac{|S_{h}|B_{h}}{m-1}.$
 Summing over all disjoint small-job segments:
-$\sum_{j\in S}F_{j}\leq F(\textup{{OPT}}_{m-1})+2n\tau+\frac{nB}{m}.$
-Using the bounds $\tau=O(\textup{{OPT}}/\sqrt{nm})$ and $B=O(\sqrt{m/n}\log m\cdot\textup{{OPT}})$, we get:
+$\sum_{j\in S}F_{j}\leq F(\textup{{OPT}}_{m-1})+2n\tau+\frac{n\cdot\max_{h}B_{h}}{m-1}.$
+Using the bounds $\tau=O(\textup{{OPT}}/\sqrt{nm})$ and $\max_{h}B_{h}=O(\sqrt{m/n}\log m\cdot\textup{{OPT}})$, we get:
 $\sum_{j\in S}F_{j}\leq\textup{{OPT}}_{m-1}+O\left(\sqrt{\frac{n}{m}}\right)\cdot\textup{{OPT}}+O\left(\sqrt{\frac{n}{m}}\log m\right)\cdot\textup{{OPT}}.$
@@
 6 Deterministic Algorithm with Kill-and-Restart
-To utilize the power of kill-and-restart without relying on randomization, we first introduce a fine-grained online partitioning framework with the following modification (see Algorithm 3 for details, and note that the threshold will be set to $\sqrt{nm}$):
-  • Modified partition rule: Let $P$ be the total size of all previously released jobs (excluding proxy jobs), which serves as a lower bound on OPT.
-We classify a new job $j$ as a small job if $p_{j}\leq 4P/\ell$, rather than using the large job queue as in the original rule.
-The modification is motivated by the need for a size-based lower bound in the restart setting, as opposed to a purely rank-based one.
-For instance, we aim to avoid killing jobs that are too small.
-These changes are mild and preserve all the original properties of the online partitioning framework.
+To utilize the power of kill-and-restart without relying on randomization, we first introduce a refined online partitioning framework with the following modification (see Algorithm 3 for details, and note that the threshold will be set to $\sqrt{nm}$):
+Modified partition rule.
+Let $P$ be the total size of all previously released jobs (excluding proxy jobs), which serves as a lower bound on OPT.
+We classify a new job $j$ as a small job if $p_{j}\leq 4P/\ell$, and large otherwise.
+The other procedures remain identical to Algorithm 1.
+This modification preserves the original properties of the framework.
+The additional size filter provides a lower bound on the processing time of large jobs, which will be used in the analysis of kill events (1).
 Data: Active large job set $A$, and every job’s other current states
-Function Partition($i,\ell$):
- // New Rule: An absolute smallness rule to filter out tiny jobs.
+Procedure Partition($i,\ell$):
    Let $P$ be the sum of processing times of all (original) jobs released so far;
  if $p_{i}\leq 4P/\ell$ then
-    return $i$ is small
- else if $|A|<\ell$ then
-      Add $i$ to $A$;
-    return $i$ is large
- else if $p_{i}>\min_{j\in A}p_{j}$ then
+      Mark $i$ as small;
+ else
+    if $|A|=\ell$ then
          Let $j=\arg\min_{j\in A}p_{j}$;
-      Remove $j$ from $A$ and add $i$ to $A$;
+       // $j$ will be retired.
+         Remove $j$ from AA;
        if job $j$ is still waiting then
             Create a new proxy job $j^{\prime}$ with $r_{j^{\prime}}=r_{i}$ and $p_{j^{\prime}}=p_{j}$;
+            Mark $j^{\prime}$ as small;
             Mark $j$ as proxied by $j^{\prime}$;
-       // The scheduler will actually process $j$ when it processes $j^{\prime}$.
+          // The scheduler will actually process $j$ when processing $j^{\prime}$.
        else
-       // $j$ is a committed large job.
+            Mark $j$ as committed large;
          end if
-    return $i$ is large
- else
-    return $i$ is small
+      end if
+     Add $i$ to AA;
+      Mark $i$ as large;
    end if
@@
 Lemma 11.
-Under the online rank-based partitioning framework (fine-grained for restart), the following properties hold:
+Under the online rank-based partitioning framework (refined for restart), the following properties hold:
   1. 1.
-    The number of large jobs is at most $\ell$ at any time.
+    The number of active large jobs (the set $A$) is at most $\ell$ at any time.
   2. 2.
-    Let $S$ be the set of small jobs, including proxy jobs. Then $S$ is strictly weaker than the full job set $J$, meaning $F(\textup{{OPT}}(S))\leq F(\textup{{OPT}}(J))$.
+    The processing time of any job classified as small, or any retired large job, is upper bounded by $O(\textup{{OPT}}/\ell)$.
   3. 3.
-    The size of any small job, including any retired large job, is upper bounded by $O(\textup{{OPT}}/\ell)$.
+    Let $S$ be the set of all jobs ever classified as small, including all proxy jobs.
+Let $J$ be the original job set.
+The instance $S$ is no harder to schedule than $J$, meaning $F(\textup{{OPT}}(S))\leq F(\textup{{OPT}}(J))$.
   4. 4.
-    New size-based lower bound: The size of any large job is at least $4P(r_{j})/\ell$, where $P(r_{j})$ denotes the total processing time of all jobs released up to $r_{j}$, excluding proxy jobs.
+    The processing time of any large job $j$ is at least $4P(r_{j})/\ell$, where $P(r_{j})$ denotes the total processing time of all jobs released up to $r_{j}$, excluding proxy jobs.
 Proof.
@@
 ∎
-After introducing the fine-grained partitioning framework, we now describe the basic design of our scheduling policy, which divides the $m$ machines into two roles:
-  • Small-Only machines (Machines $1$ to $\lfloor m/2\rfloor$): These machines accept only small jobs.
-  • Mixed machines (Machines $\lfloor m/2\rfloor+1$ to $m$): These machines can process both small and large jobs.
-Our scheduling policy is a generalized version of the NSJF algorithm, incorporating the specific machine assignment rule described above.
-In particular, the algorithm makes scheduling decisions based on the partitioning produced by Algorithm 1.
-Each job is classified as either large or small, and further as retired or active, and proxied or unproxied.
-The scheduling rule proceeds as follows: the algorithm always selects the smallest unproxied waiting job to schedule. Additionally, Small-Only machines are restricted to processing only small jobs.
-The kill-and-restart rule applies only to active large jobs.
-Intuitively, an active large job should be killed only if it is blocking too many small jobs.
-We say that an active large job blocks a small job if it is processing at the small job’s release time.
-To detect such situations, the algorithm monitors the number of blocked small jobs.
-Specifically, if at any point the number of blocked small jobs reaches the threshold $\sqrt{nm}$, we kill all currently processing active large jobs.
-At the moment of a kill event, as well as when a new large job starts processing, we reset the counter tracking the number of blocked small jobs.
-With this, we complete the high-level description of our algorithmic design. The full pseudocode is presented in Algorithm 4.
+After introducing the refined partitioning framework, we now describe the basic design of our scheduling policy, which divides the $m$ machines into two roles:
+  • Small-Only machines ($1$ to $\lfloor m/2\rfloor$): These machines accept only small jobs.
+  • Mixed machines ($\lfloor m/2\rfloor+1$ to $m$): These machines accept both small and large jobs.
+Our scheduling policy builds on the partitioning produced by Algorithm 3.
+Each job is classified as either large or small; each large job is further classified as active or retired, and as proxied or unproxied.
+Scheduling rule.
+When an Mixed machine becomes idle, it schedules the smallest waiting unproxied job (regardless of whether it is large or small).
+When a Small-Only machine becomes idle, it schedules the smallest waiting small job.
+Kill-and-restart rule.
+We say that a newly released small job is blocked if any active large job is currently being processed, and unblocked otherwise.
+The algorithm maintains a counter $\phi$, which is incremented by one for each blocked small job and reset to zero whenever an active large job begins processing or is killed.
+When $\phi$ reaches $\lfloor\sqrt{nm}\rfloor$, the algorithm kills all currently processing active large jobs and resets $\phi$ to zero.
+The full pseudocode is presented in Algorithm 4.
 Include : Algorithm 3
-Data: Potential counter $\phi$ initiated as 0; every job’s current states as determined by Algorithm 3.
+Data: Blocked small job counter $\phi$ initialized to 0; every job’s current states as determined by Algorithm 3.
 On job $i$ release:
 […]
- if Job $i$ is classified as small then
 […]
-    $\phi\leftarrow\phi+1$ if $i$ is blocked ;
-    if $\text{$\phi$}=\lfloor\sqrt{nm}\rfloor$ then
+ if $j\neq\bot$, i.e., there exists a new small job $j$ then
+[…]
+    $\phi\leftarrow\phi+1$ if $j$ is blocked ;
+    if $\phi=\lfloor\sqrt{nm}\rfloor$ then
          Kill all processing active large jobs and $\phi\leftarrow 0$;
@@
        $j\leftarrow$ the smallest waiting unproxied job ;
-         Schedule $j$ on $i$ ;
-       $\phi\leftarrow 0$ if $j$ is large ;
+         Schedule $j$ on machine $i$ ;
+       $\phi\leftarrow 0$ if $j$ is an active large job ;
       end if
@@
 Algorithm 4 Deterministic Algorithm with Kill-and-Restart on $m$ Machines
+The basic invariant we aim to keep in the algorithm is Lemma 12.
+Lemma 12.
+The number of waiting blocked small jobs at any time is at most $\sqrt{nm}$.
+Proof.
+We partition the timeline into intervals, where each interval begins when the algorithm starts processing an active large job and ends just before the next such event.
+By the algorithm, the counter $\phi$ is reset to 0 at the start of each interval.
+Within an interval, every newly released blocked small job increments $\phi$ by one, and when $\phi$ reaches $\lfloor\sqrt{nm}\rfloor$, all currently processing active large jobs are killed and $\phi$ is reset.
+Therefore, at most $\lfloor\sqrt{nm}\rfloor$ blocked small jobs are created in each interval.
+It remains to show that all blocked small jobs created in one interval have started processing before the next interval begins.
+After a kill event (or if the active large jobs complete naturally), no active large job is being processed.
+At this point, every waiting blocked small job is smaller than every waiting active large job, so by the shortest-first scheduling rule, all blocked small jobs are scheduled before any active large job can start processing again.
+∎
 Moving into the analysis, we now distinguish between several types of job sets: The real job set $J$, which includes only actual jobs and excludes proxy jobs.
-The unblocked small job set $S$, and the blocked small job set $B$, both of which may include proxy jobs.
-The large job set $L$, which includes all large jobs, even if they become retired or are proxied at some point.
-Note that we do not have $B\cup S\cup L\subseteq J$ due to the presence of proxy jobs. We analyze the total flow time contributed by $S$, $B$, and $L$ separately.
+The unblocked small job set $S$, and the blocked small job set $K$, both of which may include proxy jobs.
+The large job set $L$, which includes all large jobs, even if they become retired at some point.
+Note that $K\cup S\cup L\not\subseteq J$ with the presence of proxy jobs. We analyze the total flow time contributed by $S$, $K$, and $L$ separately.
 Lemma 13 (Unblocked Small Job).
@@
 Proof.
-In the absence of blocked small jobs and retired large jobs, the flow time of all jobs in $S$ would be bounded by $F(\textup{{NSJF}}(S))$, which is at most $O\left(\sqrt{n/m}\right)\cdot F(\textup{{OPT}}(S))$ by Lemma 2.
-The reason is that the processing of active large jobs never affects the scheduling of unblocked small jobs.
-Whenever NSJF wants to schedule an unblocked small job $j$, the algorithm cannot have a large job smaller than $j$ outside of $S$.
-If such a large job is active, it must be larger than all small jobs by our online partition rule.
-If it is retired and still waiting, it is proxied and, therefore, never considered for scheduling.
-However, the presence of blocked small jobs and retired large jobs may delay the scheduling of unblocked small jobs, because we mix them together.
-We define a set $U$, consisting of the retired large jobs, and recall the set $B$, consisting of the blocked small jobs.
-Correspondingly, we construct two modified instances $U^{\prime}$ and $B^{\prime}$ by updating the release times of all jobs $j\in U$ and $j\in B$ to $s_{j}$; note that our algorithm’s schedule on $S\cup U\cup B$ is identical to that of $\textup{{NSJF}}(S\cup U^{\prime}\cup B^{\prime})$.
-By Lemma 2, we know that
-$F(\textup{{NSJF}}(S\cup U^{\prime}\cup B^{\prime}))\leq F(\textup{{OPT}}(S\cup U^{\prime}\cup B^{\prime}))+O(n\tau)\leq F(\textup{{OPT}}(S\cup U^{\prime}\cup B^{\prime}))+O(\sqrt{n/m})\cdot\textup{{OPT}}.$
-This alone does not prove the lemma, because the jobs in $U^{\prime}$ and $B^{\prime}$ may have later release times than in the original instance, so $F(\textup{{OPT}}(S\cup U^{\prime}\cup B^{\prime}))$ is not necessarily upper bounded by OPT.
-However, for every retired job $j\in U^{\prime}$, by the same reason as Lemma 1, it is retired after $s_{j}$, and we can find a newly released job $j^{\prime}$, where $p_{j^{\prime}}>p_{j}$, and $r_{j^{\prime}}>s_{j}$, where $s_{j}$ is the release time of $j$ in $U^{\prime}$.
-So, at least $S\cup U^{\prime}$ is a weaker set of $J$, and $F(\textup{{OPT}}(S\cup U^{\prime}))\leq\textup{{OPT}}$.
-We now aim to bound the difference between $F(\textup{{OPT}}(S\cup U^{\prime}\cup B^{\prime}))$ and $F(\textup{{OPT}}(S\cup U^{\prime}))$.
-To do so, we partition the timeline into intervals free of unproxied large jobs, denoted as $t_{h},t_{h+1})$ for $h=1,2,\dots$, where every $t_{h}$ can be viewed as the time we start a currently unproxied large job (it may be killed or retired later).
-In each interval $t_{h},t_{h+1})$, no unproxied large job is processed.
-Unblocked small jobs are only released in such intervals.
-Let $S_{h}$ be the set of unblocked small jobs with $r_{j}\in[t_{h},t_{h+1})$ and retired large jobs who becomes small in $t_{h},t_{h+1})$.
-Clearly, $\bigcup_{h}S_{h}=S\cup U^{\prime}$ by definition.
-We similarly define $B^{\prime}_{h}\subseteq B^{\prime}$ to be the jobs in $B^{\prime}$ whose processing falls within $t_{h},t_{h+1})$.
-Observe that, because only $B^{\prime}_{h}$ can affect $S_{h}$, we have
-$\sum_{j\in S_{h}}F_{j}(\textup{{NSJF}}(S\cup U^{\prime}\cup B^{\prime}))=\sum_{j\in S_{h}}F_{j}(\textup{{NSJF}}(S_{h}\cup B^{\prime}_{h}))\leq F(\textup{{NSJF}}(S_{h}\cup B^{\prime}_{h})).$
-An important property of our algorithm is that $|B^{\prime}_{h}|\leq\sqrt{nm}$; otherwise, we would have killed all unproxied jobs before time $t_{h}$.
-We will use this property to bound $F(\textup{{NSJF}}(S_{h}\cup B^{\prime}_{h})$.
-Now consider an optimal preemptive solution $\textup{{OPT}}(S_{h})$.
-We insert the jobs in $B^{\prime}_{h}$ into $\textup{{OPT}}(S_{h})$ one by one, in a migratory fashion: for each job $j\in B^{\prime}_{h}$, we scan from $r_{j}$ and process it at the first available idle slot, without delaying the original schedule.
-If $j$’s processing overlaps a busy period, we interrupt it and resume at the next idle slot, possibly on a different machine.
-Let $F^{\prime}_{j}$ be the flow time of such a job $j$.
-It consists of $p_{j}$ plus some waiting time.
-This waiting time falls entirely in busy periods and is thus upper bounded by $\sum_{i\in S_{h}\cup B^{\prime}_{h}}p_{i}/m$.
-Since there are at most $\sqrt{nm}$ jobs in $B^{\prime}_{h}$ (by our restart rule), we get:
-$F(\textup{{OPT}}^{m}(S_{h}\cup B^{\prime}_{h}))\leq F(\textup{{OPT}}(S_{h}))+O(\sqrt{n/m})\sum_{j\in S_{h}\cup B^{\prime}_{h}}p_{j}.$
-Here, we use $\textup{{OPT}}^{m}$ to denote the optimal preemptive solution with migration.
-Note that migration is required in the above construction, and NSJF is compared against $\textup{{OPT}}^{m}$ in Lemma 2.
-Summing over all $h$, we obtain:
-$\sum_{j\in S}F_{j}(\textup{{NSJF}}(S\cup U^{\prime}\cup B^{\prime}))$ $\leq\sum_{h}F(\textup{{NSJF}}(S_{h}\cup B^{\prime}_{h}))$
-$\leq\sum_{h}F(\textup{{OPT}}^{m}(S_{h}\cup B^{\prime}_{h}))+O\left(\frac{|S_{h}|+|B^{\prime}_{h}|}{\sqrt{nm}}\right)\cdot\textup{{OPT}}$
-$\leq\sum_{h}F(\textup{{OPT}}(S_{h}))+O\left(\frac{|S_{h}|+|B^{\prime}_{h}|}{\sqrt{nm}}\right)\cdot\textup{{OPT}}+O(\sqrt{n/m}\sum_{j\in S_{h}\cup B^{\prime}_{h}})p_{j}$
-$\leq F(\textup{{OPT}}(S))+O(\sqrt{n/m})\cdot\textup{{OPT}}+O(\sqrt{n/m})\sum_{j\in S\cup B^{\prime}}p_{j}$
-$\leq O(\sqrt{n/m})\cdot\textup{{OPT}}.$
-The last inequality holds because all jobs in $S\cup U^{\prime}$ can be charged to a corresponding larger job in the original job set $J$.
-Notably, a proxy job appears only when a larger job arrives, and thus, we can charge it to that larger job.
-Finally, since our algorithm schedules $S\cup U\cup B$ exactly the same as $\textup{{NSJF}}(S\cup U^{\prime}\cup B^{\prime})$, the lemma follows.
-∎
+In the absence of blocked small jobs and committed large jobs, the total flow time contributed by the jobs in $S$ is simply $F(\textup{{NSJF}}(S))$, which is at most $O\!\left(\sqrt{n/m}\right)\cdot F(\textup{{OPT}}(S))$ by Lemma 2.
+Indeed, active large jobs never interfere with the scheduling of unblocked small jobs: if a job $j$ is unblocked, then no active large job is running at its release time, and no active large job can be started before $j$, since every such job is larger than $j$.
+We now account for the effect of blocked small jobs and committed large jobs.
+To analyze their effect, we partition the time line into intervals
+$I_{h}=[t_{h}^{(\ell)},t_{h}^{(r)}),\qquad h=1,2,\dots,H,$
+defined recursively as follows.
+We let $t_{1}^{(\ell)}$ be the first time at which an unblocked small job is released.
+Given $t_{h}^{(\ell)}$, we let $t_{h}^{(r)}$ be the first time at or after $t_{h}^{(\ell)}$ at which the algorithm starts processing an active large job, which may later be killed or retired.
+If no such time exists, we set $t_{h}^{(r)}=\infty$ and the construction completes.
+Otherwise, we let $t_{h+1}^{(\ell)}$ be the first time after $t_{h}^{(r)}$ at which an unblocked small job is released, and continue recursively.
+By construction, every unblocked small job is released in some interval $I_{h}$.
+The key properties of each interval $I_{h}$ are the following.
+First, no active large job is processed inside $I_{h}$: by definition, no large job is processing at $t_{h}^{(\ell)}$ since an unblocked small job is released at $t_{h}^{(\ell)}$, and no large job starts processing inside $I_{h}$.
+Second, every blocked small job released before $t_{h}^{(\ell)}$ must start processing before $t_{h}^{(r)}$: the algorithm prioritizes small jobs over the large job starting at $t_{h}^{(r)}$.
+For each $h$, let $S_{h}$ be the set of unblocked small jobs released in $I_{h}$, and let $K_{h}$ be the set of blocked small jobs that started processing in $I_{h}$.
+For every job in $K_{h}$, we modify its release time to its start time and obtain the corresponding set $K_{h}^{\prime}$.
+Shifting the time origin to $t_{h}^{(\ell)}$, the schedule of $S_{h}\cup K_{h}^{\prime}$ inside $I_{h}$ can be viewed as $\textup{{NSJF}}(S_{h}\cup K_{h}^{\prime},\vec{b_{h}})$, where $\vec{b_{h}}$ is the initial blocking vector: machine $i$ is unavailable during $0,b_{h,i})$ due to a small job or a committed large job still being processed at time $t_{h}^{(\ell)}$, with $b_{h,i}=0$ if machine $i$ is idle.
+Therefore,
+$\sum_{j\in S_{h}}F_{j}\leq\sum_{j\in S_{h}\cup K^{\prime}_{h}}F_{j}=F(\textup{{NSJF}}(S_{h}\cup K_{h}^{\prime},\vec{b_{h}})).$
+We now apply the generalized NSJF bound from Lemma 29.
+Since every small job or committed large job has size at most $O(\textup{{OPT}}/\sqrt{nm})$, the total blocking time $B_{h}=\sum_{i}b_{h,i}$ satisfies $B_{h}\leq m\cdot O(\textup{{OPT}}/\sqrt{nm})=O(\sqrt{m/n}\cdot\textup{{OPT}})$.
+Hence,
+$F(\textup{{NSJF}}(S_{h}\cup K_{h}^{\prime},\vec{b_{h}}))\leq F(\textup{{OPT}}^{m}(S_{h}\cup K_{h}^{\prime}))+O\!\left(\frac{|S_{h}|+|K_{h}^{\prime}|}{\sqrt{nm}}\right)\textup{{OPT}},$
+where $\textup{{OPT}}^{m}$ denotes the optimal preemptive schedule with migration.
+It remains to compare $F(\textup{{OPT}}^{m}(S_{h}\cup K_{h}^{\prime}))$ with $F(\textup{{OPT}}(S_{h}))$.
+For this, consider an optimal preemptive schedule $\textup{{OPT}}(S_{h})$.
+We insert the jobs in $K_{h}^{\prime}$ into $\textup{{OPT}}(S_{h})$ one by one, in a migratory fashion: for each job $j\in K_{h}^{\prime}$, we scan forward from $r_{j}$ and process $j$ at the first available idle slot without delaying the original schedule.
+If $j$ overlaps a busy period, we interrupt it and resume it at the next idle slot, possibly on a different machine.
+Let $F_{j}^{\prime}$ denote the resulting flow time of job $j\in K_{h}^{\prime}$ in this schedule.
+Since $j$ is only delayed during busy periods, its waiting time $F_{j}^{\prime}-p_{j}$ is at most the total busy time, which is $\sum_{i\in S_{h}\cup K_{h}^{\prime}}p_{i}/m$.
+It remains to bound $|K_{h}^{\prime}|$.
+By the restart rule, if a job belongs to $K_{h}$, then it must have been released while some machine was processing a currently unproxied large job, and hence before $t_{h}^{(\ell)}$.
+Therefore, every job in $K_{h}$ is already waiting at time $t_{h}^{(\ell)}$.
+By Lemma 12, the number of waiting blocked small jobs is at most $\sqrt{nm}$, so
+$|K_{h}|=|K_{h}^{\prime}|\leq\sqrt{nm}.$
+It follows that
+$F(\textup{{OPT}}^{m}(S_{h}\cup K_{h}^{\prime}))\leq F(\textup{{OPT}}(S_{h}))+O\!\left(\sqrt{\frac{n}{m}}\right)\sum_{j\in S_{h}\cup K_{h}^{\prime}}p_{j}.$
+Combining the above bounds and summing over all $h$, we get
+$\sum_{j\in S}F_{j}$ $\leq\sum_{h=1}^{H}F(\textup{{NSJF}}(S_{h}\cup K_{h}^{\prime},\vec{b_{h}}))$
+$\leq\sum_{h=1}^{H}F(\textup{{OPT}}^{m}(S_{h}\cup K_{h}^{\prime}))+O\!\left(\sum_{h=1}^{H}\frac{|S_{h}|+|K_{h}^{\prime}|}{\sqrt{nm}}\right)\textup{{OPT}}$
+$\leq\sum_{h=1}^{H}F(\textup{{OPT}}(S_{h}))+O\!\left(\sum_{h=1}^{H}\frac{|S_{h}|+|K_{h}^{\prime}|}{\sqrt{nm}}\right)\textup{{OPT}}$
+$\quad+O\!\left(\sqrt{\frac{n}{m}}\right)\sum_{h=1}^{H}\sum_{j\in S_{h}\cup K_{h}^{\prime}}p_{j}.$
+Finally, the sets $\{S_{h}\}_{h=1}^{H}$ partition $S$, and the sets $\{K_{h}^{\prime}\}_{h=1}^{H}$ partition $K^{\prime}$. Thus,
+$\sum_{h=1}^{H}F(\textup{{OPT}}(S_{h}))\leq F(\textup{{OPT}}(S)),\qquad\sum_{h=1}^{H}(|S_{h}|+|K_{h}^{\prime}|)\leq n.$
+The total processing volume of the jobs in $\bigcup_{h}(S_{h}\cup K_{h}^{\prime})$ is at most that of the original jobs in $J$, giving
+$\sum_{h=1}^{H}\sum_{j\in S_{h}\cup K_{h}^{\prime}}p_{j}\leq O(\textup{{OPT}}).$
+Therefore,
+$\sum_{j\in S}F_{j}\leq O\!\left(\sqrt{\frac{n}{m}}\right)\textup{{OPT}}.$
 Lemma 14 (Blocked Small Job).
-Let $B$ be the set of blocked small jobs. Then the total flow time of jobs in $B$ is bounded by:
-$\sum_{j\in B}F_{j}\leq O\left(\sqrt{\frac{n}{m}}\right)\cdot\textup{{OPT}}.$
+Let $K$ be the set of blocked small jobs. Then the total flow time of jobs in $K$ is bounded by:
+$\sum_{j\in K}F_{j}\leq O\left(\sqrt{\frac{n}{m}}\right)\cdot\textup{{OPT}}.$
 Proof.
@@
 By our scheduling rule, $j$ can only experience waiting time when all Small-Only machines are busy, since $j$ is classified as a small job.
-Therefore, during the entire interval $r_{j},s_{j})$, all Small-Only machines must be fully occupied.
+During the entire interval $r_{j},s_{j})$, all Small-Only machines must be fully occupied.
 We refer to this as a busy time of Small-Only machines.
-On the other hand, due to the restart rule, at any busy time $t$ of Small-Only machines, the number of blocked small jobs still waiting is at most $\sqrt{nm}$.
-This is because the number of blocked small jobs increases only when a new large job starts processing.
-However, when that happens, all previously blocked small jobs are scheduled and no longer waiting.
+Due to the restart rule, at any busy time $t$ of Small-Only machines, the number of blocked small jobs still waiting is at most $\sqrt{nm}$ by Lemma 12.
 Since there are $\lfloor m/2\rfloor$ Small-Only machines and they are fully occupied during busy periods, the total length of all busy intervals is at most
@@
 $O\left(\sqrt{\frac{n}{m}}\right)\cdot\textup{{OPT}}.$
-We remark that when $m=1$, there is only one Small-Only machine, and this argument no longer holds. Finally, the total flow time of all blocked small jobs is:
-$\sum_{j\in B}F_{j}=\sum_{j\in B}(s_{j}-r_{j}+p_{j})\leq O\left(\sqrt{\frac{n}{m}}\right)\cdot\textup{{OPT}}.$
-For each large job $j$, consider its completion time $C_{j}$, which is either its own completion time or the completion time of its proxy job $j^{\prime}$.
-If it is proxied, the delay from $r_{j^{\prime}}$ to $C_{j^{\prime}}=C_{j}$ is already charged to the small job flow time in Lemma 12 and Lemma 13.
-Thus, we only need to analyze the modified flow time where
+We remark that when $m=1$, there is no Small-Only machine, and this argument no longer holds. Finally, the total flow time of all blocked small jobs is:
+$\sum_{j\in K}F_{j}=\sum_{j\in K}(s_{j}-r_{j}+p_{j})\leq O\left(\sqrt{\frac{n}{m}}\right)\cdot\textup{{OPT}}.$
+For each large job $j$, we only need to analyze the modified flow time:
 […]
 Lemma 15 (Large Job).
-For large job set $L$. We have $\sum_{j\in L}\hat{F}_{j}\leq O(\sqrt{n/m})\cdot\textup{{OPT}}$.
+For the large-job set $L$, we have $\sum_{j\in L}\hat{F}_{j}\leq O(\sqrt{n/m})\cdot\textup{{OPT}}$.
 Proof.
-Ignoring the final $p_{j}$ units of processing time, we focus on the delay interval $r_{j},s_{j})$ (or $r_{j},r_{j^{\prime}})$ if $j$ is proxied by $j^{\prime}$).
-Define $t_{j}=s_{j}$ (or $t_{j}=r_{j^{\prime}}$ if proxied), and aim to bound $t_{j}-r_{j}$.
-We begin by showing that shifting the release time $r_{j}$ by $\sqrt{n/m}\cdot p_{j}$ is acceptable. Let $t^{\prime}_{j}=r_{j}+\sqrt{n/m}\cdot p_{j}$. Then:
+We focus on the delay interval $r_{j},t_{j})$, where
+$t_{j}=\begin{cases}r_{j^{\prime}}&\text{if $j$ is proxied by $j^{\prime}$,}\\ s_{j}&\text{otherwise,}\end{cases}$
+and aim to bound $t_{j}-r_{j}$. We begin by showing that shifting the release time $r_{j}$ by $\sqrt{n/m}\cdot p_{j}$ is acceptable. Let $t^{\prime}_{j}=r_{j}+\sqrt{n/m}\cdot p_{j}$. Then:
 $\sum_{j\in L}(t^{\prime}_{j}-r_{j})=\sqrt{\frac{n}{m}}\cdot\sum_{j\in L}p_{j}\leq\sqrt{\frac{n}{m}}\cdot\textup{{OPT}}.$ (4)
 We now analyze the remaining delay interval $t^{\prime}_{j},t_{j})$. For each time $t\in[t^{\prime}_{j},t_{j})$, we classify it into one of the following two categories:
-  1. 1.
-    Genuine busy time: All Mixed machines are busy at $t$ originally, and at least $\lceil m/4\rceil$ of the running jobs are not restarted later.
-  2. 2.
-    Fake busy time: All Mixed machines are busy at $t$ originally, but more than $\lceil m/4\rceil$ of the running jobs are restarted.
-We decompose the delay $t_{j}-t^{\prime}_{j}$ into two parts: $F_{j}^{(1)}$ and $F_{j}^{(2)}$, corresponding to time spent in genuine busy and fake busy intervals, respectively:
-$t_{j}-t^{\prime}_{j}=F_{j}^{(1)}+F_{j}^{(2)}.$
-Bounding $F_{j}^{(1)}$: At any genuine busy time, at least $\lceil m/4\rceil$ machines are processing jobs that are not restarted.
+  1. $(\checkmark)$
+    Genuine busy time: All Mixed machines are busy at $t$, and at least $\lceil m/4\rceil$ of the jobs running on Mixed machines are not restarted later.
+  2. $(\times)$
+    Fake busy time: All Mixed machines are busy at $t$, but more than $\lceil m/4\rceil$ of the jobs running on Mixed machines will be restarted later.
+We decompose the delay $t_{j}-t^{\prime}_{j}$ into two parts: $F_{j}^{(\checkmark)}$ and $F_{j}^{(\times)}$, corresponding to time spent in genuine busy and fake busy intervals, respectively:
+$t_{j}-t^{\prime}_{j}=F_{j}^{(\checkmark)}+F_{j}^{(\times)}.$
+Bounding $F_{j}^{(\checkmark)}$ (genuine busy).
+At any genuine busy time, at least $\lceil m/4\rceil$ machines are processing jobs that are not restarted.
 The total duration of such intervals is at most $O(\textup{{OPT}}/m)$.
-Furthermore, each such interval can contribute to the delay of at most $\sqrt{nm}$ different large jobs (by Lemma 11).
-Hence:
-$\sum_{j\in L}F_{j}^{(1)}\leq\sqrt{nm}\cdot O\left(\frac{\textup{{OPT}}}{m}\right)=O\left(\sqrt{\frac{n}{m}}\cdot\textup{{OPT}}\right).$ (5)
-Bounding $F_{j}^{(2)}$: Next, we analyze the contribution of fake busy intervals, which are induced by kill events.
-Let the $i$-th kill event occur at time $\beta_{i}$.
-For a time $t$ to contribute to $F_{j}^{(2)}$, the following conditions must hold:
-  1. 1.
-    $t>\beta_{i-1}$ and $t>t^{\prime}_{j}$,
-  2. 2.
-    $t$ is not a genuine busy time.
-We now prove that such $t$ satisfies:
-$\beta_{i}-t\leq p_{j}.$ (6)
-Proof of the claim: Since $t$ is a fake busy time, more than $\lceil m/4\rceil$ jobs are restarted at $\beta_{i}$.
+Since there are at most $\ell=\lfloor\sqrt{nm}\rfloor$ large jobs at any time by Lemma 11, we have:
+$\sum_{j\in L}F_{j}^{(\checkmark)}\leq\sqrt{nm}\cdot O\left(\frac{\textup{{OPT}}}{m}\right)=O\left(\sqrt{\frac{n}{m}}\cdot\textup{{OPT}}\right).$ (5)
+Bounding $F_{j}^{(\times)}$ (fake busy).
+Next, we analyze the contribution of fake busy intervals induced by the kill events. Let the $i$-th kill event occur at time $\beta_{i}$. The key observation is the following.
+Claim 1.
+For any $t\in[t^{\prime}_{j},t_{j})$ that is a fake busy time between consecutive kill events $\beta_{i-1}$ and $\beta_{i}$, we have $\beta_{i}-t\leq p_{j}$.
+Proof of 1.
+Since $t$ is a fake busy time, more than $\lceil m/4\rceil$ jobs are restarted at $\beta_{i}$.
 Suppose, for contradiction, that $\beta_{i}-t>p_{j}$.
 Then these jobs must have been running continuously since $t$, implying that their processing times are at least $p_{j}$.
-By our algorithm, it is impossible that job $j$ is already released when we decide to start one of these larger jobs.
+By our algorithm, job $j$ had not yet been released when the algorithm started one of these larger jobs.
 Therefore, letting $U$ denote this set of jobs, we have:
 $\beta_{i}<r_{j}+\min_{u\in U}p_{u}.$
-However, by definition $t^{\prime}_{j}=r_{j}+\sqrt{n/m}\cdot p_{j}$, and assuming the total workload released up to $r_{j}$ is $P(r_{j})$, we have:
-$t^{\prime}_{j}\geq r_{j}+\frac{4P(r_{j})}{m}\geq r_{j}+\frac{4\sum_{u\in U}p_{u}}{m}\geq r_{j}+\min_{u\in U}p_{u}>\beta_{i}.$
-This proves (6), by $\beta_{i}-t\leq\beta_{i}-t^{\prime}_{j}<0$.
-Thus, for each kill event, the total contribution to $\sum_{j}F_{j}^{(2)}$ is at most:
+However, by the partition rule $p_{j}\geq 4P(r_{j})/\ell$ and $\ell=\lfloor\sqrt{nm}\rfloor$, we have $t^{\prime}_{j}=r_{j}+\sqrt{n/m}\cdot p_{j}\geq r_{j}+4P(r_{j})/m$.
+Since the jobs in $U$ were released before $r_{j}$, their total size is at most $P(r_{j})$.
+Moreover, $|U|>\lceil m/4\rceil$, so:
+$\frac{4\sum_{u\in U}p_{u}}{m}\geq\frac{4\cdot|U|\cdot\min_{u\in U}p_{u}}{m}>\min_{u\in U}p_{u}.$
+Combining,
+$t^{\prime}_{j}\geq r_{j}+4P(r_{j})/m\geq r_{j}+4\sum_{u\in U}p_{u}/m>r_{j}+\min_{u\in U}p_{u}>\beta_{i}.$
+This gives $\beta_{i}-t\leq\beta_{i}-t^{\prime}_{j}<0$, a contradiction. ∎
+By 1, each large job $j$ contributes at most $p_{j}$ to $F_{j}^{(\times)}$ per kill event. Thus, for each kill event, the total contribution to $\sum_{j}F_{j}^{(\times)}$ is at most:
 $\sum_{j\in L}p_{j}\leq\textup{{OPT}}.$
 Finally, we bound the number of kill events.
-Since each kill is triggered only after at least $\lfloor\sqrt{nm}\rfloor$ new small jobs are released, becoming blocked, the total number of restart events is at most $O(\sqrt{n/m})$.
+Since each kill is triggered only after at least $\lfloor\sqrt{nm}\rfloor$ new blocked small jobs are released, the total number of kill events is at most $O(\sqrt{n/m})$.
 Therefore:
-$\sum_{j\in L}F_{j}^{(2)}\leq\sqrt{\frac{n}{m}}\cdot\textup{{OPT}}.$ (7)
+$\sum_{j\in L}F_{j}^{(\times)}\leq\sqrt{\frac{n}{m}}\cdot\textup{{OPT}}.$ (6)
 Combining (4), (5), and (6), we obtain:
-$\sum_{j\in L}\hat{F_{j}}\leq\sum_{j\in L}\left(p_{j}+t_{j}-t^{\prime}_{j}+t^{\prime}_{j}-r_{j}\right)\leq O\left(\sqrt{\frac{n}{m}}\right)\cdot\textup{{OPT}}.$
+$\sum_{j\in L}\hat{F}_{j}\leq\sum_{j\in L}\left(p_{j}+t_{j}-t^{\prime}_{j}+t^{\prime}_{j}-r_{j}\right)=O\left(\sqrt{\frac{n}{m}}\right)\cdot\textup{{OPT}}.$
 Theorem 3.
-There exists an online polynomial-time deterministic algorithm with kill-and-restart for total flow time minimization, e.g., Algorithm 4, that is $O(\sqrt{n/m})$-competitive against the preemptive offline solution, provided that $m\geq 2$.
+Algorithm 4 is an online polynomial-time deterministic algorithm with kill-and-restart for total flow time minimization that is $O(\sqrt{n/m})$-competitive against the preemptive offline solution, provided that $m\geq 2$.
 Proof.
 By Lemma 13, Lemma 14, and Lemma 15.
-$\sum_{j\in J}F_{j}\leq\sum_{j\in S}F_{j}+\sum_{j\in B}F_{j}+\sum_{j\in L}\hat{F}_{j}\leq O\left(\sqrt{\frac{n}{m}}\right)\cdot\textup{{OPT}}.$
+$\sum_{j\in J}F_{j}\leq\sum_{j\in S}F_{j}+\sum_{j\in K}F_{j}+\sum_{j\in L}\hat{F}_{j}=O\left(\sqrt{\frac{n}{m}}\right)\cdot\textup{{OPT}}.$
 7 Deterministic Algorithm with Kill-and-Restart without the Knowledge of $n$
-In this section, we extend the $\sqrt{n/m}$-competitive algorithm (which assumes the prior knowledge of $n$) to an $O(n^{\alpha}/\sqrt{m})$-competitive algorithm without such knowledge, where $\alpha=(\sqrt{5}-1)/2$.
-The overall framework still follows Section 6: we keep the same machine partition (Small-Only machines $1$ to $\lfloor m/2\rfloor$ and Mixed machines $\lfloor m/2\rfloor+1$ to $m$), and we still kill a large job once enough small jobs are blocked.
-The key difference is that the total number of jobs $n$ is unknown in advance.
-At any time $t$, we only observe the number of jobs released so far, denoted by $n(t)$, rather than the final value of $n$.
-Consequently, the rank-based partition can no longer use a fixed threshold; instead, it must classify jobs as large or small using a dynamic threshold $\ell(t)$ that depends on $n(t)$.
-This introduces a major difficulty: the partition is no longer stable as in the known-$n$ setting.
-Previously, when $n$ was known, the threshold $\ell$ was fixed.
-A small job would remain small forever, and the only source of instability was when a large job retired, which we handled by introducing a proxy job.
-In contrast, in the unknown-$n$ setting, $\ell(t)$ increases monotonically as $n(t)$ grows, and thus varies over time.
-As a result, a job may switch between being large and small multiple times.
-For instance, when $n(t)$ is small, we may treat the top $3$ jobs as large, so a job of rank $4$ is small.
-Later, as more jobs arrive and $\ell(t)$ becomes $4$, this job becomes large.
-If even more jobs arrive, $\ell(t)$ may increase to $5$, but that same job may drop out of the top-$5$ and become small again.
-This uncontrolled oscillation between the large and small states makes the partition unstable and invalidates the previous analysis.
-To this end, we make two main modifications as follow in the new algorithm as follows, and the formal description is presented in Algorithm 5.
-More Dynamic Partition Rule.
-We use a partition rule that depends on the number of jobs released so far.
-Instead of using $n(t)$ directly, we let $N=2^{k}$ be the smallest power of two such that $N\geq n(t)$, and use $N$ to define the large–small threshold $\ell$ at every time $t$, where $\ell=\lfloor N^{\alpha}\sqrt{m}\rfloor$.
-Let $T_{j}(t)$ denote the type of job $j$ at time $t$.
-We classify $T_{j}(t)$ as large if job $j$ is among the top $\ell$ jobs in size and $p_{j}\geq\frac{4P(t)}{\ell}$, where $P(t)$ is the total size of all jobs released by time $t$.
-Otherwise, $T_{j}(t)$ is classified as small.
-A New Kill Event When $N$ Is Updated.
-Previously, we triggered a kill event only when too many small jobs were waiting while a large job was being processed on a Mixed machine.
-However, in this setting, a large job may be processed on a Small-Only machine, since it might have been classified as small earlier.
-To prevent this from causing inefficiency, we introduce a new kill event that occurs whenever $N$ is updated (that is, when $n(t)>N$, and we double $N$ to $2N$).
-At this moment, we check whether any currently large jobs are running on the Small-Only machines and kill all of them if so.
-We remark that this is one reason for using a rounded value of $n(t)$ (the nearest power of two) as the benchmark, since it allows us to bound the number of such kill events.
-Data: Estimate of $n$: NN initiated as 11;
+In this section, we extend the $O(\sqrt{n/m})$-competitive kill-and-restart algorithm of Section 6 to the setting where the total number of jobs $n$ is unknown in advance.
+The resulting algorithm achieves a competitive ratio of $O(n^{\alpha}/\sqrt{m})$, where $\alpha=(\sqrt{5}-1)/2$.
+The overall framework is the same: we keep the machine partition into Small-Only and Mixed machines, and we still kill large jobs once enough small jobs are blocked.
+The key challenge is in the partition.
+Recall that in the known-$n$ setting, the threshold $\ell$ was fixed, making the partition nearly stable: a small job could never become large (since $\ell$ does not grow), and the only reclassification was in the opposite direction — a large job could be retired and handled via a proxy job.
+Without knowledge of $n$, the threshold $\ell(t)$ must depend on the number of jobs $n(t)$ released so far, and it increases over time.
+This means a job’s classification can oscillate: for example, a job of rank $4$ is small when $\ell=3$, becomes large when $\ell$ grows to $4$, and may become small again when $\ell$ reaches $5$ and larger jobs take its place.
+This instability invalidates the previous analysis.
+To address this, we make two modifications, described below and formalized in Algorithm 5.
+Dynamic partition rule.
+We round $n(t)$ up to the nearest power of two, $N=2^{\lceil\log_{2}n(t)\rceil}$, and set the threshold $\ell=\lfloor N^{\alpha}\sqrt{m}\rfloor$.
+At every time $t$, a job $j$ is classified as large if it is among the top $\ell$ jobs by processing time and $p_{j}\geq 4P(t)/\ell$, where $P(t)$ is the total processing time of all jobs released by time $t$; otherwise, it is classified as small.
+The rounding ensures that $\ell$ changes only $O(\log n)$ times.
+Additional kill event on threshold update.
+When $N$ doubles, some jobs currently running on Small-Only machines may now be classified as large (since $\ell$ has increased and new jobs may have displaced them in rank).
+We kill all such jobs immediately.
+This is a new type of kill event, in addition to the existing one triggered when too many small jobs are blocked.
+Data: Estimate of $n$: NN initialized to 11;
 On job $j$ release at $t$:
@@
    end if
-On any job completion or after any job release at $t$:
+On any job completion or after any job release:
  for each idle Small-Only machine $i$ do
@@
 $\sum_{j\in J}(t^{\prime}_{j}-r_{j})=\frac{(2n)^{\alpha}}{\sqrt{m}}\sum_{j\in J}p_{j}\leq\frac{(2n)^{\alpha}}{\sqrt{m}}\cdot\textup{{OPT}}.$ (7)
-If the total waiting time after $t^{\prime}$ is already bounded by $O(n^{\alpha}/\sqrt{m})\textup{{OPT}}$, then $\sum_{j\in J}F^{(1)}_{j}$ is also bounded by $O(n^{\alpha}/\sqrt{m})\textup{{OPT}}$.
+If the total waiting time after the shifted times $t^{\prime}_{j}$ is already bounded by $O(n^{\alpha}/\sqrt{m})\textup{{OPT}}$, then $\sum_{j\in J}F^{(1)}_{j}$ is also bounded by $O(n^{\alpha}/\sqrt{m})\textup{{OPT}}$.
 For each job $j$, let its large waiting interval be $t^{\prime}_{j},t_{j})$. For every time $t\in[t^{\prime}_{j},t_{j})$, we classify it into two types:
-  1. 1.
-    Genuine busy time: All Mixed machines are busy at $t$, and at least $\lceil m/4\rceil$ of the running jobs are not restarted later.
-  2. 2.
-    Fake busy time: All Mixed machines are busy at $t$, but more than $\lceil m/4\rceil$ of the running jobs will be restarted later.
-We decompose the delay $t_{j}-t^{\prime}_{j}$ into two parts, $F_{j}^{(A)}$ and $F_{j}^{(B)}$, corresponding to genuine and fake busy intervals:
-$t_{j}-t^{\prime}_{j}=F_{j}^{(A)}+F_{j}^{(B)}.$
-Bounding $F_{j}^{(A)}$.
+  1. $(\checkmark)$
+    Genuine busy time: All Mixed machines are busy at $t$, and at least $\lceil m/4\rceil$ of the jobs running on Mixed machines are not restarted later.
+  2. $(\times)$
+    Fake busy time: All Mixed machines are busy at $t$, but more than $\lceil m/4\rceil$ of the jobs running on Mixed machines will be restarted later.
+We decompose the delay $t_{j}-t^{\prime}_{j}$ into two parts: $F_{j}^{(\checkmark)}$ and $F_{j}^{(\times)}$, corresponding to time spent in genuine busy and fake busy intervals, respectively:
+$t_{j}-t^{\prime}_{j}=F_{j}^{(\checkmark)}+F_{j}^{(\times)}.$
+Bounding $F_{j}^{(\checkmark)}$ (genuine busy).
 During genuine busy intervals, at least $\lceil m/4\rceil$ machines are processing jobs that are never restarted.
@@
 Hence,
-$\sum_{j\in J}F_{j}^{(A)}\leq O\!\left(\frac{n^{\alpha}}{\sqrt{m}}\cdot\textup{{OPT}}\right).$ (9)
-Bounding $F_{j}^{(B)}$.
-Now consider fake busy intervals caused by type-B kill events. Let the $i$-th type-B kill event occur at time $\beta_{i}$. For a time $t$ to contribute to $F_{j}^{(B)}$, it must satisfy:
-  1. 1.
-    $t>\beta_{i-1}$ and $t>t^{\prime}_{j}$,
-  2. 2.
-    $t$ is not a genuine busy time.
-We claim that
-$\beta_{i}-t\leq p_{j}.$ (10)
-Proof of the claim.
+$\sum_{j\in J}F_{j}^{(\checkmark)}\leq O\!\left(\frac{n^{\alpha}}{\sqrt{m}}\cdot\textup{{OPT}}\right).$ (8)
+Bounding $F_{j}^{(\times)}$ (fake busy).
+Now consider fake busy intervals.
+These must be caused by type-B kill events, since type-A kill events only kill jobs on Small-Only machines.
+Let the $i$-th type-B kill event occur at time $\beta_{i}$.
+The key observation is the following.
+Claim 2.
+For any $t\in[t’_{j},t_{j})$ that is a fake busy time between consecutive type-B kill events $\beta_{i-1}$ and $\beta_{i}$, we have $\beta_{i}-t\leq p_{j}$.
+Proof of 2.
 Since $t$ is a fake busy time, more than $\lceil m/4\rceil$ jobs are restarted at $\beta_{i}$.
-Assume for contradiction that $\beta_{i}-t>p_{j}$.
-Then these jobs must have been running continuously since $t$, so their processing times are at least $p_{j}$.
-By the algorithm’s rule, job $j$ cannot have been released before starting any of these larger jobs.
-Let $U$ denote this set of running jobs.
-Then
+Suppose, for contradiction, that $\beta_{i}-t>p_{j}$.
+Then these jobs must have been running continuously since $t$, implying that their processing times are at least $p_{j}$.
+By our algorithm, job $j$ had not yet been released when the algorithm started one of these larger jobs.
+Therefore, letting $U$ denote this set of jobs, we have
 $\beta_{i}<r_{j}+\min_{u\in U}p_{u}.$
-However, by definition,
-$t^{\prime}_{j}=r_{j}+\frac{(2n)^{\alpha}}{\sqrt{m}}\cdot p_{j},$
-and let $P(r_{j})$ denote the total workload released up to $r_{j}$, then
-$t^{\prime}_{j}\geq r_{j}+\frac{4P(r_{j})}{m}\geq r_{j}+\frac{4\sum_{u\in U}p_{u}}{m}\geq r_{j}+\min_{u\in U}p_{u}>\beta_{i},$
-contradicting $\beta_{i}<t^{\prime}_{j}$. Thus, (10) holds.
-Each kill event therefore contributes at most $\sum_{j\in J}p_{j}\leq\textup{{OPT}}$ to $\sum_{j}F_{j}^{(B)}$.
+However, by the partition rule $p_{j}\geq 4P(r_{j})/\ell$ and $\ell=\lfloor N^{\alpha}\sqrt{m}\rfloor$, we have $t’_{j}=r_{j}+(2n)^{\alpha}/\sqrt{m}\cdot p_{j}\geq r_{j}+4P(r_{j})/m$.
+Since the jobs in $U$ were released before $r_{j}$, their total size is at most $P(r_{j})$.
+Moreover, $|U|>\lceil m/4\rceil$, so:
+$\frac{4\sum_{u\in U}p_{u}}{m}\geq\frac{4\cdot|U|\cdot\min_{u\in U}p_{u}}{m}>\min_{u\in U}p_{u}.$
+Combining, $t’_{j}\geq r_{j}+4P(r_{j})/m\geq r_{j}+4\sum_{u\in U}p_{u}/m>r_{j}+\min_{u\in U}p_{u}>\beta_{i}$. This gives $\beta_{i}-t\leq\beta_{i}-t’_{j}<0$, a contradiction. ∎
+Each kill event therefore contributes at most $\sum_{j\in J}p_{j}\leq\textup{{OPT}}$ to $\sum_{j}F_{j}^{(\times)}$.
 Since the total number of type-B kill events is at most $O(n^{1-\alpha}/\sqrt{m})$, we obtain
-$\sum_{j\in J}F_{j}^{(B)}\leq O\!\left(\frac{n^{1-\alpha}}{\sqrt{m}}\right)\cdot\textup{{OPT}}\leq O\!\left(\frac{n^{\alpha}}{\sqrt{m}}\right)\cdot\textup{{OPT}}.$ (11)
+$\sum_{j\in J}F_{j}^{(\times)}\leq O\!\left(\frac{n^{1-\alpha}}{\sqrt{m}}\right)\cdot\textup{{OPT}}\leq O\!\left(\frac{n^{\alpha}}{\sqrt{m}}\right)\cdot\textup{{OPT}}.$ (9)
 The last inequality holds because $\alpha$ is set to $(\sqrt{5}-1)/2>0.5$.
 Finally, combining (7), (8), and (9), we have
-$\sum_{j\in J}F^{(1)}_{j}\leq\sum_{j\in J}\bigl((t^{\prime}_{j}-r_{j})+F_{j}^{(A)}+F_{j}^{(B)}\bigr)\leq O\!\left(\frac{n^{\alpha}}{\sqrt{m}}\right)\cdot\textup{{OPT}}.$
+$\sum_{j\in J}F^{(1)}_{j}\leq\sum_{j\in J}\bigl((t^{\prime}_{j}-r_{j})+F_{j}^{(\checkmark)}+F_{j}^{(\times)}\bigr)\leq O\!\left(\frac{n^{\alpha}}{\sqrt{m}}\right)\cdot\textup{{OPT}}.$
 Next, we aim to bound the total small waiting time.
@@
   1. 1.
-    Blocked: $j\in S_{h}$, and some machine is processing a job $j^{\prime}$ at $\theta_{j}$, where $\theta_{j}<\theta_{j^{\prime}}$. We use $B_{h}$ to denote the set of blocked small jobs in $S_{h}$.
+    Blocked: $j\in S_{h}$, and some machine is processing a job $j^{\prime}$ at $\theta_{j}$, where $\theta_{j}<\theta_{j^{\prime}}$. We use $K_{h}$ to denote the set of blocked small jobs in $S_{h}$.
   2. 2.
@@
 We focus on the number of waiting blocked jobs at time $t^{*}$, which is at most $n^{\alpha}\sqrt{m}$.
-Let $B^{*}_{h}$ denote this set.
-View $t^{*}$ as time 0 and form two shifted instance $U^{\prime}_{h}$ and $B^{\prime}_{h}$ by moving every release time in $U_{h}$ and $B^{*}_{h}$ by $t^{*}$, repectively.
-The schedule of $U_{h}$ in our algorithm is a shifting to the schedule of $U^{\prime}_{h}$ under $\textup{{NSJF}}(U^{\prime}_{h}\cup B^{\prime}_{h},\vec{b})$, where $\vec{b}$ is the initial blocking vector from the jobs processing at $t^{*}$.
+Let $K^{*}_{h}$ denote this set.
+View $t^{*}$ as time 0 and form two shifted instances $U^{\prime}_{h}$ and $K^{\prime}_{h}$ by moving every release time in $U_{h}$ and $K^{*}_{h}$ by $t^{*}$, respectively.
+The schedule of $U_{h}$ in our algorithm is a shifting to the schedule of $U^{\prime}_{h}$ under $\textup{{NSJF}}(U^{\prime}_{h}\cup K^{\prime}_{h},\vec{b})$, where $\vec{b}$ is the initial blocking vector from the jobs processing at $t^{*}$.
 By the definition of $t^{*}$, these processing jobs are already in their last small period, so each component of $\vec{b}$ is at most $O\!\left(\frac{\textup{{OPT}}}{n^{\alpha}\sqrt{m}}\right)$.
 By Lemma 29,
-$\sum_{j\in U_{h}}F^{(2)}_{j}\;\leq\;F(\textup{{NSJF}}(U^{\prime}_{h}\cup B^{\prime}_{h},\vec{b}))\;\leq\;F(\textup{{OPT}}^{m}(U^{\prime}_{h}\cup B^{\prime}_{h}))\;+\;O\!\left(\frac{|U^{\prime}_{h}\cup B^{\prime}_{h}|}{n^{\alpha}\sqrt{m}}\right)\textup{{OPT}},$ (12)
+$\sum_{j\in U_{h}}F^{(2)}_{j}\;\leq\;F(\textup{{NSJF}}(U^{\prime}_{h}\cup K^{\prime}_{h},\vec{b}))\;\leq\;F(\textup{{OPT}}^{m}(U^{\prime}_{h}\cup K^{\prime}_{h}))\;+\;O\!\left(\frac{|U^{\prime}_{h}\cup K^{\prime}_{h}|}{n^{\alpha}\sqrt{m}}\right)\textup{{OPT}},$ (10)
 where $\textup{{OPT}}^{m}$ denotes the optimal preemptive, migratory schedule.
-The next step is to bound the gap between $\textup{{OPT}}^{m}(U^{\prime}_{h}\cup B^{\prime}_{h})$ and $\textup{{OPT}}^{m}(U_{h})$ using the fact that $|B^{\prime}_{h}|$ is small.
+The next step is to bound the gap between $\textup{{OPT}}^{m}(U^{\prime}_{h}\cup K^{\prime}_{h})$ and $\textup{{OPT}}^{m}(U_{h})$ using the fact that $|K^{\prime}_{h}|$ is small.
 Consider an optimal preemptive, migratory schedule $\textup{{OPT}}^{m}(U^{\prime}_{h})$.
-Insert the jobs of $B^{\prime}_{h}$ one by one with migration: for each $j\in B^{\prime}_{h}$, we process it in the earliest idle slots without delaying already scheduled work; if the machine is busy, pause and resume at the next idle slot (possibly on another machine).
+Insert the jobs of $K^{\prime}_{h}$ one by one with migration: for each $j\in K^{\prime}_{h}$, we process it in the earliest idle slots without delaying already scheduled work; if the machine is busy, pause and resume at the next idle slot (possibly on another machine).
 Let $F^{\prime}_{j}$ be the flow time of such an inserted job $j$. Then
-$F^{\prime}_{j}\;=\;p_{j}+\text{wait}_{j},\qquad\text{wait}_{j}\;\leq\;\frac{1}{m}\sum_{i\in U^{\prime}_{h}\cup B^{\prime}_{h}}p_{i}.$
-By the lemma’s condition, $|B^{\prime}_{h}|=|B^{*}_{h}|\leq n^{\alpha}\sqrt{m}$, hence
-$F(\textup{{OPT}}^{m}(U^{\prime}_{h}\cup B^{\prime}_{h}))\;\leq\;F(\textup{{OPT}}^{m}(U^{\prime}_{h}))\;+\;O\!\left(\frac{n^{\alpha}}{\sqrt{m}}\right)\sum_{i\in U^{\prime}_{h}\cup B^{\prime}_{h}}p_{i}.$
+$F^{\prime}_{j}\;=\;p_{j}+\text{wait}_{j},\qquad\text{wait}_{j}\;\leq\;\frac{1}{m}\sum_{i\in U^{\prime}_{h}\cup K^{\prime}_{h}}p_{i}.$
+By the lemma’s condition, $|K^{\prime}_{h}|=|K^{*}_{h}|\leq n^{\alpha}\sqrt{m}$, hence
+$F(\textup{{OPT}}^{m}(U^{\prime}_{h}\cup K^{\prime}_{h}))\;\leq\;F(\textup{{OPT}}^{m}(U^{\prime}_{h}))\;+\;O\!\left(\frac{n^{\alpha}}{\sqrt{m}}\right)\sum_{i\in U^{\prime}_{h}\cup K^{\prime}_{h}}p_{i}.$
 Combining with Equation 10,
-$\sum_{j\in U_{h}}F^{(2)}_{j}$ $\leq\;F(\textup{{OPT}}^{m}(U^{\prime}_{h}\cup B^{\prime}_{h}))\;+\;O\!\left(\frac{|U^{\prime}_{h}\cup B^{\prime}_{h}|}{n^{\alpha}\sqrt{m}}\right)\textup{{OPT}}$
-$\leq\;F(\textup{{OPT}}^{m}(U^{\prime}_{h}))\;+\;O\!\left(\frac{n^{\alpha}}{\sqrt{m}}\right)\sum_{i\in U^{\prime}_{h}\cup B^{\prime}_{h}}p_{i}\;+\;O\!\left(\frac{|S_{h}|}{n^{\alpha}\sqrt{m}}\right)\textup{{OPT}}$
+$\sum_{j\in U_{h}}F^{(2)}_{j}$ $\leq\;F(\textup{{OPT}}^{m}(U^{\prime}_{h}\cup K^{\prime}_{h}))\;+\;O\!\left(\frac{|U^{\prime}_{h}\cup K^{\prime}_{h}|}{n^{\alpha}\sqrt{m}}\right)\textup{{OPT}}$
+$\leq\;F(\textup{{OPT}}^{m}(U^{\prime}_{h}))\;+\;O\!\left(\frac{n^{\alpha}}{\sqrt{m}}\right)\sum_{i\in U^{\prime}_{h}\cup K^{\prime}_{h}}p_{i}\;+\;O\!\left(\frac{|S_{h}|}{n^{\alpha}\sqrt{m}}\right)\textup{{OPT}}$
 $\leq\;F(\textup{{OPT}}(S_{h}))\;+\;O\!\left(\frac{n^{\alpha}}{\sqrt{m}}\right)\sum_{i\in S_{h}}p_{i}\;+\;O\!\left(\frac{|S_{h}|}{n^{\alpha}\sqrt{m}}\right)\textup{{OPT}}.$
@@
 To bound $\sum_{j\in J}F^{(2)}_{j}$, it remains to control
-$\sum_{h\in H_{2}}\sum_{j\in B_{h}}F^{(2)}_{j}$.
+$\sum_{h\in H_{2}}\sum_{j\in K_{h}}F^{(2)}_{j}$.
 Lemma 20.
 For the waiting time of blocked jobs during uncrowded periods,
-$\sum_{h\in H_{2}}\sum_{j\in B_{h}}F^{(2)}_{j}\;\leq\;O\!\left(\frac{n^{1-\alpha}}{\sqrt{m}}\right)\textup{{OPT}}.$
+$\sum_{h\in H_{2}}\sum_{j\in K_{h}}F^{(2)}_{j}\;\leq\;O\!\left(\frac{n^{1-\alpha}}{\sqrt{m}}\right)\textup{{OPT}}.$
 Proof.
@@
 Each killed job was small just before the update, hence its size is at most $\textup{{OPT}}/(N^{\alpha}\sqrt{m})$.
-Thus each such event contributes to $\sum_{h\in H_{2}}\sum_{j\in B_{h}}F^{(2)}_{j}$ at most
+Thus each such event contributes to $\sum_{h\in H_{2}}\sum_{j\in K_{h}}F^{(2)}_{j}$ at most
 $O\!\left(\frac{\textup{{OPT}}\cdot N^{1-\alpha}}{\sqrt{m}}\right).$
@@
 Theorem 4.
-There exists an online polynomial-time deterministic algorithm with kill-and-restart for total flow time minimization, e.g., Algorithm 5, that is $O(n^{\alpha}/\sqrt{m})$-competitive against the preemptive offline solution without the prior knowledge of $n$, where $\alpha=(\sqrt{5}-1)/2$, provided that $m\geq 2$.
+Algorithm 5 is an online polynomial-time deterministic algorithm with kill-and-restart for total flow time minimization that is $O(n^{\alpha}/\sqrt{m})$-competitive against the preemptive offline solution without prior knowledge of $n$, where $\alpha=(\sqrt{5}-1)/2$, provided that $m\geq 2$.
 Proof.
@@
 $\sum_{j\in J}F_{j}^{(1)}\;\leq\;O\!\left(\frac{n^{\alpha}}{\sqrt{m}}\right)\textup{{OPT}}.$
-By Lemmas 17, 18 and 19,
+By Lemmas 18, 19 and 20,
 $\sum_{j\in J}F_{j}^{(2)}\;\leq\;O\!\left(\frac{n^{\alpha}}{\sqrt{m}}\right)\textup{{OPT}}.$
@@
 Moving to the analysis of an arbitrary online algorithm, we aim to prove that the algorithm’s flow time is at least $\Omega(mk^{2})$.
-First, we will show that when a large job is placed such that it conflicts with the small jobs (having a intersection of at least $0.5$), it significantly increases the flow time of those small jobs.
+First, we will show that when a large job is placed such that it conflicts with the small jobs (having an intersection of at least $0.5$), it significantly increases the flow time of those small jobs.
 Definition 1.
@@
 Theorem 5.
-Any randomized online algorithm for minimizing the total flow time on parallel machines has a competitive ratio of at least $\Gamma(n,m)$, for every fixed pair $(n,m)$, where $\Gamma(n,m)=\Omega(\sqrt{n/m})$.
+Any randomized online algorithm for total flow time minimization on parallel machines has a competitive ratio of at least $\Omega(\sqrt{n/m})$, for every fixed pair $(n,m)$.
 Proof.
@@
 Therefore,
-$\mathbb{E}(F^{(L)})$ $\geq\sum_{i=0}^{k-1}\left((i+1)\cdot m-\sum_{j=0}^{i-1}\mathbb{E}(N_{b})\right)\geq\mathbb{E}\left(\sum_{i=0}^{k-1}\sum_{b=0}^{i}\left(m-\mathbb{E}(N_{b})\right)\right)$
+$\mathbb{E}(F^{(L)})$ $\geq\sum_{i=0}^{k-1}\left((i+1)\cdot m-\sum_{b=0}^{i-1}\mathbb{E}(N_{b})\right)\geq\mathbb{E}\left(\sum_{i=0}^{k-1}\sum_{b=0}^{i}\left(m-\mathbb{E}(N_{b})\right)\right)$
 $=\sum_{b=0}^{k-1}\left(m-\mathbb{E}(N_{b})\right)\cdot(k-b)$
@@
 $\mathbb{E}(F^{(L)})+\mathbb{E}(F^{(S)})\geq\frac{k}{16}\cdot\sum_{b=0}^{\lfloor 15k/16\rfloor}\frac{m}{2}=\Omega(k^{2}m).$
-By Lemma 21 $\textup{{OPT}}=O(mk)$, any deterministic algorithm is then $\Omega(k)=\Omega(\sqrt{n/m})$ competitive on the counterexample, and we can conclude the theorem by Yao’s minimax principle. ∎
+By Lemma 22, $\textup{{OPT}}=O(mk)$, then any deterministic algorithm is $\Omega(k)=\Omega(\sqrt{n/m})$ competitive on the counterexample, and we can conclude the theorem by Yao’s minimax principle. ∎
 8.3 Generalization to Multiple Machines with Kill-and-Restart
@@
 Theorem 6.
-Any randomized online algorithm for minimizing the total flow time on parallel machines has a competitive ratio of at least $\Gamma(n,m)$, for every fixed pair $(n,m)$, where $\Gamma(n,m)=\Omega(\sqrt{n/m})$, even with the capability of kill-and-restart.
+Any randomized online algorithm for total flow time minimization on parallel machines has a competitive ratio of at least $\Omega(\sqrt{n/m})$, for every fixed pair $(n,m)$, even with the capability of kill-and-restart.
 Proof.
@@
 $\geq\frac{k}{16}\cdot\left(\sum_{b=0}^{\lfloor\frac{7}{8}k\rfloor}m/2\right)=\Omega(k^{2}m).$
-By Lemma 23, $\textup{{OPT}}=O(km)$, hence any deterministic algorithm is $\Omega(k)=\Omega(n/m)$ competitive and the theorem is proved. ∎
+By Lemma 24, $\textup{{OPT}}=O(km)$, hence any deterministic algorithm is $\Omega(k)=\Omega(\sqrt{n/m})$ competitive and the theorem is proved. ∎
 8.4 Deterministic Algorithms: Single Machine with Kill-and-Restart
@@
     $\tau\leftarrow\min\{x\in T\mid x>t-t^{\prime}\}$ ;
-    if ALG hasn’t kill job $u$ before time $t^{\prime}+\tau$ then
+    if ALG hasn’t killed job $u$ before time $t^{\prime}+\tau$ then
          Release a job at time $t^{\prime}+\tau$ with processing time $\varepsilon$ ;
@@
 Theorem 7.
-All deterministic algorithms with kill-and-restart on single-machine total flow time minimization have a competitive ratio of at least $\Omega(n/\log n)$.
+Any deterministic online algorithm with kill-and-restart for total flow time minimization on a single machine has a competitive ratio of at least $\Omega(n/\log n)$.
 Proof.
@@
 8.5 Deterministic Algorithms: Multiple Machines without Kill-and-Restart
-The $\Omega(n/m^{2}+\sqrt{n/m})$ follows by combining our randomized lower bound of $\Omega(\sqrt{n/m})$ with the deterministic $\Omega(n/m^{2})$ lower bound from DBLP:journals/dam/EpsteinS06, as presented below.
+The lower bound $\Omega(n/m^{2}+\sqrt{n/m})$ follows by combining our randomized lower bound of $\Omega(\sqrt{n/m})$ with the deterministic $\Omega(n/m^{2})$ lower bound from [10], as presented below.
 Theorem 8.
-Any deterministic online algorithm for minimizing the total flow time on parallel machines has a competitive ratio of at least $\Gamma(n,m)$, for every fixed $n$ and $m$, where $\Gamma(n,m)=\Omega(n/m^{2}+\sqrt{n/m})$.
+Any deterministic online algorithm for total flow time minimization on parallel machines has a competitive ratio of at least $\Gamma(n,m)$, for every fixed pair $(n,m)$, where $\Gamma(n,m)=\Omega(n/m^{2}+\sqrt{n/m})$.
 Proof.
-We first show that for every $n$, $m$, we can construct an adversary to prove $\Omega(n/m^{2})$, the same as DBLP:journals/dam/EpsteinS06.
+We first show that for every $n$, $m$, we can construct an adversary to prove $\Omega(n/m^{2})$, the same as [10].
 We still present it for completeness.
@@
 Hence, OPT is $O(1)$ in both cases.
-For ALG, as it can only use m1m-1 machines to solve the small jobs, at most m1m-1 jobs can be solved between the releases of two small job batchs.
+For ALG, as it can only use m1m-1 machines to solve the small jobs, at most m1m-1 jobs can be solved between the releases of two small job batches.
 The total flow time for ALG is at least $\sum_{i=1}^{n/m}i/n=\Omega(n/m^{2})$.
@@
 At time 0, the adversary releases a single job of size $2$.
-For any algorithm, let $p(t)$ denote the probability density function that ALG starts this job at time $t$ if only the single job is present, which is known by the oblivious adversary since the probability distribution solely depend on the description of the algorithm and does not depend on algorithm’s randomized results.
+Let $p(t)$ denote the probability density that ALG starts this job at time $t$, assuming no other jobs are present.
+Since $p(t)$ depends only on the algorithm’s description, it is known to the adversary in advance.
 The algorithm must fall into one of the following categories (the parameters $x$ and $y$ will be fixed later):
@@
     There exists an interval $[t,t+1)$, where $t\in\{1,2,\dots,x\}$, such that
-    $\int_{t}^{t+1}p(t)\,dt\geq y.$
+    $\int_{t}^{t+1}p(s)\,ds\geq y.$
     In this case, the adversary releases $n_{1}$ dummy jobs of size 0 at time $t+1$.
@@
   2. B.
-    For all $t\in\{1,2,\dots,x\}$, it holds that $\int_{t}^{t+1}p(t)\,dt<y$.
+    For all $t\in\{1,2,\dots,x\}$, it holds that $\int_{t}^{t+1}p(s)\,ds<y$.
 In this case, the probability that the size-$2$ job has not been started by time $x$ is at least 1xy1-xy.
 The adversary releases $n_{0}$ dummy jobs of size 0 at time $x+1$.
-If ALG delays the size-$2$ job until after time $x$, each dummy job has flow time at least $x$, while $\textup{{OPT}}\leq 2$.
+If ALG delays the size-$2$ job until after time $x$, its flow time alone is at least $x+2$, while $\textup{{OPT}}\leq 2$.
 Therefore,
@@
 for some instance with $n\geq n_{0}$. Consequently, no randomized algorithm can achieve a competitive ratio of $o(n)$. ∎
-Corollary 26.
+Theorem 10.
 No randomized algorithm can achieve a competitive ratio of $o(n/m^{2}+\sqrt{n/m})$ without prior knowledge of $n$ in the multiple-machine setting.
+The $\sqrt{n/m}$ term follows from the known-$n$ lower bound (Theorem 5).
+For the $n/m^{2}$ term, the intuition resembles Lemma 9: losing one machine out of $m$ incurs an $\Omega(n/m^{2})$ penalty in flow time.
 Proof.
-It is known that even when $n$ is known in advance, there exists a lower bound of $\Omega(\sqrt{n/m})$.
-Therefore, it suffices to generalize Theorem 9 to obtain a hard instance with ratio $\Omega(n/m^{2})$, from which the theorem follows.
+Since the $\Omega(\sqrt{n/m})$ lower bound holds even when $n$ is known (Theorem 5), it suffices to construct a hard instance achieving ratio $\Omega(n/m^{2})$.
 Adversary construction.
-We use the same setup as in the single-machine case: the adversary first releases a size-2 job at time 0 and observes the algorithm’s random behavior.
-The only difference lies in how we construct the hard instance for Type A algorithms in the multi-machine setting.
-Hard instance for Type A.
-If ALG is of Type A, the adversary releases $n_{1}$ additional jobs of size $1/k$, where $k=n_{1}/m$. The release pattern is as follows:
+We use the same setup as in the single-machine case: the adversary first releases a size-2 job at time 0 and classifies the algorithm into Type A or Type B based on the density $p(t)$.
+The only difference lies in how we construct the hard instance for Type A.
+  1. A.
+    If ALG is of Type A, the adversary releases $n_{1}$ additional jobs of size $1/k$, where $k=n_{1}/m$. The release pattern is:
     $t+1,\;t+1+\frac{1}{k},\;t+1+\frac{2}{k},\;\dots,\;t+1+\frac{k-1}{k},$
@@
     where the last inequality uses $m\geq 1$. As $n_{1}\to\infty$, this ratio approaches $yn/(5m^{2})$.
-Hard instance for Type B.
+  2. B.
     The construction and argument are identical to the single-machine case: if the algorithm delays the size-2 job beyond time $x$, the expected ratio satisfies
@@
 for some instance with $n>n_{0}$. Consequently, no randomized algorithm can achieve a competitive ratio of $o(n/m^{2})$. ∎
-Appendix A Schedule Small Jobs: Non-Preemptive Shortest Job First
+Appendix A Generalized Version of Non-Preemptive Shortest Job First
 In this section, we analyze the performance of the Non-Preemptive Shortest Job First (NSJF) algorithm in a generalized version, with the concept of blocking period: for blocking vector $\vec{b}$, where each $b_{i}$ denotes the initial blocking time of machine $i$ — that is, machine $i$ is unavailable for processing during the interval $0,b_{i})$, we aim to prove the following bound:
@@
 $A(t)=\int_{0}^{t}a(y)\,dy.$
-Next, for any threshold $p$ and time $t$, we define $J_{\leq p}(t)$ as the set of jobs with $p_{j}\leq p$ and release time $r_{j}\leq t$.
-We are interested in how NSJF and OPT handle the jobs in $J_{\leq p}(t)$, though we measure progress differently for each:
-  • For OPT, let $V_{\leq p}(t)$ denote the total workload of jobs in $J_{\leq p}(t)$ that has been processed by time $t$.
-  • For NSJF, let $V^{\prime}_{\leq p}(t)$ denote the total processing time of jobs in $J_{\leq p}(t)$ that have started under NSJF by time $t$; that is, the total processing time of job $j$ in the set $\{j\in J_{\leq p}(t)\mid s_{j}(\textup{{NSJF}})\leq t\}$.
-Lemma 27.
-For every $t\geq 0$, define $t^{\prime}=A^{-1}(m(t+\tau))$. Then for every upper bound $p$, we have: $V^{\prime}_{\leq p}(t^{\prime})\geq V_{\leq p}(t).$
+For any processing time threshold $p$ and time $t$, let $J_{\leq p}(t)$ be the set of jobs with $p_{j}\leq p$ and release time $r_{j}\leq t$.
+Step 1: Volume Matching.
+For each algorithm, we track the volume of work associated with jobs in $J_{\leq p}(t)$.
+Let $V_{\leq p}(t)$ denote the total volume of work processed on jobs in $J_{\leq p}(t)$ by time $t$ under OPT, and let $V^{\prime}_{\leq p}(t)$ denote the total volume of jobs in $J_{\leq p}(t)$ that have been started by time $t$ under NSJF.
+Lemma 27 (Volume Matching).
+For every $t\geq 0$, define $t^{\prime}=A^{-1}(m(t+\tau))$. Then for every threshold $p$, we have $V^{\prime}_{\leq p}(t^{\prime})\geq V_{\leq p}(t)$.
 Proof.
-Fix an arbitrary $p$, let $t_{0}$ be the last time some machine is idle by NSJF before $t^{\prime}$.
-Define $s_{j}$ as the latest start time, within the interval $(t_{0},\,t+\tau)$, at which NSJF schedules a job with $p_{j}>p$.
-If no such job exists, we set $s_{j}=t_{0}$ and conceptually assume that an infinitely large job starts at time $s_{j}$.
+Fix an arbitrary $p$, and let $t_{0}$ be the last time some machine is idle under NSJF before $t^{\prime}$.
+Let $j$ be the job with $p_{j}>p$ that is started latest by NSJF within the interval $(t_{0},t^{\prime})$.
+If no such job exists, we let $s_{j}=t_{0}$ for a fictitious job $j$ of infinite size.
 Since NSJF starts job $j$ with $p_{j}>p$ at time $s_{j}$, it must have already started all jobs with size at most $p$ released up to $s_{j}$ (due to its SJF policy). Thus, we have:
 $V^{\prime}_{\leq p}(s_{j})=\sum_{h\in J_{\leq p}(s_{j})}p_{h}\geq V_{\leq p}(s_{j}).$
-We now analyze two cases:
-Case 1: $s_{j}\geq t$. We straightforwardly have:
+We now analyze two cases. If $s_{j}\geq t$, we immediately have:
 $V^{\prime}_{\leq p}(t^{\prime})\geq V^{\prime}_{\leq p}(s_{j})\geq V_{\leq p}(s_{j})\geq V_{\leq p}(t).$
-Case 2: $s_{j}<t$. Then the time interval $s_{j}+\tau,t^{\prime})$ must be fully used by NSJF to process jobs with size at most $p$ (if any).
+Otherwise, if $s_{j}<t$, then since $s_{j}$ is the latest start of a job with $p_{j}>p$ after the last idle time $t_{0}$, all processing in $s_{j}+\tau,t^{\prime})$ is on jobs of size at most $p$, and no machine is idle in this interval.
+Therefore:
 $V^{\prime}_{\leq p}(t^{\prime})$ $\geq V^{\prime}_{\leq p}(s_{j})+\int_{s_{j}+\tau}^{t^{\prime}}a(y)\,dy$
@@
 $\geq V_{\leq p}(t).$
-In both cases, we conclude:
-$V^{\prime}_{\leq p}(t^{\prime})\geq V_{\leq p}(t),$
-which proves the lemma. ∎
-Denote $N_{\leq p}(t)$ (or $N^{\prime}_{\leq p}(t)$ below) as the number of jobs in $J_{\leq p}(t)$ completed by time $t$ by OPT (or NSJF, respectively).
-Note that without the subscription $\leq p$, we remove the condition on job size.
-We prove:
-Lemma 28.
-For every $t\geq 0$, define $t^{\prime}=A^{-1}(m(t+2\tau))$. we have $N^{\prime}_{\leq p}(t^{\prime})\geq N_{\leq p}(t)$.
+Step 2: Completion Matching.
+Let $N_{\leq p}(t)$ and $N^{\prime}_{\leq p}(t)$ denote the number of jobs in $J_{\leq p}(t)$ completed by time $t$ under OPT and NSJF, respectively.
+When the subscript $\leq p$} is omitted, we count all completed jobs regardless of size.
+Lemma 28 (Completion Matching).
+For every $t\geq 0$, define $t^{\prime}=A^{-1}(m(t+2\tau))$. For every threshold $p$, we have $N^{\prime}_{\leq p}(t^{\prime})\geq N_{\leq p}(t)$.
 Proof.
-Define $t_{1}=A^{-1}(m(t+\tau))$ and let $t_{2}=t^{\prime}$. By Lemma 27, we have the volume bound: $V^{\prime}_{\leq p}(t_{1})\geq V_{\leq p}(t)$ for every $p$.
-We claim the following: the $i$-th smallest job (in terms of size) that has started (either in processing or completed) under NSJF by time $t_{1}$, denoted by size $p^{\prime}_{i}$, must not exceed the size of the $i$-th smallest completed job under OPT by time $t$, denoted $p_{i}$.
-That is, we aim to prove: $p^{\prime}_{i}\leq p_{i}$ for all $i$.
-To prove this, assume for contradiction that there exists the smallest index $i$ such that $p^{\prime}_{i}>p_{i}$. Under this assumption, we have the following bound for OPT:
-$V_{\leq p_{i}}(t)\geq\sum_{j=1}^{i}p_{j},$
-since at least $i$ jobs of size at most $p_{i}$ are completed by time $t$ in OPT.
-Now consider NSJF.
-Since the $i$-th smallest started job under NSJF has a size strictly greater than $p_{i}$, all started jobs of size $\leq p_{i}$ must be among the first i1i-1 jobs.
-Therefore, we have:
+Define $t_{1}=A^{-1}(m(t+\tau))$.
+For each rank $i$, let $p_{i}$ be the size of the $i$-th smallest job completed under OPT by time $t$, and let $p^{\prime}_{i}$ be the size of the $i$-th smallest job started under NSJF by time $t_{1}$.
+Let $k$ and $k^{\prime}$ denote the total number of such jobs, respectively.
+We claim that $k^{\prime}\geq k$ and $p^{\prime}_{i}\leq p_{i}$ for all $i\leq k$.
+Assume for contradiction that $i\leq k$ is the smallest rank such that either $i>k^{\prime}$ or $p^{\prime}_{i}>p_{i}$.
+For OPT, at least $i$ jobs of size at most $p_{i}$ are completed by time $t$, so:
+$V_{\leq p_{i}}(t)\geq\sum_{j=1}^{i}p_{j}.$
+For NSJF, all started jobs of size $\leq p_{i}$ are among the first i1i-1 jobs (either because only i1i-1 jobs have started, or because the $i$-th smallest is strictly larger than $p_{i}$). Therefore:
 $V^{\prime}_{\leq p_{i}}(t_{1})\leq\sum_{j=1}^{i-1}p^{\prime}_{j}\leq\sum_{j=1}^{i-1}p_{j},$
-where the second inequality uses the assumption that $p^{\prime}_{j}\leq p_{j}$ for all $j<i$.
-Combining these two, we get: $V^{\prime}_{\leq p_{i}}(t_{1})<V_{\leq p_{i}}(t),$ which contradicts the earlier volume bound from Lemma 27.
-Hence, our assumption must be false, and we conclude: $p^{\prime}_{i}\leq p_{i}$ for all $i$.
-Finally, this implies that the number of started jobs with $p_{j}\leq p$ by time $t_{1}$ under NSJF is at least the number of completed jobs with $p_{j}\leq p$ by time $t$ under OPT.
-Moreover, since the number of active machines in NSJF is non-decreasing over time, all jobs started by time $t_{1}$ must be completed by time $t_{2}$.
-This completes the proof of the lemma.
+where the second inequality uses $p^{\prime}_{j}\leq p_{j}$ for all $j<i$. This gives $V^{\prime}_{\leq p_{i}}(t_{1})<V_{\leq p_{i}}(t)$, contradicting Lemma 27.
+Therefore $k^{\prime}\geq k$ and $p^{\prime}_{i}\leq p_{i}$ for all $i\leq k$, which implies that for every threshold $p$, the number of jobs of size $\leq p$ started by NSJF by time $t_{1}$ is at least the number completed by OPT by time $t$.
+Moreover, since the number of active machines in NSJF is non-decreasing over time, all jobs started by time $t_{1}$ are completed by time $t^{\prime}$.
+This completes the proof.
+∎
+Step 3: Bounding the Flow Time.
 Lemma 29.
@@
 Proof.
-Let $u$ denote the maximum completion time under NSJF. The flow times of NSJF and OPT can be expressed as:
+Let $u$ denote the maximum completion time under NSJF. The total flow times can be expressed as:
 $F(\textup{{OPT}})=\int_{0}^{\infty}\left(|J(t)|-N(t)\right)\,dt,\quad F(\textup{{NSJF}})=\int_{0}^{u}\left(|J(t)|-N^{\prime}(t)\right)\,dt,$
@@
 Appendix B Randomized Non-Preemptive Algorithm: Multiple Machines
-In this section, we extend Algorithm 2 from single-machine case to multi-machine case. The intuition is already introduced. We detail the dynamic but online stable algorithm in Algorithm 10.
+In this section, we extend Algorithm 2 from the single-machine case to the multi-machine case.
+The intuition is introduced in Section 4.
+We detail the dynamic but online stable algorithm in Algorithm 10.
 Include : Algorithm 1
@@
 On job $i$ release:
-   Classify job $i$ large or small based on Partition($i,\ell$) with $\ell=\lfloor\sqrt{nm}\rfloor$;
- if $i$ $i$ is classified as large then
+   Call Partition($i,\ell$) with $\ell=\lfloor\sqrt{nm}\rfloor$ to classify job $i$ and update related jobs;
+ if $i$ is classified as large then
       Sample $w_{i}\sim\mathrm{Unif}\{1,2,\dotsc,\lfloor\sqrt{n/m}\rfloor\}$ ;
@@
       Insert $j$ into $\S_{2}$ to start at $t$ on $m_{j}$ ;
+    // Overlapping jobs are shifted right.
    end for
@@
 Theorem 11.
-There exists an online polynomial-time randomized non-preemptive algorithm for total flow time minimization, e.g., Algorithm 10, that is $O(\sqrt{n/m})$-competitive against the preemptive offline solution.
+Algorithm 10 is an online polynomial-time randomized non-preemptive algorithm for total flow time minimization that is $O(\sqrt{n/m})$-competitive against the preemptive offline solution.
 Proof.
-The algorithm is online simulatable and runs in polynomial time, for the same reason as in Lemma 5.
+The algorithm is online stable and runs in polynomial time, for the same reason as in Lemma 5.
 We use $S$ and $L$ to denote the sets of small and large jobs, where $S$ also includes proxy jobs.
@@
 For small jobs, by Lemma 1, and Lemma 2, we have:
-$F(\S_{1})=O\left(n\cdot\frac{1}{\sqrt{nm}}\right)\cdot F(\textup{{OPT}}(S))\leq O(\sqrt{n/m})\cdot\textup{{OPT}}.$
+$F(\S_{1})=O(\sqrt{n/m})\cdot F(\textup{{OPT}}(S))=O(\sqrt{n/m})\cdot\textup{{OPT}}.$
 For large jobs, we bound both the external delay and the self-delay.
@@
 Then:
-$\mathbb{E}[|D_{k,l}|]=\sum_{k=1}^{\lfloor\sqrt{n/m}\rfloor}\sum_{l=1}^{m}|D_{k,l}|\cdot\Pr[w_{j}=k]\cdot\Pr[m_{j}=l]\leq O(\sqrt{n/m}),$
+$\mathbb{E}[|D_{w_{j},m_{j}}|]=\sum_{k=1}^{\lfloor\sqrt{n/m}\rfloor}\sum_{l=1}^{m}|D_{k,l}|\cdot\Pr[w_{j}=k]\cdot\Pr[m_{j}=l]=O(\sqrt{n/m}),$
 so $\mathbb{E}[\delta_{j}]=O(\sqrt{n/m})\cdot p_{j}$.
-For the self-delay, if $j$ is proxied by $j^{\prime}$, we count $r_{j^{\prime}}-r_{j}$; otherwise, we count $C^{\prime}_{j}-r_{j}$, where $C^{\prime}_{j}$ is the completion time of $j$ immediately after insertion.
+For the self-delay, if $j$ is proxied by $j^{\prime}$, we count $r_{j^{\prime}}-r_{j}$; otherwise, we count $\hat{C}_{j}-r_{j}$, where $\hat{C}_{j}$ is the completion time of $j$ immediately after insertion.
 This self-delay is denoted by $\hat{F}_{j}$.